A Complete Bibliography of The Bell System Technical Journal, 1960–1969

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: http://www.math.utah.edu/~beebe/

09 April 2024 Version 1.08

Title word cross-reference

1/f [NM67]. _{2} [NG67]. L_{2} [San64d]. DC [SW69b]. M [Jac67, Kad65b]. μ [Sem65]. N [Llo67, Wyn67]. N = 2^{n} [Dra67b]. z [RSR69].

-ary [Jac67, Kad65b]. -boundedness [San64d]. -sphere [Wyn67]. -transform [RSR69].

0.2-Megabit [GHS64]. 000-cps [HR63].

1-Mev [RGS63]. 1/4 [MM66a]. 1/4-Watt [MM66a]. 1 [?]TLFriis:1966:TTS. 10-kw [SB61]. 100Å [GN65]. 125-Kmc [Ste65]. 14-Watt [Kib65a]. 140-Kilomegacycle [Sha63]. 15-Watt [Kib65b]. 150-mc [MV61]. 150-megacycle [KS63].

2 [CCJ60, Cur60, FJJK66]. 2.5-mc [War60]. 20-20 [HR63]. 2390-mc [DKS61]. 2A [Sta65].

3 [HHJP68]. 39 [Ano60g]. 3A [BGLS68].

4 [Ano69k, BBHH69]. 4-gc

[Gig65, UCE^{+}63]. 4-Watt [MM66a]. 42 [Ano64j]. 43 [Ano64k, Ano64l]. 45 [Ano66l]. 47 [Ano68k]. 48 [Ano70]. 4A [PS69].

5-Gigacycle [Sch63]. 50-mc [Kib65a].

6328Å [Col65].

80-Megabit [Kib65b].

9-Milliwatt [Sch63]. 90- [Sha63]. 960-mc [SB61]. 961-mc [US61].

A5 [BH62a]. Aberrated [Mar67a]. Aberration [Mar67b]. Above [Til69]. AC [Hei68, Log61]. Accumulation [Man69, San67a]. Accuracy [Des65, San67c]. Acoustic [CG65a, CG66, DG67, Mat61, Rab68]. Acting [Wel66a]. Active [Ano61h, Eld68, FJL68, Par68b, Row64a, Row64b, San60, San61]. Activity [Mau67]. Adaptive [Ger69a, Luc66, Luc68, Mor65, Rud69, SP66, Son67]. Added [Bra67]. Adding [Wol69]. Additive [Wyn69a]. Adjustment [Kur62]. Administration [BFH^{+}69]. Adopting [Ano66a]. after [Lan60]. Against [Ano62a, KK62]. Agent [Ber65a]. Aggregates [Bur60]. Aging [DGS68]. Aided [KML69, ZI68]. Air [Lep68]. Air-Insulated [Lep68]. Albersheim [Ano60g]. Algebra [Bro63, BHT64, Hyd64]. Algebraic [Ben62a]. Algorithm [Kat65, RSR69]. Algorithms [Kru69]. All-Weather [CG65b]. ALPAK [Bro63, BHT64, Hyd64]. alphabets [Mac65]. Alternate [Web62]. Alternating [Mil64a]. Alternating-Gradient [Mil64a]. Alternative [FS60]. Aluminum [GK62]. AM [Cro66, Gib60, Sun61]. AM/PM [Cro66]. Ambiguously [JJ68]. among [GK64b]. Amplifier [BR63b, BG69, Bra68a, Ham64, Kib65b, Kib65a, KU61, UCE^{+}63, Mac60]. Amplifiers [Kur62, Kur65, SBF68, Tho68]. Amplitude [Dra67a, Dra67b, Dra69, Fen69, GP60, Rus66, Tho63, You63b]. Analog [CF68, Goo69a, Sun64b, Wyn68, WZ69]. Analog-to-PCM [Goo69a]. Analogues [Mur68]. analogy [Ric62a]. Analysis [AG68, Ano64k, Blu69, BH62b, BH65, Bra68b, Bur65, CEZ65, Dra68, FRY63, GWB67, Goe69, Gol62, GS64a, GS64b, Hae63, Hig62, Lee67b, Leg60, Liu68b, Luc63, MM64, Mor62a, Mur68, Nar67, Par68a, PW68, Rut68a, Sim67, Sle64, WC69, Wyn69c, ZI68, LP61, LP62, SP61]. Analytical [O’N66a]. Andover [Cur63, Gig65, SCM^{+}63]. Angle [Bar64b, DH63, KR62, Nel63, PE69, Ung66]. Angle-Error [Nel63]. Angle-Modulated [Bar64b, PE69]. Anisotropic [MM68]. Annular [Leg60]. Anomalies [Gra66]. Antenna [Bay68, CEZ65, Coy65, CHH61, CT69, DGPS65, DBKW63, GT65, GKLL63, GP63, HLT63, HW65, KNG61, Lee67a, Lee67b, LNI63, LM69b, Mor65, NC65]. Antennas [CG65b, HS64, Jag65, Pop68, VZ67, Wu69, ZI68, Zuc68, BEH^{+}63]. Antireflection [Sch66]. Aperture [Li63, McC69]. Aperture-Limited [Li63]. Apertures [McC65, Nag68]. Apparatus [DR69, FGK^{+}64, LMS^{+}69]. Apparent [MP69]. Application [BR66, CMS^{+}64, Goo69a, Kah64, KS67, Kru69, San65a, Sch62b, RSR69, Ric68]. Applications [Ben61b, BS64, Bob67, Che64, Hin60, Kla61, Liu69, McD66, Mil69b, Til69, Wol69]. Applied [Bar62, Suy65, Ano70, Lee69b]. Applying [Arm61]. Approach [Buc69a, FS60, HU67, Par68a, Rab68, SS69b]. Approximate [Blu69, Rai68a, Row62, You63a]. Approximating [Fla60, Fla62]. Approximation [AS66, HS69, Jag65, Kyo69]. Approximations [Spa68]. Arbitrary [Bur65, Llo67, Mor65, Row64b]. Argon [MLW67]. Arise [SB64]. Arises [Ano66l, San66a]. Arising [NM64, Ben66a]. Armorless [Ano64l, Bhu60, BNR64, LP64]. Arrangements [AEH^{+}69]. Array [BCDY68, CLW69, CBL^{+}67, CL69, CKT63, Jag65, VZ67]. Arrays [AG68, GW68a, Mor65, WG68]. Art [Ger66]. ary [Jac67, Kad65b]. Aspects [BST65, HF61, KL61, Kla61]. Assembler [MOP64]. Associated [BR63a, Fri68, Rai66a, SS65]. Astigmatic [Mar64a]. Asymmetric [McC69]. Asymptotic [Ben61a, Ric68]. Asynchronous [AMS65, BH62b]. Atmosphere [Bul63, CH66, Hog63]. Atmospheric [SC65, SC67]. Atomic [Ber65a]. Attaching [Coy63]. Attack [Ste61]. Attenuation [CH66, Con66, Hog69, KM61, ST69, Ste65, Wil69, You65a]. Attitude [Gla67, LZ64, PWY63, TC64]. Au-n-Type [Kah64]. Autocorrelation [BR63a]. Automata [HU67, Lee60]. Automated [HR63]. Automatic [BR66, GL61, IKM67, JM62, Luc65, LR67, TU62, Web62]. Automatically [Eva69]. Autotrack [CL63, NC65]. Auxiliary [HW61]. Avalanche [GS66, MM66a, SR67]. Average [Cut61, Rap67]. averages [Ric62c]. Axes [NC65]. Axial [Whi66]. Axis [Ber65b, Hil63, Mar67b, Mar64b, Rai65, Rai67a, Rai67b, Yu63]. Axis-Crossing [Rai65, Rai67b]. Axis-Crossings [Rai67a].

B.S.T.J. [AS66, Ano61g, Ano64a, Ano66a, AK65, Bar65, Bec64b, Bec64a, BKS66, BR66, Ber64b, Ber64a, Bur66, CG64, Cop67, CBL^{+}67, Dal67, Dar63a, EMR66, Enl69, Ger69b, GS69b, Gor64b, GL64, Hei68, Hog63, Hub67a, JDC65, KS67, Kah67, Kog64, Kog65a, Lep68, LPSL66, MLW67, MM66a, Pfa64, Pro67a, Rai65, Rai67b, Rai68a, San64a, San64b, San65a, San66a, Sca63, Sch63, SH67, Sha63, SP66, SC65, Wel66a, Whi66, Wyn64, Zad66a]. Back [RC65]. Back-scattering [RC65]. Background [CCKT63]. Bacteria [Ste61]. Balanced [Kur65, RE62]. Balking [Lin64, Tak61]. Band [Cha66, MB61, MM66a, Rai66b, Ste65, Wan68, Wyn66b, Lan60, LP62].

Band-Limited [Cha66, Wyn66b, Lan60, LP62]. Band-Rejection [Wan68]. Band-Splitting [MB61]. BandGaussian [Rob67a]. Bandlimited [Rut68b, SK65, San66b]. Bandwidth [Kib65a, Mou67]. Bank [BH62a]. Barrier [AART68, Ano68k, Kah64, KL65, LS68, Sch66, Sle62]. Base [BR63b]. Based [KH66, Par68a]. Basic [GLR69, Kur69]. Basilar [Fla60, Fla62]. Basis [Bra65a, Dar63a]. Battery [BM63]. BCH [Ber67a, Wol69]. Beam [Chu66, CG65a, Dai68, Dal68, Dal69, GW68b, GS69a, Lep66, Lep68, Mar66b, Mar65b, MM66b, Ste66, Gor60]. Beam-Lead [Lep66, Lep68]. Beam-Waveguides [Mar66b, Mar65b]. Beams [SC65, SC67]. Behavior [Fla62, GWES64, Llo67, Ben61a]. Bell [Ado63, And63, Ano60h, Ano60i, Ano60j, Ano60k, Ano60l, Ano60m, Hig65, Jak61, Kel64, Lap13b, Lap13a, Nas68, NEH68, WC69]. BELLBOY [KS63]. Bends [Con66, Mar69a]. Between [Bat69, Cut61, Dar63a, Koe67, LK61, Lee66, Mar66c, Cur62, Fla62, HGG62, Ric62a, Ric62c]. Beyond [SS69b]. Bias [Ziv69]. biggest [Lap13a]. Binary [Aik67, ABDS65, BS63, BR63a, Buc66, Hub66, Hub67a, Hub67b, Kad64, McC68, SK65, Sal67, Ton66, Kad65a, Mac65, Sch65]. Binocular [Jul60, Jul67]. Bipolar [May62]. Bird [Ung66]. Bit [BC69, SS69b]. Block [Gol66, Joh64, KLV61, Pil68, Sav66b]. BLODIB [Gol66]. blue [Lap13a]. Blurred [Sle67]. Bode [Lee69a]. Body [FRY63, Yu65]. Bombardment [RGS63]. Bond [FGVH62]. Boolean [Sch62a]. Boosting [Mea68]. Boresight [War61]. Both [Kla60]. Bound [Sav66b, SK68, Wyn64]. Bounded [Wyn65, WZ69]. Boundedness [San65a, San64d, San65b]. Bounds [Dar63b, Gra66, PR69, Sle63, Wyn66a, Ziv69].

Bowker [Ano64l]. box [Lap13a]. Breakdown [Bug65, SR67]. Bremsstrahlung [Mar62, Mar63]. Bridge [Log61]. Brief [AK65, Bar65, Gig65, JDC65, KL65, Kog65a, Rai65, San65a]. Briefs [AS66, Ano64a, Bec64b, Bec64a, BKS66, BR66, Ber64b, Ber64a, Bur66, CG64, Cop67, CBL^{+}67, Dal67, Dar63a, EMR66, Enl69, Ger69b, GS69b, Gor64b, GL64, Hei68, Hog63, Hub67a, KS67, Kah67, Kog64, Lep68, LPSL66, MLW67, MM66a, Pfa64, Pro67a, Rai67b, Rai68a, San64a, San64b, San66a, Sca63, Sch63, SH67, Sha63, SP66, SC65, Wel66a, Whi66, Wyn64, Zad66a]. Broadband [BH62a, BG69, Bod69, DA63, EH65a, Kur69, Osb69, Pra68, SG61, Tho68]. broke [Lap13a]. BSTJ [Ano60g, Ano64k, Ano64l, Ano64j, Ano66l, Ano68k, Ano70, Gig65, KL65]. buccaneering [Lap13a]. Building [Oss64, SW68]. built [Lap13a]. Bulk [BCDY68, Irv62, Kur67]. Burst [Ell63, Gil60]. Burst-Noise [Ell63, Gil60]. Bus [CHK64]. Business [WC69]. Busy [Ric62b, Rio61].

C [Ano60g, Ano70]. Cable [And69, Bhu60, BM64, BG61a, BNR64, BDV64, BAMS64, DR69, Ehr64, EFK^{+}64, FC68, GPZ63, Gre64, Gri64, HVW64, LP64, Pie66, Ano64l]. Cadmium [BM63]. Calculated [Iwe69]. Calculation [Hil63, Man63]. Calculations [GM67, Kla61, LW60, RJ61]. Calibration [Fre69]. Call [ADH^{+}69, CDGH64, GHS64]. Calling [Sch60]. Camera [BE69, CLW69, CBL^{+}67, CL69]. Cameras [War61]. Canceller [SP66, Son67]. Cantilever [Pee63]. Capabilities [AGN60, Mor62a, Wyn65]. Capacitance [Bur65, KU61, Kur62, MS60, Mor62b]. Capacitor [Kah67]. Capacity [Ano64a, CCJ60, Gil60, GW68b, Kur66, Ski66, Tab67, Wyn66b, San66b]. Capture [Bar62]. Carbon [Sla64]. Carlo [FGVH62].

Carrier [AEGH63, BST65, BI66, BH65, CC63, Cur62, Ell62, FP65, GR67, GBW66, HW66, LW65, SD61, Til69, TY65, Van60]. Cascade [Wyn69c]. Cascaded [Kin66]. Cassegrain [CEZ65, DGPS65, ZI68]. Cassegrainian [HS64]. Cathode [Coo61, Mau67, MP67, Pfa63]. Cathode-Ray [Pfa63]. Caused [Mar69c, MD69]. Cavities [Guc62, Whi66]. Cell [Kle61, RGS63]. Cells [BM63, GS64b, SGB^{+}63]. Center [DRRS64]. Centimeter [HS61]. Central [CMS^{+}64, FGM64, GS64a, HTU64, KH66]. Certain [Buc69b, Gra66, Mac60, Wol69, Ric62a]. Chain [BKR63, Kin62a, Kin62b, Kin64]. Changes [MD69]. Channel [Aik67, BS63, BH62a, DKS61, Fen69, FA64, Gil60, Mar61a, McC68, McD66, Pra69, Sle68, Sta67, Wyn66b, Wyn68, Wyn69a, Pil69, San66b, Wyn69b]. Channel-Dropping [Mar61a, Sta67]. Channels [CCJ60, Ell63, For69, Ger69a, HW61, LR67, Rud69, Wyn64, Fal69]. Characteristics [Bor63, Eva69, FBL62, Gre69, HLT63, Hsi67, Kur69, LS68, Nas68, Pee61, RR62, Rob67a, Suy65, SR67, Web62, Ano68k]. Characterizing [Lee66]. Charge [Dew60, GC68, Kah67, Koe67, SS67]. Chi [Rob69a]. Chi-Square [Rob69a]. Chirp [Ano60g, KPDA60, RSR69]. Circle [Gor64b]. Circuit [AL60, CE64, CL60, HW66, Hin60, KML69, Kur62, Lew64, LW65, Mur68, Par68a]. Circuits [Ado63, BBC^{+}64, CMS^{+}64, CUM68, DDH64, Dar63a, Dar63b, Gia60, GS69b, Hel66, Koe67, Kur69, Lep68, McC62, Sch69a, SGB69, Sch69b, Sin66, SS60, TU62, UW69]. Circular [AG68, Con66, Enl63, Goe69, Li65, LU64, Mar61a, NS65, Sta67, Wei65]. Circular-Electric [Mar61a, Sta67]. Circular-Harmonic [Goe69]. Class [AMS65, Bar65, Hyd64, McC62, Mel68, NM64]. Classes [For69]. Classic [Koe67, San67b]. Click [Sch68a]. Clogston [Man63]. Clutter [Rai66a]. Co [Pra69]. Co-Channel [Pra69]. Coated [Con66]. Coating [Sch66]. Coaxial [Ano69k, BBHH69, DR69, FC68, Man63, Pie66]. Code [Buc66, Dav62, LS66, May62, McD66, O’N66b, Pro67b, Sch60, Mac63]. Codecs [EH65a]. Coder [BC69, Bro69, MSV62]. Coders [Iwe69, Pie68]. Codes [Ber67a, Buc69b, Eck60, Ell63, Ell66, Mac64, Pie68, Sle60, Sle68, Ton66, Wel66b, Wol69, ET69, GM66, Mac61]. Coding [Buc67, Fra68, McM69, Sav66b, Ski66, Wyn66a, Fal69]. Coefficients [Bro63, BHT64, Rob67b, Mea62]. Coherent [DeL65, Enl67, Enl69, Hub66, Hub67a, Hub67b, HM67, Nel63, Pra69, Rus66]. Coinage [KS66]. Coincidence [Mel68]. Color [Ski66]. Comb [Che64, HDSD64]. Comb-Type [Che64]. Combinatorial [Ben67a]. Combined [Rap65]. come [Rio61]. Command [CCM63, EMT63, MMW^{+}63]. Comment [Ano62a]. Common [Cur62, Til69]. Communicaions [RK63]. Communication [Ano66l, CG65b, CHH61, Cur62, Gil67, Gol66, HBK63, Hub67a, HGW^{+}67, Luc65, Luc66, LR67, McM69, Pop68, Ros62b, RP68, Rud69, SK65, San66a, Sle63, Til68, TW64, UCE^{+}63, Wol65, Wyn66a, Wyn68, Ric62a, WZ69, Wyn69b]. Communications [BH63, DHWM63, Ell65, GT65, Kar66, Lub68, Mag68, OS63, Rap67, RR62, SCM^{+}63, TS63, Tho65, Web62, Web64]. Compact [GT65]. Companded [BC69, MSV62]. companding [Lan60]. Compandor [GR67]. Compandored [Dos65]. Comparator [Byr62, Gol62]. Comparison [Bro67b, Jac67, Mar66c, Mil64c, Sch68a]. Comparisons [Ben66b]. Compatible [KS66]. Compiler [Gol66, KLV61, MOP64, Mat61].

Compiler-Assembler [MOP64]. Complex [TU62]. Complexes [Lie60, Ben64b]. Component [PW63]. Components [Hit63, Ros62b, Ste66, Tay68a, Tay68b]. Composed [Ric63]. Compression [Mar61b]. Computable [Rad62]. Computation [Gra65b, Gra65c, Rob69a, Rut68b, Sav66a, Tay68a, Tho63]. Computer [BF60, Bro63, BHT64, CBHR63, GHS69, Gib60, Goe69, Gol66, GS64a, Hyd64, Jul60, KML69, Lin65, O’N66a, Rap64, Rob69b, ZI68]. Computer-Aided [KML69, ZI68]. Computer-Generated [Jul60]. Computer-Operated [GHS69]. Computers [And63]. Computing [Buc66, Tay68a]. Concentrated [SK61b]. Concentration [Cut61]. Concentrator [JJ60, MR60, Sta65]. Concentric [Li63]. Concept [Koe67]. Concepts [EH65b]. Concerning [Enl69, Rai68a, San65a]. Condition [San64a, San64c]. Conditional [Mou69]. Conditions [Pro67a, SK61a, SK61b, Zad66a, San69a]. Conductance [NG67]. Conducting [MM66b]. Conduction [FP69, Hei67]. Conductivity [Cut61]. Conductor [Bhu60, Man63]. Configuration [Mel68]. Confocal [BG61b, BK62, McC65, McC69]. Congestion [Wol65]. Conical [HLT63, Nag67]. Conjugate [AK65]. Connecting [Ben62c, Ben64a, Ben65b, Ben66b, Ben67c, Ben67b, Ben62a, Ben62b, Ben63a, Ben63b, Ben64b]. Connection [Pau62]. Connections [Mil64c, Nas64, Nas68]. Connector [GPZ63]. Connectors [CE64]. Considerations [Bus60, Cur60, DP68, GS61, HKM^{+}63, KS66, Kle61, Pra67, Pra68, Pra69, PE69, SW68, SG67]. Constant [FHS67, Lee69a]. Constants [Kla61]. Construction [Eck60, PW63, RC61, SR63]. Contact [Kel64, Sha63]. Contacts [HP60]. Contained [FA64]. Containing [Hei67, San64a, San64c]. Continuous [Enl63, SH67]. Continuously [GL64, Rud69]. Contraction [Hol67, San65a]. Contraction-Mapping [San65a]. Contrast [HH65, Hei66]. Contributors [Ano60a, Ano60b, Ano60c, Ano60d, Ano60e, Ano60f, Ano61b, Ano61c, Ano61d, Ano61e, Ano61a, Ano61f, Ano62b, Ano62c, Ano62d, Ano62e, Ano62f, Ano62g, Ano63a, Ano63b, Ano63c, Ano63d, Ano63e, Ano63f, Ano64b, Ano64c, Ano64d, Ano64e, Ano64f, Ano64g, Ano64h, Ano64i, Ano65a, Ano65b, Ano65c, Ano65d, Ano65e, Ano65f, Ano65g, Ano65h, Ano65i, Ano65j, Ano66b, Ano66c, Ano66d, Ano66e, Ano66f, Ano66g, Ano66h, Ano66i, Ano66j, Ano66k, Ano67a, Ano67b, Ano67c, Ano67d, Ano67e, Ano67f, Ano67g, Ano67h, Ano67i, Ano67j, Ano68a, Ano68b, Ano68c, Ano68d, Ano68e, Ano68f, Ano68g, Ano68h, Ano68i, Ano68j, Ano69a, Ano69b, Ano69c, Ano69d, Ano69e, Ano69f, Ano69g, Ano69h, Ano69i, Ano69j, Cro67]. Control [AC69, BD60, Bri68, BQTY69, CMS^{+}64, Dal68, DRRS64, Gla67, Kyo69, LZ64, MM69, Mil64b, NC65, PWY63, Rap68, San64f, TW64, Web64, Wel66a, Ben66a]. Controlled [Bar62, Byr62, Gol62, Koe67, LM69b, SBF68, You65a]. Controller [MR60]. Controls [CK68]. Convection [Hei67]. Convectively [Ber64b]. Conventional [Mil64a]. Convergence [Lyt68]. Convergent [Mil64a]. Conversation [Bra69, Cro66, Mar69c, MD69]. Conversations [Bra68b, EM63, Kle67]. Conversion [Enl63, Glo67c, Log67, Mar61d, Mil63, NS65, Ung61a, Ung61c]. Converter [US61]. Convolutional [ET69]. Cooling [Hei67]. Copper [Kla61]. Core [GS65b, Leg60]. Cores [Sla64]. correcting [Mac61]. Correction [Arm61, Dal69, Enl69]. Correlated [Cro67, Mus63]. Correlation [Hub67a, SK65, Ano70, Lee69b]. Cosine [Jag65]. Cost [Gil67, Rap67]. Counter [Ber64a]. Counter-Flowing [Ber64a].

Counters [Ble68]. Coupled [BR63b, Kog69, Mil69b, Row62, You63a, You63b]. Coupler [Mar69b]. Couplers [ML68]. Coupling [Bat69, Lie60, McC65, McC69, Mil68]. Covariance [Ben61b, Kad67b]. Covered [GW68a, Nag68]. coverings [Wyn67]. CPC [MP67]. cps [HR63]. Criteria [Kyo69, Lyt68]. Criterion [KG66, Shn67]. Critique [Pro67a]. Crossbar [Zup60]. Crossing [Rai65, Rai67b, SK68]. Crossings [Lee67b, Rai67a]. Crossovers [Lep68]. Crosstalk [Mar61c, Mar61e]. Crystal [Bur60, FGVH62, War60]. Crystallization [SK61a, SK61b]. Crystals [Lew61]. Cube [SS67]. Current [MP67, MLW67, SR67, Van60]. Current-Carrier [Van60]. Current-Voltage [SR67]. Curved [Bat69, Mar64b]. Curves [Goe66, TK60]. Customer [Gre69, Hel66]. Cutoff [HDSD64, Kib69]. CW [Cop67, Kib65a, Rob67a]. Cyclic [CM69, Ton66, Wel66b, GM66, Mac65]. Cylindrical [Con66, McC65, McC69, Thi69].

D. [Ano64k]. Damping [Yu63]. Darlington [Ano60g]. Data [AGN60, BKS66, BS63, Buc67, Cha66, CT67, Cha67, Dav64, Ell65, Fre69, FA64, GWB67, Ger69a, Gib60, GK64a, Gol66, Hae63, Hel65, Hil63, KH66, Liu68b, Luc63, Luc68, Maz68, Mor62a, Rap64, Rap65, Sal67, SS69a, Sav67, Sch68b, SH67, Smi65, Smi68, TW64, TY65, Wyn68, Zad66b, Gop69, Lev61, WZ69]. Data-Rate-Limited [Hel65]. DC [SW69a, Zad66b]. Decay [Pfa63, Yu63]. Decision [Hub66]. Decoding [Mac64, Sav66a, Wyn65]. Decomposition [Gra65a]. Definitions [Wol65]. Deflecting [Kur66]. Deflection [Nel64]. Deflector [Ano64a, Ski66, Tab67]. Deformation [Glo67a, Mar66d]. Deformations [Dai68]. Degradation [Gig65, Rob68, RGS63]. Delay [EM63, GW68b, Hae63, Hel66, Kle67, Lee66, Luc63, Mil63, Rap65, RK63, Tak62, Tak63a, Var66, Sch65]. Delays [Rio61]. Delta [AFMP69, Goo69a, Goo69b, Iwe69, O’N66a]. Delta-Modulation [Iwe69]. Demagnetizing [Dov67]. Demodulation [Dar64, PW68]. Demodulator [GC63]. Demodulators [Rut61]. Demultiplexer [Wit65]. Dense [RT69, SK69]. Densities [Bar64a]. Density [BR63a, Cro67, Lee67a, Lee67b, MP67, MLW67, PR69]. Dependence [GN67]. Dependent [Cha69a]. Depletion [LW60]. Deposition [BCD64]. Depth [Jul60, SC67]. Describing [MM69]. Description [FG69, HM63, SH63]. Design [Ano60g, Ano64l, Bay68, BG69, Bod69, BKRS68, BNR64, BH63, Bus60, BAMS64, Byr62, CMS^{+}64, CK68, CG61, CEZ65, Coo61, Dei60, DBKW63, GT65, GBW66, Gla67, GK64a, HM63, HW66, Hig62, HKM^{+}63, KS63, KS66, Kla60, KPDA60, KML69, Kur65, Kur66, Lim69, LW65, Nel63, Osb69, Par68b, PW63, SK65, SR63, Smi68, SG67, Wan68, War60, Yu65]. Designed [Tay68a, Tay68b]. Designs [RC61]. Detail [Hei66]. detectability [MM61]. Detecting [Rai66b]. Detection [ABDS65, CG64, FA64, Mar61c, Nel63, O’N69, SK65, You65b]. Detector [Rob67a, Wyn69c]. Detector-Filter [Wyn69c]. Detectors [MS65, MS62]. Determinability [For69]. Determination [Bri66, CBHR63, Suy65, TC64]. Determined [NG67]. Development [GPZ63, Zup60]. Developments [Kel64]. Deviations [Cro66, You65a, Zuc68]. Device [Bob67, FLLR60]. Devices [CDJ69, Eld68, Gla67, KS67, MW67]. Diagnosis [TU62]. Diagnostic [CT67, KH66]. diagonalization [Kad67b]. Diagram [Gol66, Gor64b, KLV61]. Diagrams [Dav64]. Diameter [MD69]. Dielectric [Bhu60, Con66, GW68a, GCE69, Goe69, MS64, Mar66a, Mar69a, Mar69b, Mar69c, MD69, Mar69d, McK67, MM68, Mor64b, Nag68, WG68, Wu69]. Dielectric-Coated [Con66]. Dielectrically [HDSD64]. Difference [Bay68, Wel66b, GM66]. Difference-Set [Wel66b, GM66]. Differential [ABDS65, LM69a, McD66, O’N66b, Pro67b, You63a, Ben62d, Mea62, San67c, SS68]. Differentially [Hub66, Hub67a, Hub67b, HM67]. Differentially-Coherent [Hub67b]. Diffraction [GK62, Li65]. Diffused [BF60, Cut61, Irv62, LW60]. Diffusely [Enl67]. Diffusion [Ros62a]. Digit [Bor63, GP64]. Digit-Stimulating [Bor63]. Digital [AS65, Ano64a, Arm61, Bro63, BHT64, BKR63, CT67, CF68, Dav64, DSW66, Dos65, EMT69, Eng65, Fra68, FC68, GP63, Gol66, Hyd64, KH66, Kur66, LM69a, Luc65, Luc66, Lun69, MS66, MRS69, Mou67, Nel64, PW68, Pie69, Rap64, San67a, San69c, Sav67, Sch68a, Ski66, Sun64a, Tab67, Wit65, San69a, Wyn69b]. Digital-Filter [San67a]. Digitally [Cha69b]. Dilute [SK61a]. dimension [LP62]. Dimensions [Sle64, Woo64]. Diode [Ano68k, BCDY68, Cha69b, CLW69, CBL^{+}67, CL69, Dra69, GS61, Hin60, JDC65, Koe67, LS69, LS68, MM66a, Mur68, San62b, Sch63, San62a]. Diodes [BF60, Cop67, Dra68, GR67, GS66, Ham64, KD64, KL65, Mor62b, Sha63]. Dip [Mau67]. Diperiodic [Woo64]. Direct [Bro69, Ell62, Ott62, PDS64]. Direct-Feedback [Bro69]. Direct-Reading [Ell62]. Directed [Mag68]. Directional [Abe67, Mar69b, MM69, Mil64b, ML68]. Discharges [MLW67]. Discontinuity [CNN67]. Discrepancy [Wyn65]. Discrete [Bar64a, Mus63, Rai68c, Fal69, Wyn69b]. discrete-time [Wyn69b]. Discrimination [KK62, Ano62a]. Discussion [Mar63]. Disk [FL64]. Dispersive [Ger69a]. Displacement [Fla60, Fla62]. Display [Bur66]. Distance [Gar60, MS64]. Distort [San64e, San63a]. Distorted [Ber64b]. Distortion [BH62b, Bun69, Hol68, Mar69d, Mil63, Nar67, Pil68, Rap65, Ric69, Rut68b, San63c, Sun61, Sun64b, Tho68]. Distributed [KG66, ML68, Par68b, Spa68]. Distribution [AG62, Bar65, GP60, MS60, Rob69a, SK68, Mac63, Ric68]. Distributions [Dew60, Fen69, Tak62, Tak63a]. Distributor [FGM64]. Dither [Lim69]. Diversity [Gil69, Rap68]. Division [BWY69]. Domain [Rab68, Spa69, San64c]. Domains [Bob67, Kur67, Thi69]. Domestic [Til68]. Doped [BF60, GR67]. Double [Ung60]. Doubling [CCJ60]. Downconverter [OKS69]. Drift [Yu63]. Drive [Bur65]. Driving [Lew61, San60]. Driving-Point [San60]. Dropping [Mar61a, Sta67]. Dual [DKS61]. Dubrovskii [BS64]. Due [Cro66, DH63, Gig65, Lio66, Mil63, Mor64b, NM67, Oss64, BT63]. Duration [Lee67b, Rai66a]. During [Jul67]. Dynamic [Bri67, San69b]. Dynamical [SB64]. Dynamics [FRY63, Kur67].

E6 [BGK60]. Ear [Fla62]. Early [And63, Ung66]. Earth [SCM^{+}63, CG65b]. Echo [BR66, BH63, RK63, SP66, Son67, CHH61, DKS61, De 61, Jak61, Kib61, KNG61, Ohm61, Rut61, RJ61, SB61, US61, War61]. Echoes [EM63]. Economical [SS69b]. Editor [Ano61g]. Effect [HS61, Hub66, Hub67a, Hub67b, Lew61, Mar67a, MS62, You63a]. Effectiveness [TW64]. Effects [Bra67, BMP67, Bug65, CH68, Coy65, Dos65, EM63, Fla62, Fri68, LS69, MM68, MW67, PBBS63, SK61a, SK61b, Ste66, Var66, Weg60, WG68]. Efficiency [Dra67a, Dra67b, Pfa64]. Efficient [BKS66, Bod69, Osb69, RP68]. Eigenmodes [McC65, McC69]. Eigenvalues [NM64, SS65, Coc65]. Elastic [Wei65]. Electric [Con66, Enl63, Lee67a, LU64, Mar61a, Sta67]. Electrical [BF60, HLT63, Lew61, Llo67, Mil64c, NG67]. Electro [DA63, Kur66]. Electro-Optic [DA63, Kur66]. Electrode [Dew60, FL64, Lew61]. Electrodes [Llo67]. Electroluminescence [Pfa64]. Electroluminescent [Ger66]. Electromagnetic [CS65, CEZ65, MM69, MM68, RC65]. Electron [GC68, HH65, Hei66, Hei68, HVW64, MM66b, RGS63, Gor60]. Electronic [Hig65, KH66, HTZ64, UV64]. Electronically [VZ67]. Electronics [SR63]. Electrooptic [GR63]. Electrostatic [Coo61]. Element [Mou69, San62b, San64a, San64c]. Elementary [Man63]. Elements [Tho65, Tru60]. Eliminating [Tho68]. Embedded [Mar61c]. Emission [KR62, Mar62, Mar63]. Enclosed [DeL65, DD68]. Encoding [Goo69a, Mou69, Pil68]. Ended [BCD64]. Energy [GWB67, Gil65, Lee67a, Lee67b]. Energy-Density [Lee67a]. engineer [Lap13a]. Engineering [CC63, Dei60, EH65b, KL61]. Enumeration [Ber67a]. Environment [Mor65]. Epitaxial [KD64, KL65]. Epoch [You65b]. Equalization [Ger69a, Luc65, Luc66]. Equalizer [AC69, LR67, Rud69, SS60]. Equalizers [BDV64, Joh64, Lyt68]. Equalizing [KSW69]. Equation [Liu69, Ben61c, Ben62d, Kad67b]. Equations [Ger69b, Hel65, Hyd64, Kle64, Llo67, NM64, Row62, SB64, San65a, You63a, You63b, Coc65, Mea62, San64d, San65b, San65c, San67c, SS68, SW69b, Wil68]. Equipment [BST65, CKS68, FGK^{+}64, GBW66, GP63, HF61, HD61, LMS^{+}69, SW68]. Equivalence [Buc69b, GK64b]. Equivalent [Hol67, Koe67, Mar66c]. Errata [Ano64j, Ano66l]. Erratum [Ano60g, Ano64k, Ano64l, Ano68k, Ano70]. Error [AS66, Aik67, Arm61, Buc69a, Ell63, Ell66, FA64, Hub67b, HM67, Mac61, MS66, Nel63, Pra69, Ros69, SS69a, SBF68, TW64, Wel66a, Wyn64, Wyn69a, Zad66b]. Error-Control [Wel66a]. Error-Controlled [SBF68]. Error-correcting [Mac61]. Error-Marking [Ell66]. Error-Rate [HM67]. Error-Voiding [Ell66]. Errors [Dos65, KR62, Kin62a, Kin62b, Kin64, Mar61c, MS65, Mor62a, Ung66]. Esaki [GS61, Hin60]. ESS [AGGK64, BBC^{+}64, BWY69, CMS^{+}64, CDGH64, CE64, CHK64, DDH64, DRRS64, DNT64, FH64, FGK^{+}64, FGM64, GHS64, HTU64, HHS64, KKV64, Kru69, MOP64, PS69]. essentially [LP62]. Estimates [Des65, Ell63]. Estimation [BDEL66, Dal67]. Estimators [Ziv69]. Evaluating [Mou67]. Evaluation [AL60, DJS63, DGPS65, Gib60, GS64b, HH65, Hel66, HGG62, Kle67, PW63, RK63]. Ewald [Gra65b, Gra65c]. Exact [Rai68a]. Examples [Ben66b, Kad66]. Exchange [Aar62, EH65b]. Excitation [McK67]. Excited [Col65, MLW67]. Excluding [Rap67]. Exhibiting [Cha69a]. Existence [NM64, Coc65]. Expansion [Kad66]. expansions [Ric68]. Experiment [JJ60, MR60, Man63, ROB69c, TS63, CCKT63, Dic63]. Experimental [Dav62, DSW66, EH65a, GS69a, HS64, HM67, Jak61, Man69, MSV62, May65, Mou67, Oss64, Wit65]. Experiments [BGR63, DD68, Glo67b, Hig62, HS63, Weg60, BBM^{+}63]. Exploding [Lap13b]. Exponential [Bul63, Hel65, Des63, Tak62]. Expression [SS67]. extended [Ano70, Lee69b]. Extension [Lee69a]. Extensions [GS65a, Sim67, Sle64]. External [Guc62, SCM^{+}63].

Fabrication [Man63]. Fabry [Ano62a, CG66, GR63, KK62]. Facilities [BG61a]. Facility [BM64, Ehr64]. Factor [CT64, Leg60]. Factoring [Ber67b]. Factors [Dei60, Log67]. Fades [Lee67b, Rai66a]. Fading [Oss64]. Failures [Cha69b]. Fast [Pfa63, Gen68]. Fast-Decay [Pfa63]. Faults [CT67]. Features [ADH^{+}69, BI66]. Featuring [FHS67]. Feedback [AS66, Ano64j, Ben65a, Bro69, GC63, Hol68, Rut61, San63c, San63b, San64a, Sav66b, Sim67, Spa68, Tak63b, San64c, Wyn69b]. Ferreed [FLLR60, Fei64]. Ferrimagnetic [VCDK63]. Ferrite [GHS64]. Field [Aul65, HS64, Kur67, Lee67a]. Fields [Ber67b, Dov67, Kro67, Mar66d, Gor60]. Figure [KU61]. Filling [CT64]. Film [MM66b, Sin66, SSD69]. Films [BCD64, CUM68, Dov67]. Filter [Abe67, FS60, MB61, San67a, Sch68a, Smi65, Sta67, Wyn69c]. Filtered [Ric69]. Filtering [Mar61b]. Filters [BR66, Bog69, EMT69, GK64a, Mar61a, Mar61d, Mar67a, Mos65, Mus63, Par68b, SSD69, Wan68]. Finite [Ben67c, Ber67b, Hub66, Lee60]. Finite-Width [Hub66]. First [Rai68b, Rio61]. first-served [Rio61]. Fits [Sch62b]. Fixed [San65a]. Fixed-Point [San65a]. flexibility [Lev61]. Flexural [Pee63]. Floating [KS67, San67a]. Floating-Point-Roundoff [San67a]. Flow [Bug65, McC62]. Flowing [Ber64a]. Fluctuating [Rap67]. Flying [Coo61]. FM [ABDS65, AMS65, AS65, BGLS68, BT63, Ben65a, BS63, Cro66, Cro67, Dar64, Enl63, GC63, HH61, Lio66, MS66, PW68, Ric69, Rut61, Rut68b, Rut68a, Sun61, Sun64b]. Focusers [Mar64a]. Focusing [Hei68, Mil64a]. Fokker [Liu69]. Folded [GT65]. Foreword [Ano62h, Ano69k]. Form [Koe67]. Format [BKS66]. Formation [AT60]. Formula [San67b]. Forward [Ott62, Wel66a]. Forward-Acting [Wel66a]. Foundation [Wei66]. Four [Log61, Log67, Rap64, Rap65, Var66]. Four-Phase [Rap64, Rap65]. Four-Point [Log61]. Four-Wire [Var66]. Fourier [Gen68, LP61, LP62, SP61, Sle64]. Frame [BMP67, GP64, MP69]. Frame-Repeated [MP69]. Framers [DDH64]. Free [Gar60, Mag68]. Frequencies [CG64, Ell62, HDSD64, Tho63, Til69]. Frequency [BR63a, Bog69, BH62b, Bri66, Bur65, CDW63, Col65, Cur60, Dra67a, Dra67b, DP67, Eva69, FHS67, GN67, Guc62, Hin60, KD64, Kes68, LS69, Mar61c, Mar61e, MRS69, Pra67, SS69a, Sch60, Sha63, SSD69, Woj60, San64c]. frequency-domain [San64c]. Frequency-Shift [BR63a, BH62b]. Frequency-Synthesizing [Woj60]. Fronts [DH63]. FSK [Sch68b]. Function [Ben61b, BHT64, Bun69, Lee69a, Mar69d, Pil68, Pro67a, Spa68, Zad66a, Ano70, Lee69b]. Functional [CG61, Luc63, San65a, San64d, San65c]. Functions [BR63a, BHT64, Dar63a, FS60, Gra65a, JM62, Rad62, Sch62a, Sle64, Hsu62, LP61, LP62, SP61, SS65]. Fundamental [Bus60, Mor62b]. Further [Mar63, MM61, Mor62a, Sle60]. Fusion [Jul67].

G [Ano64l]. G. [Ano61g]. GaAs [Kah64]. Gain [CS65, HW65, Rap68, SD64, Zuc68]. Gains [Rap67]. GaP [Ger66, Pfa64]. Garnets [GWES64]. Gas [Ano64k, Bec64b, Bec64a, Ber64a, Col65, Dai68, FP69, Glo67a, GL64, MM64, Mar65a, Mar66c, Mar66d]. Gaseous [Mar66a]. Gases [Ber64b, Ber64a, Bug65]. Gate [GS61, KS67]. Gauge [Fre69]. Gaussian [AFMP69, Bun69, Chu66, Dal67, Hal69, Kad64, Kad65b, Kad65a, Kad66, Kad67a, MM61, O’N66a, Pil69, Rai66b, Rap65, Ros69, San66b, Sle62, Sle68, Wyn64, Wyn66b, Wyn68, Wyn69b, Wyn69a]. GaussianNoise [Rai68a]. gc [Gig65, Lee66, UCE^{+}63, Con66]. Gc/s [Con66]. General [Arm61, Ben60a, Dra67a, Gor66, GS64a, Log67, LR67, Rud69, SH63, Suy65, Ben61a, Tak63a]. General-Purpose [GS64a, LR67, Rud69]. Generalization [Gra65b, Gra65c]. Generalized [BK62, Kad67a, Mil65, Shn67, Sle64]. Generated [Jul60]. Generating [Bra69]. Generation [LS69, Man69, Tho65]. Generator [AL68]. Generators [DP68]. Geographically [GK66]. Geometric [KH66]. Geometrical [Aul65, Chu66, Lie60]. Geostationary [LM69b]. Germanium [Cut61, TK60, Tru60]. GHz [Cop67, ST69, Til69, WY68, Wil69]. Gigacycle [Sch63]. Glass [GS69b, HP60]. Glass-Sealed [HP60]. Gold [BF60, GR67, KD64]. Gold-Doped [BF60]. Gold-Epitaxial [KD64]. governed [San65c]. Governing [Llo67]. Grade [Ott62]. Graded [MS60]. Gradient [LZ64, Mil64a]. Gradients [Ber64b]. gradings [Ben67a]. Granular [Goo69b]. Gratings [Kog69]. Gravitational [PWY63]. Gravitationally [FRY63, Yu65]. Gravity [Glo67a, LZ64]. Gravity-Gradient [LZ64]. Green [Pfa64]. Ground [CHLS63, GPW63, GT65, Liu68b, Pop68, UCE^{+}63]. Ground-Motion [Liu68b]. Grounded [BR63b]. Grounded-Base [BR63b]. Group [Ben61b, Buc66, GAPB63, Lin64, Sch60, Sle60, Sle68]. Groups [Woo64, Ben64b, Des63, Tak62]. Growth [Ber65b]. Guanidinium [GK62]. Guidance [Eck60, EMT63, MM69, Mil64b]. Guide [Ber64b, Ber65b, Dai68, Kai68, Mar66a]. Guided [HGW^{+}67]. GuideI [GS68]. Guides [Glo67c, Mar69a]. Guiding [Kai68]. Guilding [Gor66]. Gyro [LZ64].

hacked [Lap13b]. Half [GS68]. Half-Mile [GS68]. Hall [Kla61]. Handling [Gri64]. Hands [Gar60]. Hands-Free [Gar60]. Hard [Hal69]. Hard-Limited [Hal69]. Harmonic [DP68, Goe69, LS69]. Harper [Buc69b]. Haul [Dav62, FJJK66]. Having [Kla60, Llo67]. Healing [Bri68]. Heat [Hei67, HGG62]. Heating [Mor64a, MM66b]. Helix [Ste65, Ung61a, Ung61b, Ung61c, Ung61d, WY68, You65a]. Heterodyne [DD68, DG67, EJR68, Lar69, Pra68]. Heuristic [Ben62b]. Hexahydrate [GK62]. High [Abe67, Ano64a, Bos62, CHLS63, Coo61, Dra67a, Dra67b, Ell62, FHS67, GWB67, Hin60, KD64, Kla60, Kur67, Kur66, MP67, MLW67, SBF68, SSD69, Tab67]. High-Capacity [Ano64a, Kur66, Tab67]. High-Efficiency [Dra67b]. High-Energy [GWB67]. High-Field [Kur67]. High-Frequency [FHS67, Hin60, KD64]. High-Power [CHLS63, SSD69]. High-Precision [Ell62]. High-Quality [Abe67]. High-Resolution [Coo61]. High-Speed [Kur66]. Highly [Ger69a, MM66b]. Hilbert [Bog69]. holding [Tak63a]. Hollow [MS64]. Holmdel [Jak63, Ste65]. Hologram [EMR66, EJR68, Kog69, Lar69]. Holograms [Bur66, BE69]. Holographic [Kog65a, LPSL66]. Holography [JJ68]. Hop [CT69, ROB69c]. Horn [CHH61, DBKW63, GT65, HLT63, HW65]. Horn-Reflector [CHH61, DBKW63, HLT63, HW65]. Horns [CS65]. Human [Dei60]. Hybrid [CF68, FC68, Mar61a, SGB69, Fal69]. Hydrodynamic [Coy65]. Hydrogen [Ber65a]. Hz [GHS69].

I. [Ano64j, Ano66l]. Ideal [Ano68k, Goe66, LS68]. Identification [Mar68, Gop69]. Idle [McD66]. IF [BR63b, BG69, FJL68]. IGFET [Kah67]. II [BHT64, DGPS65, Fla62, FC68, Gra65c, Kin62b, LP61, Ric62b, SK61b, San64e, Wan61]. III [Hyd64, Kin64, LP62, MOP64]. Illinois [KH66]. Illumination [Enl67, Enl69, LPSL66]. Image [BCDY68, Enl67, GC68, HH65, Kog65a, LPSL66, Sle67]. Images [Hei66]. Imaging [Cha69a, CLW69, Kog65b]. Imbedding [KML69, San62a]. IMPact [MM66a]. IMPATT [Blu69, GS66, LS69, MM66a]. Impedance [BGK60, Lee69a, Mea68, Zad66a]. Impedances [San60]. Impendance [Pro67a]. Imperfect [Glo67c, Lio66]. Imperfections [KU62, Mar69c, RW62, Row64a, Row64b]. Importance [GWES64]. Improved [FHS67, KM61, MM69, Wan68, Wyn64]. Impulse [Fen69, Rap65]. Impulsive [Eng65, Wei65]. Impurity [AT60, Tru60]. Inch [KM61]. Incompletely [McC62]. Increase [MP69]. Increased [Whi66]. Increasing [Ski66, Lev61]. Independent [Lee67a, Mar67c]. Index [Ben65a]. Inductors [Leg60]. Inequalities [Ben65c]. Inferences [Mil63]. Infinite [FL64]. Influence [BCDY68, SK61a, SK61b]. Information [Ber67a, FC68, Pie66, Tay68b, Wol69, GM66]. Inherent [BH62b]. Inhomogeneities [Mor64b]. Inhomogeneous [Kog65a]. Injection [Rut68a]. Injection-Locked-Oscillator [Rut68a]. Innage [Ric63]. Inner [Bhu60]. Inorganic [Lie60]. Input [Cha69b, O’N66a, San64f, Gop69, Tak62, Tak63a]. input-output [Gop69]. inputs [Des63]. Insertion [Lee66]. Inspections [Rob62]. instantaneous [Lan60]. Insulated [Lep68]. Insulator [NG67]. Integral [Hsu62, Kad67b, NM64, SB64, Ben61c, Coc65, San65b]. integrals [Ric68]. Integrated [CUM68, Gia60, GS69b, Lep68, Mar69b, Mil69a, Sch69a, SGB69, Sch69b]. Integration [Iwe69, San67b, SS68]. integro [Ben62d]. integro-differential [Ben62d]. Intense [FG69, Hog69]. Interaction [Mor65]. Interactions [GWES64]. Interchannel [BT63, Enl63, PE69]. Interface [Bhu60, NM67, NG67]. Interference [Cur60, Cur62, Cur63, Eck60, Enl63, Hub67b, Luc63, LM69b, Pra69, PE69, Ros69, RT69]. Interferometer [FL61, Li63, RB64]. Interlaced [Bro67b]. Intermodal [Bat69]. Intermodulation [Cro66, Cro67, Sun64b]. Internal [Kog65b]. International [Ano66a]. Interpolation [MS62, SH67]. Interpretation [KH66]. Interpreting [CT67]. Interstitial [CCJ60]. Intersymbol [Hub67b, Luc63]. Intertoll [NEH68]. Intervals [Ben67c, Rai65, Rai67b, Ric62a, Ric63]. Intrasublattice [GWES64]. Introduction [Mil69a]. invariant [Gop69]. Inversion [GN67]. Inversion-Layer [GN67]. Investigating [BKRS68, Bra65b]. Ion [GWES64, GL64]. Ionic [Lie60]. Ionizing [Ros62a]. Iron [GWES64]. Irregularly [Ber65b]. Isolator [CT64]. Issue [Ano60a, Ano60b, Ano60c, Ano60d, Ano60e, Ano60f, Ano61b, Ano61c, Ano61d, Ano61e, Ano61a, Ano61f, Ano62b, Ano62c, Ano62d, Ano62e, Ano62f, Ano62g, Ano63a, Ano63b, Ano63c, Ano63d, Ano63e, Ano63f, Ano64b, Ano64c, Ano64d, Ano64e, Ano64f, Ano64g, Ano64h, Ano64i, Ano65a, Ano65b, Ano65c, Ano65d, Ano65e, Ano65f, Ano65g, Ano65h, Ano65i, Ano65j, Ano66b, Ano66c, Ano66d, Ano66e, Ano66f, Ano66g, Ano66h, Ano66i, Ano66j, Ano66k, Ano67a, Ano67b, Ano67c, Ano67d, Ano67e, Ano67f, Ano67g, Ano67h, Ano67i, Ano67j, Ano68a, Ano68b, Ano68c, Ano68d, Ano68e, Ano68f, Ano68g, Ano68h, Ano68i, Ano68j, Ano69a, Ano69b, Ano69c, Ano69d, Ano69e, Ano69f, Ano69g, Ano69h, Ano69i, Ano69j]. Iterative [CNN67]. IV [Sle64].

J [Ano60g]. J. [Ano60g]. Jitter [AG62, BKR63]. Joint [CF68, Smi65]. Journal [Ano60h, Ano60i, Ano60j, Ano60k, Ano60l, Ano60m]. Jr. [Ano61g]. Jump [Kyo69]. Junction [AT60, Bat69, LW60, Mar61a, MS60, SR67]. Junctions [Ger66, McK67, MM68].

Junctor [BBC^{+}64].

kernels [Kad67b]. Keyed

[Hub67b, HM67, Pra69]. Keying [BR63a]. Kilomegacycle [Sha63]. Klauder [Ano60g]. Klystron [Guc62]. Kmc [KM61, Ste65]. Known [RC65, MM61]. kw [SB61].

L [AEGH63, Ano69k, BBHH69, CDW63, GAPB63, HM63, Kes68, Suy65]. L-4 [Ano69k, BBHH69]. L-Type [AEGH63]. L8 [San64a]. L8-Stability [San64a]. Laboratories [And63, Jak61]. Laboratory [HR63]. Laminar [Bug65]. Laminated [Sla64]. Language [FGP66]. Large [Blu69, GM67, HW65, Koe67, Kro67, Mor64b]. Large-Signal [Blu69, GM67]. Laser [CH66, CUM68, Col65, McC65, McC69, NM64, SC65, SC67, Whi66, Coc65]. Lasers [GL64, MS64, SSD69]. last [Rio61]. last-come [Rio61]. Latching [GS65b, HP60, Zup60]. Lattice [Gra65a, Gra65b, Gra65c, Lee69a]. Lattice-Periodic [Gra65a]. Lattices [FGVH62]. Launch [EMT63]. Launching [UDKM63]. Law [Ben65a]. Laws [Kro67]. Layer [GN67, LW60]. Layers [Cut61, Irv62]. Laying [Ehr64]. Lead [Lep66, Lep68]. least [Lev61]. Lee [Ano70]. Length [Lub68, Pil68, Ros62a]. Lens [Bec64b, Bec64a, Ber64b, Ber64a, Ber65b, FP69, Glo67b, Glo67c, GS68, Hei68, Mar64a, Mar67b, MM64, Mar64b, Mar66c, Mil64a, Mil65, Mor64a, Mor64b, Ano64k]. Lens-Like [Mar64a, Mar67b, Mil65]. Lens-Waveguide [Mar64b]. Lenses [Dai68, Glo67a, Glo67c, Kog65b, Mar67a, Mar65a, Mar66d]. Lepselter [Ano68k]. Level [Bur65, Lee67b, MLW67, MP69, Mac61]. Levels [Bra65a]. Light [Ano64a, Ber64b, Ber65b, Bur66, CG64, CG66, DeL65, DA63, DG67, Kur66, LPSL66, Mar66a, Mar67a, Mar64b, Mar65b, Mil64b, Mil65, Nel64, Ski66, Tab67]. Light-Ray [Mar65b]. Light-Wave [Mil64b]. Lightning [BG61a]. Like [Enl67, Mar64a, Mar67b, Mil65]. limit [Mar67c]. Limitations [Mar66e, SG62, VZ67]. Limited [Cha66, Hal69, Hel65, Li63, Wyn66b, Lan60, LP62]. Limiter [VCDK63, Wyn69c]. Limiter-Phase [Wyn69c]. limiting [Lan60]. Line [BBC^{+}64, BAMS64, DSW66, JJ60, MR60, Man63, Rob62, Row62, Spa68, Sta65, You63a, Zad66a, Pro67a, Tak63a]. Linear [Ano64j, Bur60, Dal68, Dar63b, Eva69, HS69, Hyd64, Lio66, Mar68, Pie68, Rob67a, Rut68b, San62b, San63b, SBF68, Shn67, Gop69, Mea62, Pil69]. Linear-Real [Pie68]. Linearization [Hol67, Kyo69]. Lined [KU62, Ung62]. Lines [CC63, GW68b, Hsi67, Kin66, Row64a, Row64b, Sch69a]. Lining [Ung60]. Link [Koe67, Lub68]. Links [Ric63]. Liquidus [TK60]. little [Lap13a]. Load [GS64a, Hsi67]. Load-Loss [GS64a]. Loaded [GW68a, HDSD64]. Loads [Wei65]. Locating [CT67]. Location [Rai68c]. Lock [Nel63]. Locked [CDW63, Mos65, PW68, Rut68a]. Logic [CMS^{+}64, Gia60, TU62]. Long [BAMS64, Guc62, MS64]. Long-Term [Guc62]. Loop [Ano64j, Gol62, Gre69, Mag68, Mos65, Nel63, PW68, San63b, WY68]. Loop-Free [Mag68]. Loss [Des65, Ell62, FHS67, GS64a, Kai68, LK61, Lee66, Li65, Nas64, RB64, Sch69b, You63a]. Losses [DeL65, Kib69, Mar69d, Mor64b]. Lossless [DP67, Kin66, Pra67, Pro67a, Zad66a, San62a]. Lossy [DP68]. Low [Bay68, Bra67, BMP67, Bro67b, Dar64, Kai68, Mou67, OKS69, SS69a, Sla64, Ste61]. Low-Carbon [Sla64]. Low-Frequency [SS69a]. Low-Loss [Kai68]. Low-Power [Dar64]. Low-Resolution [Bro67b, Mou67]. Lower [US61]. Lower-Sideband [US61]. LSA [Cop67]. Lumped [ML68]. Luneberg [Mor64a, Mor64b].

M [Ano64l, Ano68k, San65d]. M.

[Ano64l, Ano68k]. Ma [Lap13b, Lap13a]. Machine [KH66]. Machines [CT67, Gia60, Mar68]. Machining [CUM68]. Made [Dai68]. Magnetic [Aul65, Bob67, Dov67, GWES64, Gia60, Gla67, GS65b, HP60, Kro67, Lee67a, Pee61, Thi69, Zup60, Gor60]. Magnetization [Pee61]. Magnetoelastic [Aul65]. Main [KSW69]. Maine [Cur63, Gig65]. Maintenance [BFH^{+}69, DNT64]. Malfunction [MMW^{+}63]. Manipulations [Dar63b]. Manufacture [Joh64, LP64]. Many [DeL65, Sle64]. Mapping [San65a]. Maps [Hol67]. Marcuse [Ano64k]. Margin [GS61]. Marine [Ste61]. Marking [Ell66]. Markov [Ben63a, Bar64a, Bun69]. Maser [CT64, DKS61, FL61, GS62, Gor64a, Kle64, Li63, Li65]. Masers [BG61b, HDSD64, SD64, TS63]. Master [DRRS64]. Matched [Sch68a]. Matching [Kog64]. Materials [GS65b, Ham67, Ste61]. Mathematical [Hae63, Wei66, Ben62b]. Mating [Pee61]. Matrices [Rao69]. Matrix [Dar63b, Gen68, ML68]. Maximally [Hel65]. Maximization [Hol66, Mor62b]. Maximum [GM67, Hol66]. Mb [DSW66, EH65a, May65, Wit65]. Mb/s [DSW66, EH65a, May65, Wit65]. mc [DKS61, Kib65a, LK61, MV61, SB61, US61, War60]. mean [BDEL66, Mac60]. Means [GS64b, Ros62a]. Measured [Dai68, You65a]. Measurement [AK65, Ben61b, Bra65a, HW65, KR62, Lee67a, Mau67, Ros62a, SK69]. Measurements [Bec64b, Bec64a, Kur62, Log61, Log67, RC65, Rus66, Sem65, ST69, WY68, Wil69]. Measuring [CL60, Ell62, Eva69, GHS69, HR63, IKM67, LK61, Lee66, RB64, Rai68c]. Mechanical [DGPS65, DBKW63]. Media [Gor66, Kog65a, Mil65]. Medium [Aik69, DeL65, Kai68, Mar67b, MP67]. Megabit [GHS64, Kib65b]. megacycle [KS63]. Melt [SK61a]. Membrane [Fla60, Fla62]. Memories [DGS68, Tay68b]. Memory [AC61, GHS64, KS67, Kah67, Ski66, SG67, Hsu62]. memoryless [Fal69]. Mental [JJ68]. Message [Ado63, AL60, CL60, Lew64]. Metal [NG67, SSD69]. Metal-Insulator-Silicon [NG67]. Metallic [Kla61, MS64, Sem65, Ung61a]. Methane [CH66]. Method [Arm61, BT63, Cha69b, Gra65b, Gra65c, Par68a, Sch62a, You65b]. Methods [Bla64, CT67, Gey64]. Mev [RGS63]. MHz [BG69, GHS69, SS69b]. Microns [CH68]. Microscope [Hei68]. Microstrip [Sch69a]. Microwave [AL68, Bra68a, CG64, CT69, FJJK66, HHJP68, JRS68, JDC65, KL61, Mor64a, Sch69a, SGB69, Sch69b, SW68, SD61, Ung66]. Microwaves [Hog69]. Middle [Fla62]. Middle-Ear [Fla62]. Mile [GS68, WY68]. Miller [Ano64k]. Millimeter [BG61b, HGW^{+}67, LS69, SGB69, Sta67]. Millimeter-Wave [HGW^{+}67, LS69]. Milliwatt [Sch63]. Miniature [Bos62, Bro67a, Kel64]. Miniaturized [Mos65]. Minimal [Rao69]. Minimally [RC61]. Minimize [McM69]. Minimum [Gil67, KU61, McC62, Rap67]. Minimum-State [McC62]. Mirror [GK64b, RB64]. Mirrors [GCE69, Li65]. Misaligned [Dal69, Mar67a]. Missile [Eck60]. Mixed [Ott62]. Mixers [Dra69]. Mixing [DG67]. Mixture [Bec64a]. Mobile [Cla68, Gil65, Gil69, Lee67b, Oss64, Ano70, Lee69b]. Mobile-Radio [Cla68]. Mode [Enl63, Glo67c, Li63, Mar61d, Mar67a, Mar69c, MD69, Mil63, NS65, SSD69, Sta67, Ung61a, Ung61c, Whi66]. Mode-Conversion [Mar61d]. Mode-Selection [SSD69]. Model [Bra69, Ell65, Fen69, Fra66, Kar66, Oss64, San69c, Til68, Wei66, You63b]. Modeling [GWB67]. Models [Ben67b, Fla60, Fla62, Hae63, Koe67].

Modes [FL61, Kog64, Kog65b, Li65, Mar64a, MM68, Ung61c]. Modification [Mau67]. Modulated [Bar64b, CG64, Mar61b, Mar61e, PE69, Ric69]. Modulation [AFMP69, BT63, Dav62, Dra67a, Dra67b, Dra69, Goo69a, Goo69b, Iwe69, Jac67, LS69, May62, MRS69, McD66, O’N66a, O’N66b, Pie66, Pro67b, Ric69, Sch68b, Shn67, SC65, SC67, Sun64a]. Modulator [AART68, GR63, HP68, Pil69]. Modulators [BH62b, DA63, DG67]. Molecular [Bur60, Hei66]. Monitoring [Coy63]. Monocularly [Jul67]. Monographs [Ano60h, Ano60i, Ano60j, Ano60k, Ano60l, Ano60m]. Monopulse [Bay68]. Monte [FGVH62]. Morris [KH66]. MOS [Goe66]. Motion [Liu68b, Sle67, Wei65]. Mounting [SGB^{+}63]. Mounts [NC65]. Moving [Wei65]. Multi [Ano64j, San63b]. Multi-Loop [Ano64j, San63b]. Multichannel [Cha66]. Multicolor [LPSL66]. Multimode [BG61b, Mil63, RW62]. Multimoding [Ble68]. Multipair [GPZ63]. Multipath [Aik67, BT63]. Multiple [Cha67, GW68b, GS69a, Hei67, Mar66a, Mac61]. multiple-level [Mac61]. Multiple-Speed [Cha67]. Multiplex [AEGH63, CDW63, GAPB63, HM63, Kes68, O’N69, Suy65]. Multiplexer [Wit65]. Multiplexer-Demultiplexer [Wit65]. Multiplication [Gen68]. Multipliers [Bur65, Dra67a, Dra67b, DP67, Pra67]. multiport [San62a]. Multiprocessing [Gra66]. Multistage [Ben64a, GS64a]. Music [Mat61]. Mutual [GK66]. Mutually [Bri66, Bri67].

n [Kah64, McK67, MM68, MS60, SR67, Ano61h, FS60, San61]. N-Path [FS60]. N-Port [Ano61h, San61]. N2 [BST65, BH65, GR67, LW65]. N3 [BI66, GR67, GBW66, HW66]. Narrow [Rai66b, Rob67a]. Narrow-BandGaussian [Rob67a]. Nature [Mil63]. Near [Ano68k, HS64, LS68, SR67]. Near-Field [HS64]. Near-Ideal [Ano68k, LS68]. NEASIM [GS64a]. Negative [BGK60, GC63, Hin60, Kur69, Mea68, Rut61]. Negative-Resistance [Hin60]. Net [FHS67, HGG62]. Network [AGN60, DDH64, Ell65, FH64, FS60, Ger69b, Mor62a, RT69, San62b, SW69a, Sca63, SK69, TW64, Wel66a, Woj60, Lap13a, San62a]. Network-Theoretic [SW69a]. Networks [Ano61h, Ben62c, Ben64a, Ben65b, Ben66b, Ben67c, Ben67b, DK65, DMS68, Eva69, Gil67, GS64a, Joe68, Kat65, LK61, Lee66, Mag68, Pau62, Pie69, Rut68b, San60, San61, SW69a, San69c, Spa69, Web62, Web64, Wol65, Ben62a, Ben63a, Ben63b, Ben64b, Ben66a, San69b, SW69b, Wil68]. Nickel [BM63]. No [AGGK64, BBC^{+}64, BWY69, CMS^{+}64, CDGH64, CE64, CHK64, DDH64, DRRS64, DNT64, FH64, FGK^{+}64, FGM64, GHS64, HTU64, HHS64, HTZ64, KKV64, Kru69, MOP64, PS69, UV64, Ric62a]. Nodal [Par68a]. Noise [AL60, Bra67, CL60, Cro66, Cro67, Dra67b, Dra68, Ell63, Eng65, Enl67, Fen69, GP60, Gil60, Goo69b, Gor64a, HS61, Hol66, Hub67a, Iwe69, KU61, Lew64, Lio66, Man69, Mar61b, Mar61c, Maz68, McD66, McM69, MP69, Mus63, Nas64, NM67, O’N66a, OKS69, Pra68, Pro67b, Rai67a, Rai67b, Rai68b, Rap67, Rap65, Rob67a, Ros69, SS69a, Sle62, Wyn69a, Kad65b, MM61, Ric68]. Noise-Like [Enl67]. Noisy [For69, Wyn68, Wyn69b]. Non [NC65, Rad62]. Non-Computable [Rad62]. Non-Orthogonal [NC65]. Nonbinary [Wol69]. Nonblocking [Ben65b]. Noncentral [Rob69a]. Noncylindrical [Ung61b]. Nondissipative [Dar63a]. Noninterlaced [Bro67b]. Nonlinear [DK65, Ger69b, Hol68, Kat65, Kyo69, Liu68a, Liu69, Mor62b, Rob67b, San63c, San64a, SB64, San64f, San65a, SW69a, San69c, Ben61c, Ben62d, BS64, Hsu62, San64d, San64c, San65b, San65c, SS68, San69b, SW69b, Wil68]. Nonlinearities [Ben65a, Rob68]. Nonlinearity [SR67]. Nonnumerical [Bro63, BHT64, Hyd64]. Nonprobabilistic [Kru69]. Nonuniform [Aul65]. Nonuniformly [KG66]. Norm [San65a]. Norm-Boundedness [San65a]. Normal [Nas67, Ung61c, Mar67c]. Note [Ano66l, Bar65, San64b, San66a]. number [GM66]. Numerical [Buc67, HH65, San67b, SS68, Tho63, San67c]. Nutt [Ano64l]. Nyquist [GS65a, Shn67].

Objective [Bra65a, Hei68, Lew64].

Objectives

[BI66, BH65, HM63, KKV64, SV69]. Objects [Enl67, Enl69]. Observables [AK65]. Observation [San65a]. Observations [Ung66]. Observed [Con66, KM61]. obtained [San62a]. occurring [Ric68]. Ocean [Ano64l, Bhu60, BNR64, Joh64]. Ocean-Block [Joh64]. Off [Bra65b, Bra68b, Bra69, Mar67b]. Off-Axis [Mar67b]. Office [GS64a, KH66]. Offices [WC69]. On-Off [Bra65b, Bra68b, Bra69]. One [Bar65, BC69, Sle62, Ben60a, Ben61a, Tak63a]. One-Bit [BC69]. One-Sided [Bar65, Sle62]. Open [BCD64, CEZ65, DGPS65]. Open-Ended [BCD64]. Operated [GHS69, GL64]. Operating [GM67, Rob67a]. Operation [Cop67, FHS67, SCM^{+}63, Wyn69c]. Optic [DA63, Kur66]. Optical [BG61b, CG65a, Dal68, DD68, DG67, GS62, GW68b, GS68, GS69a, GS69b, Gor64b, GK64b, Gor64a, Hil63, Hog63, Kai68, Kog64, Kog65b, MS64, Mar69a, Pra68, RB64, Rig65, Rob67b, Rus66, SG67, Tho63]. Optics [Aul65, Gor66, Mar67b, Mar69b, Mil69a]. Optimal [Ben64a, Ben67c, Ben66a, Ben67a]. Optimality [LS66]. Optimization [Ano66l, CF68, Ger69b, MRS69, San66a, Smi65].

Optimum [Kad64, Kad65b, Kad65a, Kad67a, Shn67, Yu65, Pil69]. Orbit [Bla64, CBHR63, Lie60]. Orbital [Tho65]. Orbits [Gey64, RR62]. Order [CM69, Dra67b, Ric69, Wyn69c, SS65]. Orders [Ano62a, KK62, Tak63a]. ordinary [Mea62, San67c]. Organic [BKRS68, CK68, Kar66, Ste61]. Organization [GKLL63, HTU64, HHS64, KKV64, SV69]. Orientation [Hil63]. Oriented [FRY63, Yu65]. Orthoferrites [Bob67]. Orthogonal [Cha66, Kad66, NC65]. Oscillating [RB64]. Oscillation [Mar66e]. Oscillations [Ber65b, EMT69]. Oscillator [Byr62, Cop67, JDC65, Kur69, LS69, Rut68a, Sch63]. Oscillators [Bar62, Blu69, Bri66, Bri67, GK66, Gor64a]. Other [Ham67]. Outage [Ric63]. outlaws [Lap13b]. Output [GM67, McC65, McC69, Rob69b, Sch63, Gop69]. Output-Coupling [McC65, McC69]. Over-All [FBL62]. Overdriven [GM67]. Overflow [EMT69, Lin64, Des63]. overlapping [You65b]. Overload [And69, Pro67b]. Oxidation [AT60]. Oxide [Dew60]. Oxidized [NM67].

P [Ano68k, McK67, MM68, MS60, SR67]. Package [SR63]. Packaged [BST65, CT69]. packings [Wyn67]. Packs [CE64]. Pair [AK65]. Paired [BG61a]. Papers [Ano60h, Ano60i, Ano60j, Ano60k, Ano60l, Ano60m]. Paramagnetic [Weg60]. Parameter [KG66, Ziv69]. Parameters [Gar60, LK61]. Parametric [Ham64, Pop68, UCE^{+}63]. Part [CF68, CEZ65, DGPS65, Fla62, FC68]. Participation [Jak61, Jak63]. Particular [McD66]. Particularly [Kel64]. Passage [Rai68b, Weg60]. Passive [Kur62, PWY63]. Path [FS60, Lub68]. Paths [DD68]. Pattern [Hig62, Wan61]. Patterns [Bay68, Bra65b, Bra68b, Bra69, Jul60]. Payout [Gre64]. PCM [AG62, Aar62, Bro69, Dos65, EH65a, Goo69a, GP64, Man69, MSV62, May65, SG62]. Perceivable [Jul67, JJ68]. Perception [Jul60]. Percolation [FGVH62]. Percussive [Coy63]. Performance [CT64, Ell66, Gib60, GKLL63, HM67, McD66, NEH68, Pra68, Rob68, Ros69, SG62, War60, Wel66a]. period [Rio61]. Periodic [Dal67, Gra65a, Hsi67, LU64, Mar65a, Mil68, Rut68b, San64f, BDEL66, Ben62d, BS64, Mea62]. Periodically [Dar63a, Mil69b]. Periods [Ric62b, Ric62a]. Peripheral [DFHT69]. Permanent [Mil64c]. Permeability [Ham67]. Permutation [Ben64b, Joe68, Mac64, Sch68b]. Perot [Ano62a, CG66, GR63, KK62]. Personal [KS63, MV61]. Perturbation [Gey64, McC65]. Phase [AG62, Bar62, Byr62, CDW63, DH63, Dra67a, Dra67b, Dra69, Ell62, FG66, Gol62, HH65, Hei66, Hub67b, HM67, LK61, Lee66, Mos65, Nag67, Nel63, PW68, Pra69, Rai66b, Rai67a, Rai68c, Rap64, Rap65, Rus66, Sun61, Tho63, Wyn69c]. Phase-Controlled [Bar62, Byr62, Gol62]. Phase-Lock [Nel63]. Phase-Locked [CDW63, Mos65, PW68]. Phase-Shift [Pra69]. Phase-Shift-Keyed [HM67]. Phase-Shifted-Keyed [Hub67b]. Phased [AG68, GW68a, VZ67, WG68]. phone [Lap13b]. Phosphors [Pfa63]. Photoconductivity [Van60]. Photodiodes [Sch66]. Photographs [Sle67]. Phreaking [Lap13a]. Physical [Gre69, Mar66e, San65c]. Physiological [Fla62]. Picture [BMP67, Mou69]. Picture-Element [Mou69]. Pictures [Bro67b, MP69]. Piezoelectric [Lew61, Llo67]. Plan [ADH^{+}69, BFH^{+}69, DNT64, FH64]. Planar [KL65, McK67]. Planck [Liu69]. Planning [SCM^{+}63]. Plant [Aar62, EH65b, Gre69]. Plastic [Ham67]. Plated [DGS68]. Plates [HGG62, Wei65]. Plugs [Wu69]. Plus [Rai67a, Rai67b, Rai68b]. PM [BT63, Cro66, Lun69, PR69, Sun61]. Pocket [KS63]. Point [Aik69, Log61, Log67, San60, San65a, San67a, Sha63, Ric68]. Point-Contact [Sha63]. Pointing [GKLL63, GP63, LM69b]. Poisson [Des63, Tak63a]. Polarization [Kur66]. Pole [Rob62, Sta67]. Polyethylene [Bur60, Ham67, SK61a, SK61b]. Polynomial [Bro63]. Polynomials [Ber67b, Bro63]. Polyphase [Wyn66a]. Pontryagin [Hol66]. Popov [San65d]. Populations [MLW67]. Port [Ano61h, Eva69, LK61, Lee66, San61]. Portraying [JJ68]. Position [Dal68]. Positioning [LNI63]. Possible [Kru69, San69a]. Postdetector [Rob68]. Potential [Dew60, MS60]. Power [Ben65a, BM64, BM69, BFT^{+}63, Bro63, BHT64, CHLS63, Cro67, Dar63b, Dar64, GHS61, GM67, Hal69, JM68, Mar69d, Mor62b, Nas67, Oss64, SBF68, SSD69, VCDK63, You63b, Mar67c]. Power-Law [Ben65a]. Powering [BAMS64]. Practical [SG62]. Precession [Yu63]. Precipitation [CH68]. Precise [HW65, WY68]. Precision [AHMS63, Ell62]. Precoding [Cha67]. Predicting [BT63]. Prediction [CBHR63, Hel65, TC64]. Predictive [O’N66b]. Presence [Eng65, SS69a, Sun61, Sch65]. Presentation [JM62]. Pressure [Bhu60, Sch62b]. Price [Ano60g]. Primary [CDW63]. Principle [Rai66b, Rai68c, Hol66]. Principles [Gla67, Hin60]. Prisms [Ano64a, Tab67]. Probabilities [Ben60b, SS69a, Zad66b]. Probability [AG62, AS66, Aik67, MS65, Wyn69a, Ric68]. Probe [Log61, Log67]. Probing [CG65a]. Problem [And69, Ano66l, BR66, Lin65, San64b, San66a, Sav66a, Sle62, Ben67a]. Problems [CNN67, Mel68, Ben62b, Ben66a]. Procedures [Mag68]. Process [Coy63, Dal67, Fra66, BDEL66, MOP64]. Processes [Bun69, FGVH62, Hal69, Kad66, Rai65, Ben60a, Ben63a, Des63, Tak61].

Processing [ADH^{+}69, CDGH64]. Processor [HTU64]. Produced [Enl63, Ric69]. Program [AGGK64, GS64a, HHS64, DJS63]. Programming [Ben66a, FGP66]. Programs [BHR69, CBHR63]. Progress [Hig65]. Progression [Nag67]. progressive [Ben67a]. Project [CHH61, DKS61, De 61, Hit63, Jak61, Jak63, Kib61, KNG61, Ohm61, Rut61, RJ61, SB61, US61, War61]. Projection [Kog65a]. Prolate [LP61, LP62, SP61, Sle64, SS65]. Propagated [Hog63]. Propagating [Kur67, Mar66d]. Propagation [Aik69, Aul65, BT63, Chu66, CH68, LU64, Mar66b, Mar64b, Mar65b, Mil65]. Properties [BF60, Bar64b, Ben65b, Bob67, BCDY68, Byr62, Dra67a, Kib69, Lew61, Mar65a, Mil64a, Nag68, Nas67, NG67, Pfa63, SB64, San64e, San67b, SW69a, San69c, VZ67, Zup60, Ben62a, Mac65, San63a, SW69b]. Proposals [KS66]. Proposed [Ben67b]. Propped [Pee63]. Protection [GL61, GN65]. Proton [GWB67]. Provided [RR62]. Proving [Wan61]. PSK [Hub66, Hub67a, Ros69]. Psychological [Mat61]. Published [Ano60h, Ano60i, Ano60j, Ano60k, Ano60l, Ano60m]. Pull [Pee61]. Pulse [Bar64a, Dav62, FGM64, GS68, Kib65b, Mar61b, Mar61e, May62, McD66, O’N66b, Pro67b, SD64, Shn67, Spa69, Sun61, Wit65, Zad66b]. Pulse-Modulated [Mar61b, Mar61e]. Pulse-Shaping [Spa69]. Pulses [Mar61c]. Pumped [Dra68]. Pumping [GL64]. Purpose [GS64a, LR67, Rud69]. Pushbutton [Dei60, Sch60]. PvN [MM66a].

Quadratic [FP69, MS65]. Quality [Abe67, Leg60, MS62]. Quantitative [NM67]. Quantization [Buc69a]. Quantized [AS66, Lim69, Sim67]. Quantizer [LM69a, Rob69b]. Quantizing [Goo69b, Iwe69, Maz68, O’N66a, O’N66b]. Quantum [Pfa64]. Quartz [War60]. Quasi [Kro67]. Quasi-Stationary [Kro67]. Queue [Tak63b]. Queues [CM69, Ben61a]. Queuing [Kru69].

R [Ano60g, Ano64l]. Radar [De 61, KR62, Kla60, Rai66a, Rap67, US61]. Radars [KPDA60, Ano60g]. Radiant [HGG62]. Radiating [Nag68]. Radiation [BGR63, BBM^{+}63, CH66, DH63, Mar69d, MW67, PBBS63, Ros62a, SK61a, SK61b]. Radii [Lie60]. Radio [BSW62, BT63, Cla68, CT69, CCJ60, Cur60, GH60, GHS61, GL61, Gil65, Gil69, HW61, HSW63, HHJP68, HF61, HH61, HD61, KR62, KL61, Lee67b, LM69b, MV61, Oss64, Rai68c, RT69, ROB69c, SGB69, SW68, SG61, SD61, Ano70, Lee69b]. Radio-Relay [LM69b]. Radiometer [OS63]. Radiotelephone [FHS67]. Radome [DBKW63]. Rain [Fre69, Gig65, HS61, Hog69, Wil69]. Rainfall [FG69, SK69, ST69]. Random [Ben65b, Dal67, Eck60, Enl63, Fra66, Liu69, Mor64b, Nas67, Rap67, RW62, Ste66, Ung66, Wyn67, Zuc68, Ano70, BDEL66, Lee69b, Mar67c, Sch65]. Randomly [Aik69, Rap67]. Range [Kla60, Sha63]. Rapid [SK69]. Rate [Bun69, FC68, Hel65, Hub67b, HM67, Kle64, MRS69, Pie66, Pra69, SK69]. Rate-Distortion [Bun69]. Rates [Ell63, MS66, SS69b]. Ratio [Hol66, Mar61b]. Rational [BHT64, Spa68]. Rational-Function [BHT64]. Ray [Ber65b, CLW69, Coo61, GK62, Mar66b, Mar65b, Mar66e, Pfa63]. Rayleigh [Rai65]. Rays [Mar64b]. RC [Ano61h, Mos65, San60, Sin66, Spa68]. Reactance [Rao69]. Read [Kah67]. Read-out [Kah67]. Reading [Ell62]. Real [BS63, Pie68]. Realizability [Pro67a, Zad66a, San62a]. Realization [FS60]. Realizations [San67a].

Rearrangeable [Ben62c, Ben64a, Ben64b]. Receiver [AART68, BSW62, BGLS68, FJL68, GPW63, HP68, HH61, JRS68, KS63, Kib61, OS63, Rap67, Rut68a, Shn67, SG61, UCE^{+}63]. Receivers [Kad67a, Pra68, Rap68, Sch68a]. Receiving [Ohm61, OKS69, Smi65]. Reception [Cla68, Gil65, Gil69, Kad64, Kad65b, Kad65a, Sch65]. reciprocal [San62a]. Recirculating [SS69b]. Recognition [Hig62, Wan61]. Recommendation [Ano66a]. Reconstruction [LPSL66]. Reconstructive [AG62]. Recovery [Sal67, San64b, Ton66, Lan60]. Rectangular [GW68a, Goe69, Kib69, Log67, Mar69b, WG68]. recurrent [Tak62]. recursive [Lev61]. Red [Pfa64]. Redirected [DeL65]. Redirectors [Mar66b, Mar67a]. Redistribution [AT60]. Reduced [BE69]. Reducing [Ber65a]. Reduction [Ben65a, Mou67]. Redundancy [Luc68]. Redundant [RC61]. Reed [Bos62, ET69]. Reeds [Pee61]. Reference [Kes68]. Refinement [Bla64]. Reflecting [Enl67, Enl69, HGG62]. Reflections [Bul63, Oss64]. Reflector [CHH61, DBKW63, GT65, HLT63, HW65]. Reflex [Guc62]. Refocused [DeL65]. Refocusers [Mar67a]. Refracting [Aik69]. regarding [Ben62b]. Regenerative [AS66, HM67, HGW^{+}67, Kin62a, Kin62b, Kin64, McC68, Sim67]. Regenerator [Zad66b]. Regenerators [BKR63]. Region [GS66, Kib69]. Regular [LS66]. Regulating [GLR69]. Rejection [Wan68]. Related [Gar60, LS69, Mil64a, San65d]. Relating [Luc63, You63b]. Relation [Cut61, Dar63a]. Relations [Dar63b, DP67, Fla62, GK64b, Hyd64, MM69, Ric62c]. Relay [Cur60, GH60, HW61, HSW63, HHJP68, KL61, LM69b]. Relays [Bro67a, GS65b, HP60, Kel64, Sla64]. Reliability [MP67, Ros62b, Sav66b].

Reliable [Hel65, RC61, Tay68a, Tay68b]. remarks [Ben62b]. Remote [AC69, JJ60, MR60]. Remote-Control [AC69]. Removal [Luc68, SS69a]. Reorganizing [Sca63]. Repeat [Wyn69a]. Repeat-Request [Wyn69a]. Repeated [MP69]. Repeater [AS66, BGK60, CKT63, Gri64, HGW^{+}67, May62, DHWM63]. Repeatered [CC63, DSW66]. Repeaters [AG62, BR63b, BDV64, CF68, GLR69, HM67, Joh64, KSW69, Kin62a, Kin62b, Kin64, Sim67]. Repetition [BMP67]. Replenishment [BMP67, Mou69]. Reported [Mor62a]. Representation [Chu66, Nar67, Pop68, Hsu62]. Representing [Koe67, Ben63a]. Request [Wyn69a]. Requirements [BE69]. Research [CCKT63]. Resistance [FL64, Hin60, Kur69, Lee69a, Ste61, SS67]. Resistant [Eck60]. Resistive [Dra68, Dra69, Wil68]. Resistivity [Irv62, Log61, Log67]. Resistor [GS61, Kat65]. Resolution [Bra67, BMP67, Bro67b, BE69, Cha69a, Coo61, Kla60, Mou67]. resolving [You65b]. Resonance [Weg60]. Resonances [Sch69b, Wu69]. Resonant [FL61]. Resonator [Ano62a, BG61b, BK62, CG66, KK62, Llo67, McC65, McC69, Rig65]. Resonators [Bog69, Gor64b, GK64b, Kog65b, Li65]. Response [AFMP69, Bri67, Fri68, GN67, San64f, San69b]. responses [BS64]. Restoration [Sle67, Zad66b]. Restricted [McC62]. Result [San65a]. Results [BGR63, CK68, HBK63, HS63, Jak61, O’N66a, Rai68a, MM61, San65c, San65d]. Reswitching [Pau62]. Retransmission [FA64]. Reverse [SR67]. RF [Col65, Mar61c]. RF-Excited [Col65]. Rigid [Joh64]. Riley [Ano64l]. Ring [Ble68, Rig65]. RLC [DK65]. Rooting [Mag68]. Rough [DH63, RC65]. Round [MD69, Ung60]. Roundoff [San67a].

Routing

[Ben67c, Web62, Web64, Ben66a, Ben67a]. Rule [Rab68].

S [Ano60g, Ano64k, Ano68k, Con66, DSW66, EH65a, May65, Wit65]. saddle [Ric68]. Salesman [Lin65]. Sampled [GK64a, Gol66, KG66]. Sampled-Data [GK64a, Gol66]. Sampling [Bog69, Dal67, BDEL66]. Sandberg [Ano64j, Ano66l]. Satellite [BLOR63, CG65b, Cur62, Cur63, De 61, FRY63, GWB67, Gey64, Gig65, GT65, HOW63, HS63, LZ64, Lub68, MMW^{+}63, Nel63, SH63, TS63, Til68, US61, UCE^{+}63, Ung66, UDKM63, VCDK63, Yu63, Yu65, Ric62a]. Satellite-Tracking [US61]. Satellites [Gla67, Hil63, LM69b, PWY63, PW63, RR62, Ros62b, RP68, TC64, Tho65]. Saturation [SD64]. Sawtooth [Byr62, Gol62]. Scale [Bro69]. Scaling [Kro67]. Scanner [FGM64]. Scanners [EJR68]. Scanning [Lar69]. Scatter [Sem65]. Scattering [CG66, Dra67a, DP67, GCE69, Mor64b, RC65]. Scheduling [Kru69, Rob62]. scheme [Fal69]. Schindler [Ano61g]. Schottky [Ano68k, AART68, Kah64, KL65, LS68, Sch66]. Scramblers [Sav67]. SD [BDV64, EFK^{+}64, HVW64]. Sealed [HP60, Kel64]. Seals [RE62, Sch62b]. Second [Rai68b, Ric69, Wyn69c]. Selection [Li63, Li65, SSD69]. Selector [Bos62]. Self [AG62, Bri68, CKT63, FA64, Mur68, Sav67, Sca63, SP66]. Self-Adaptive [SP66]. Self-Analogues [Mur68]. Self-Contained [FA64]. Self-Healing [Bri68]. Self-Reorganizing [Sca63]. Self-Steering [CKT63]. Self-Synchronizing [Sav67]. Self-Timed [AG62]. Semiconductor [GF60, Lin62, Log61, MW67]. Semiconductors [Kur67, PDS64, Van60]. Semipermanent [Kah67, SG67]. Sensing [BCDY68, GC68]. Sensitivity [Bos62, San66b]. Separated [GK66].

Separation [JJ60, MR60, Whi66]. Sequence [Fra68, Mar64a]. Sequence-State [Fra68]. Sequential [Gia60, Mar68, McC62, Sav66a, SS69b, UW69]. Series [Bro63, BHT64, Bro67a, Nar67]. Served [CM69, Rio61]. Server [Ric62b, Tak63b, Ben60a, Ben61a, Ric62c]. Service [ADH^{+}69, BHR69, BBC^{+}64, Ott62, RR62, Des63, Ric62a, Rio61, Tak62, Tak63a]. Servo [LNI63]. Servomechanisms [Hel65, Ben61c]. Set [CL60, Ell62, GHS69, HR63, LK61, Lee66, Wel66b, GM66]. Sets [Dei60]. Several [Bro63, BHT64]. Shape [Lew61, Rai68c]. Shaping [Spa69]. Sharpening [SD64]. Sheathed [WG68]. Sheet [GHS64, Log67]. Shields [Kro67]. Shift [BR63a, BH62b, HP68, HM67, Pra69]. Shifted [Hub67b]. Ship [BAMS64]. Short [AC61, CT69, Dav62, FJJK66, ROB69c]. Short-Haul [Dav62]. Short-Hop [CT69]. Short-Term [AC61]. Shorted [Cha69b]. Shot [Dra68, GP60]. Shrink [Sch62b]. Shuttling [GS68]. SI [Ano66a, MM66a, NG67]. Side [Hyd64]. Sideband [Bar64b, GM67, US61]. Sided [Bar65, Sle62]. Sidelobes [Bay68]. Signal [BKS66, Blu69, Bra67, Dav64, FGM64, GM67, Hol66, Koe67, Lee67b, Mar61b, McD66, San63c, San64b, Smi65, Ziv69, Lan60]. Signal-Noise [Hol66]. Signal-to-Noise [Mar61b, McD66]. Signaling [Aik67, BD60, MV61]. Signalling [KS63]. Signals [AFMP69, Bar64b, Cha66, Dar64, GP64, HM67, Kad64, Kad67a, Kla60, Lim69, May62, McD66, O’N66a, O’N66b, Rai66b, Rap67, Rob67a, Rut68b, SK65, San64e, San64f, SH67, Kad65b, Kad65a, LP62, MM61, San63a, Sch65, You65b]. Silicon [Ano68k, AT60, BF60, BCDY68, CBL^{+}67, CL69, GR67, Goe66, GN67, Irv62, JDC65, KD64, KL65, LS68, NM67, NG67, Sla64, TK60, Tru60]. Simple [Ben61b, Kai68, Liu68a, SK65, Sav67, Tak62].

Simplex [LS66]. Simplifying [Sch62a]. Simulating [Cha69b]. Simulation [CK68, Gib60, GS69a, Gol66, GS64a, Liu68b, O’N66a, Rap64, Web64]. Simulator [BKRS68]. Simultaneous [AK65, Kad67b]. Simultaneously [Kad66]. Sine [Rai67a, Rai67b, Rai68b]. Single [Bar64b, Col65, Iwe69, McC69, Ric62c, Ric62b, San64a, Sch63, SSD69, Tak63b, San64c]. Single-Frequency [Col65, SSD69]. Single-Integration [Iwe69]. Single-Server [Tak63b]. Single-Sideband [Bar64b]. Singular [KML69]. SiO [NG67]. Sixteen [Bra68b]. Sky [HS61]. Slab [FL64, Mar69c]. Sliding [Bro69]. Sliding-Scale [Bro69]. Slope [Pro67b]. Slow [Che64]. Slow-Wave [Che64]. Small [RB64, Sch65]. Smoothing [Mus63, Lev61]. SNOBOL3 [FGP66]. Soil [Wei66]. Soil-Foundation [Wei66]. Solar [GS64b, KR62, Kle61, RGS63, SGB^{+}63]. Solderless [GPZ63]. Solid [BSW62, CDJ69, Eld68, HGW^{+}67, Mar66a, Tru60]. Solid-State [BSW62, Eld68, HGW^{+}67]. Solomon [ET69]. Solubilities [Tru60]. Solution [CNN67, FGVH62, Liu69, Log67, Mea62, SK61a, Ben67a]. Solutions [Lin65, Mil68, Row62, SK61b, San65a, You63a, Ben62d, San64d, San65b, San67c, Wil68]. Solving [Ger69b, Kat65]. Some [Ben65c, Ben66b, Ben67b, Bug65, CK68, Dar63b, DP68, Fla62, GS65a, Kur69, Mag68, Mar61a, Mil69b, Nas67, Nas68, Pra68, San64e, San65c, San65d, San67b, SW69a, San69c, San69b, SW69b, Sav67, ST69, Sle60, VZ67, Web62, Wol69, Ric62c, San63a]. Sounds [Bor63]. Source [Aik69, Pil68, Rai68c, Wyn68, Pil69, WZ69]. Sources [Hei67]. Space [CHH61, Dav64, HS63, OS63, SS67, LP62, WZ69, BR63b]. Space-Charge [SS67]. Spacecraft [BEH^{+}63, BM63, BFT^{+}63, BBM^{+}63, DHWM63, DJS63, HKM^{+}63]. Spacing [RP68]. Speakerphone [Bus60, CG61]. Special [Bar65]. Specified [McC62].

Spectra [AMS65, AS65, Lun69, Oss64, Rob69b]. Spectral [Bar64a, Bar64b, BR63a, Bor63, GS64b, PR69, Rai68c]. Spectrum [Buc66, Cro67, Hal69, Hog63, Liu68a, Mar69d]. Specularly [HGG62]. Speech [Ado63, Bor63, Bra65a, Bra65b, Bra69, Dos65, MS62, Rab68, SH67]. Speed [Cha67, Kur66, O’N69]. Speed-Up [O’N69]. sphere [Wyn67]. Spherical [GK64b, RB64]. Spheroidal [Sle64, LP61, LP62, SP61, SS65]. Spiking [Kle64]. Spin [Hil63, Lie60, TC64, Yu63]. Spin-Axis [Hil63, Yu63]. Spin-Orbit [Lie60]. Spin-Precession [Yu63]. Spin-Stabilized [TC64]. Splicing [GPZ63]. Splitting [MB61]. Spot [Coo61]. Spring [Bro67a, GS65b, Sla64]. Sputtered [GS69b]. Square [Rob69a, RH60]. squares [Lev61]. Stability [And69, Bos62, Dar63b, Dra67b, DP68, Guc62, KG66, Lie60, Pra67, Row64b, San64a, San64c, San65d]. Stabilized [TC64]. Stable [Col65]. Stage [Ben62c]. Stages [BR63b]. Standard [Kes68]. Standby [Kib61]. State [BSW62, CDJ69, Eld68, Fra68, Ger66, HGW^{+}67, McC62, Par68a]. States [NM67]. Station [CHLS63, CG65b, Gig65, GT65, KSW69, UCE^{+}63, Jak63, SCM^{+}63]. Stationary [Dal67, Kad66, Kro67, BDEL66]. Statistical [Bra65a, Bra68b, Cla68, FG69, Fre69, GWB67, Kyo69, Lee67b, Liu68b, Mar65b, Ste66, Suy65, Wyn69c]. Statistics [GP64, Hog69]. Steel [Sla64, Zup60]. Steerable [VZ67]. Steering [CKT63, KNG61]. Step [Dal69, DGS68]. Stereograms [JJ68]. Stereoscopic [JM62]. stiff [SS68]. Stimulated [Mar62, Mar63]. Stimulating [Bor63]. Stimuli [Mat61]. Stochastic [Liu68b, Tak61, Ben60a]. Storage [GC68, Kah67, SS69b, Tay68b]. Store [AGGK64, Coo61, GHS64, Ott62]. Store-and-Forward [Ott62]. Stored [HHS64]. Stores [SS69b]. story [Lap13b].

Straight [Bat69]. Straightness [You65a]. Strain [GF60, MM68]. Strategy [Wyn69a]. Stream [Kai68]. Stress [DGS68]. Stresses [RE62]. Strip [Hsi67]. Structural [DGPS65]. Structure [Bur60, Che64, Enl67, GK62, HDSD64, Sta67, Mac65, HKM^{+}63]. Structures [Coy65, Kin66, LU64, San62a]. Studies [Dei60, Pop68]. Study [Coy65, Gar60, GK62, HS64, KH66, Kru69, Man69, Rob69b, Web64]. Stuffing [Wit65]. Subjected [Wei65]. Subjective [Bra67, BMP67, Bro67b, Fla62, Kle67, RK63]. Submarine [And69, BDV64, EFK^{+}64, HVW64]. subsequent [Lan60]. Subsidiary [KS66]. Substituted [GWES64]. Substitution [Sla64]. Substitutional [GWES64]. Substrate [MM66b]. Substrates [Hei67]. Suffered [DeL65]. Sulfate [GK62]. Sum [Cro67, Jag65]. Sums [Bar65, Gra65b, Gra65c, Nas67, Mar67c]. Sun [Wil69]. Superconductor [Hsi67]. Supergroup [GAPB63]. Supplies [AEGH63, BM69]. Supply [BFT^{+}63, CDW63, GBW66, SD61]. Suppression [BR66, Ble68, Jul67]. Suppressor [BH63]. Suppressors [Mar66e, RK63]. sure [Kad65a]. Surface [BCDY68, Cur62, Cut61, Lin62, Mar69c, MW67, PBBS63, WG68, Zuc68]. Surface-Wave [WG68]. Surfaces [JJ68, Log67, RC65, Sem65]. Surges [BG61a]. Survey [Ado63, Hig65, Nas64]. Switch [BWY69, Kur66, PS69]. Switched [Bus60, CG61, Ell65, TW64, Wel66a]. Switches [Zup60]. Switching [And63, BD60, DDH64, FLLR60, FH64, GL61, GS64a, GN65, Hae63, HTZ64, Hig65, JJ60, Joe68, MR60, UV64]. Symbols [Ber67a, Wol69, GM66]. Symmetric [McC65]. Symmetry [Jul67]. Synchronization [BKRS68, CK68, GK66, Kar66, Sca63, Ton66, Wit65]. synchronize [San69a]. Synchronized [Bri66, Bri67].

Synchronizing [Pie69, San69c, Sav67]. Synchronous [Arm61, Gia60, Lub68, RP68, Sal67, Smi68]. Synthesis [Ano61h, Cha66, Jag65, Lee69a, Mar68, PW68, Rab68, Rao69, San60, San61, Spa68, Spa69, UW69]. Synthesizing [Woj60]. System [AC69, Ano66a, Ano69k, BCD64, BSW62, BST65, BBHH69, BI66, BDV64, Bro63, BHT64, BQTY69, Buc69a, Cha69a, CHK64, CL63, CC63, CCJ60, Cur60, Cur63, Dav62, DFHT69, EFK^{+}64, FBL62, FHS67, FP65, GH60, GHS61, Gib60, GL61, GKLL63, GP63, Gre64, GN65, Gri64, HM63, HW61, HBK63, HSW63, HHJP68, HTZ64, HF61, HVW64, HOW63, HH61, HD61, Hyd64, IKM67, JM68, KKV64, KS63, Kes68, Kib61, KL61, KNG61, Kur66, Lar69, LZ64, Lio66, Liu68a, LNI63, Mar61b, MMW^{+}63, Mil63, MV61, Mou67, Mou69, Nel63, Ohm61, PWY63, Rap64, RC61, RJ61, ROB69c, San69c, Shn67, Smi68, SV69, SD61, Til68, Wei66, Wel66a, Zad66b, Ric62a, Ado63, Ano60h, Ano60i, Ano60j, Ano60k, Ano60l, Ano60m, BFT^{+}63, HTZ64, Hig65, Kel64, Nas68, NEH68, UV64]. System [WC69]. Systematic [BKR63, Mac64, Mac63]. Systems [And69, Arm61, BT63, BH62a, BKRS68, BD60, Bri66, Bri67, CF68, CCM63, CT69, Cro66, Cro67, Cur62, Enl63, FC68, FJJK66, GR67, GHS61, Gol66, Hae63, HS69, Hol68, Hub66, Hub67a, Hub67b, HGW^{+}67, Hyd64, Jac67, Kar66, KG66, Kyo69, Luc66, Man69, Mar61e, McD66, McM69, Mil63, NC65, O’N66b, Pie66, Pop68, Pra69, PE69, Pro67b, Rap65, Ric62b, Ric63, San63c, San63b, San64a, SB64, San64e, San64f, SGB69, Smi65, Sun64b, Tay68a, Ben60a, Ben62b, BS64, Gop69, Mea62, Ric62a, Ric62c, San63a, San64c, San65c, San67c, SS68, San69a, Ano64j]. Sze [Ano68k].

T1 [CC63, Fri68, FP65, TY65]. Tables [McC62]. Talking [Gar60]. Tantalum [BCD64]. Tapered [Pro67a, Zad66a]. Tapped [Spa68]. Target [CBL^{+}67, GC68]. TASI [FBL62, MS62]. TD [CCJ60, Cur60, HHJP68]. TD-2 [CCJ60, Cur60]. TD-3 [HHJP68]. TE01 [You65a]. Technical [Ano60h, Ano60i, Ano60j, Ano60k, Ano60l, Ano60m]. Technique [Bra65b, NG67, RB64, Lev61]. Techniques [Bar62, CG65a, Coy65, Ger69b, Luc66, Lun69, Mou67, Ton66]. Technology [Lep66]. teenagers [Lap13b]. Teflon [Sem65]. Teleconferencing [Var66]. Telegraph [GS65a]. Telemetry [CCM63]. Telephone [AGN60, And63, Ben60b, Ben65b, Ben65c, BG61a, BH63, BD60, Dei60, Ell65, EM63, Fri68, Hel66, IKM67, Jak61, Kle67, Mor62a, RK63, TW64, Wel66a, Ben66a, Tak61]. Telephony [Gar60]. Television [BC69, Bro69, BE69, EMR66, Fen69, LM69a, MP69, O’N66a, O’N66b, SS60]. TELSTAR [Hil63, Tho65, BGR63, CCKT63, Dic63, EMT63, GWB67, HBK63, Hit63, HOW63, HS63, Jak63, MMW^{+}63, Nel63, SH63, TS63, UDKM63, Yu63]. Temperature [GN67]. Temperatures [Ste61]. Term [AC61, Guc62]. Terminal [AEH^{+}69, BGLS68, BH65, EH65a, HW66, HH61, LW65, MSV62, SG62]. Terminals [BST65, Coy63, GBW66, GAPB63, May65]. Terms [Mar69d, Ric69]. Terrestrial [BR63b]. Test [CKS68, DJS63, HD61, Mau67]. Testing [HTZ64]. Tests [Cur63, HBK63]. TH [GHS61, GL61, HW61, HF61, HH61, HD61, KL61, SG61, SD61]. Their [CMS^{+}64, McC65, SGB^{+}63]. Theorem [San65a, BS64, Mac63, Mar67c]. Theorems [Mag68, Wan61, San67c, San69b, SW69b]. Theoretic [SW69a]. Theoretical [Man69, Oss64]. Theory [Ano60g, Ano64j, Ben65c, BK62, Cla68, GS65a, Ham64, HU67, Kin66, KPDA60, Kog69, Kur65, MS66, Mil69b, ML68, NM64, NM67, PDS64, San63b, SB64, San66a, Sch62b, Sle60, Sun64a, Thi69, Van60, Woj60, Ben61c, Ben62b, Ben63b, Coc65, Mac60, Ric68, San65c, Tak61, Ano66l]. Thermal [AT60, Bec64b, Ber64b, Dra68, Fri68, HKM^{+}63]. Thermally [NM67]. thermodynamic [Ben63b]. Thermoelastic [RE62]. Thick [Kog69, Mar64a]. Thin [CUM68, Dov67, Mar66c, MM66b, Sin66, SSD69]. Third [Ric69]. those [San65d]. Three [Ben62c, Man63, Woo64]. Three-Conductor [Man63]. Three-Stage [Ben62c]. Threshold [Hub66]. Time [Ben67c, BWY69, Dar63b, EM63, JJ60, MR60, Mar61c, Mar61e, O’N69, San64a, Spa69, WC69, Gop69, LP62, San64c, Wyn69b, MM66a]. time- [LP62]. time-invariant [Gop69]. Time-Separation [JJ60, MR60]. Time-Varying [Dar63b, San64a, San64c]. Timed [AG62]. Times [DeL65, Rai68b, Des63, Tak62, Tak63a]. Timing [Kin62a, Kin62b, Kin64, Man69, Sal67]. TJ [GH60]. TL [BSW62, HSW63]. TL-2 [FJJK66]. TM [FJJK66]. TM-1 [FJJK66]. TM-1/TL-2 [FJJK66]. Tolerances [Ung61d]. Toll [Nas64, Nas68]. Tone [O’N69]. Tool [Zup60]. Topological [Kla61, Ben62a]. Topology [Sin66]. Touch [O’N69]. Touch-Tone [O’N69]. Tracker [AHMS63, Wil69]. Tracking [De 61, Hel65, Nel63, US61]. Traffic [Ben60b, Ben61b, Ben65b, Ben65c, Ben67b, Lin64, Ott62, Web62, Ben60a, Ben63a, Ben63b, Ric62a, Tak61]. Trains [Bar64a]. Tranmitter [FJL68]. Transducer [Lio66]. Transducers [GF60]. Transfer [FS60, HGG62, Lee69a, Spa68]. Transform [Lun69, RSR69]. Transformer [BR63b]. Transformer-Coupled [BR63b]. Transformerless [San61]. Transformers [Bog69]. Transforms [Gen68]. Transient [Wei65]. Transistor [BR63b, Kib65b, Kib65a, Koe67, Kur65, LK61, Mur68, Nar67, SW69a, Tho68, San69b]. Transistorized [BH62a]. Transistors [Lee66, PBBS63]. Transit [MM66a]. Transition [Ben60b]. Transitions [PDS64]. Translations [UV64]. Transmission [Aar62, AGN60, BKS66, BS63, BI66, BC69, Buc67, Buc69a, BE69, Cha66, Cha67, CF68, Cro66, DD68, Dos65, Eng65, EMR66, Fla62, Fra68, FC68, FA64, GHS69, Ger69a, GS65a, Gig65, Gre69, HR63, Hsi67, IKM67, Kin66, Kle67, Lar69, Luc63, Luc68, Man63, MS64, Mar66a, Mar67b, Mar67a, Maz68, Mil63, Mor62a, Nas68, NEH68, O’N66b, Pil68, Pro67a, Rap64, Ric63, RW62, Row64a, Row64b, Sal67, SS69a, San69c, Sch68b, Smi65, Smi68, Sun61, Sun64a, Suy65, Var66, Zad66a, Lee69b, Mac61, San69a, Ano70]. Transmitted [Smi65]. Transmitter [BGLS68, Bra68a, Fri68, GPW63, HP68, HH61, JRS68, SB61, SG61]. Transport [Van60]. Transversal [BR66, Lyt68, SS60]. Transverse [Gor60]. Trapping [Van60]. Traveling [BLOR63, Bra68a, CT64, CHLS63, DKS61, DA63, GS62, HDSD64, Lin65, SD64, Mac60]. Traveling-Wave [BLOR63, Bra68a, CT64, CHLS63, DKS61, DA63, GS62, HDSD64, SD64, Mac60]. Treatment [Fre69, Mar65b]. Tree [RH60]. Triply [GT65]. Triply-Folded [GT65]. Troposcatter [Sun64a, Sun64b]. Tropospheric [BT63, Ung66]. Trouble [TU62]. Truncated [Bro63, BHT64, Nas67]. Trunk [Ben61b, BBC^{+}64, Lin64, Des63, Tak62]. Trunks [Dav62, IKM67, NEH68]. Tube [BLOR63, Bra68a, BE69, CHLS63, Coo61, CBL^{+}67, CL69, Pfa63]. Tubes [CLW69, HVW64]. Tubular [Ano64k, Dai68, MM64]. Tuned [Bos62]. Tunnel [San62b, Sch63, San62a]. Turbulence [SC65, SC67]. Turbulent [Bug65]. TV [Bra67, BMP67, Bro67b, Mou67]. Twisted [Ble68]. TWM [Che64]. Two [Bra69, Cro67, Eva69, FRY63, Ham64, HM67, JM62, Kad66, KM61, LK61, Lee66, LZ64, Mil68, Rao69, San67c, Sch60, Sta67, WY68, Wol69, Yu65, Ano70, Kad67b, Lee69b]. Two-Body [FRY63, Yu65]. Two-Group [Sch60]. Two-Gyro [LZ64]. Two-Inch [KM61]. Two-Mile [WY68]. Two-Pole [Sta67]. Two-Port [Eva69, LK61, Lee66]. Two-Variable [Rao69]. Two-Way [Bra69]. Type [AEGH63, Ano66l, Che64, GWES64, Kah64, NM67, San66a]. Types [HM67].

U.S [KS66]. Ultimately [Ben62d]. Ultraprecise [War60]. Ultrasonic [SS69b]. Unbalanced [RE62]. Uncertainty [Sle64, LP61, LP62, SP61]. Uncontrolled [RR62]. Underground [Glo67b]. Unidirectional [GS62]. Unified [HU67, ML68, Smi68]. Uniform [HS69, Ric68, UW69]. Unilateral [Ham64]. Unit [BQTY69, PS69, SS67, Wyn67]. Unit-Cube [SS67]. Units [Ano66a, BWY69, FJL68, War60]. Unlike [Ber64a]. Unreliable [Tay68a, Tay68b]. untold [Lap13b]. Unwanted [Ano62a, KK62]. Up-Converter [US61]. Upconverter [GM67]. Upconverters [Osb69]. Upper [GM67, SK68]. Upper-Sideband [GM67]. Usage [WC69]. Use [Ano64a, Dei60, Ell66, Guc62, HR63, KR62, Kur62, Sha63, Til69]. Used [HDSD64, Pil69]. Uses [Kib69]. Using [Ber64b, Ber64a, BR63b, CG65a, DG67, Ger69b, Gol66, GP64, Ham64, Hol66, HP60, Kah67, Log61, Mos65, Nar67, O’N69, RB64, SK65, SSD69, Spa68, Tab67, Wit65, Wyn69a].

V [Ano68k, LS68]. V. [San65d]. Vacuum [BCD64]. Validity [Kyo69]. values [Mac60]. Vapor [Ham67, HS61]. Varactor [Bur65, Dra67a, Dra67b, DP67, DP68, GM67, Kah64, Lin62, Osb69, Pra67].

Variable [KU61, Kur62, Par68a, Rao69]. Variable-Capacitance [KU61]. Variables [Bro63, BHT64, JM62, Nas67, Ano70, Lee69b, Mar67c]. Variance [Dal67]. Variation [Bur65, Lub68, Luc63]. Variational [Bar62]. Variations [Ste66]. Variolossers [Bod69]. Various [FGVH62, Pie66, Sem65, Tak63a]. Varying [Dar63a, Dar63b, San64a, Gor60, San64c]. Vehicle [EMT63]. Velocity [Kla60]. Verification [HM67]. VHF [SBF68]. via [EMR66]. Vibrating [Wei66]. Vibration [Liu69]. Vibrations [Pee63]. Video [Fra66, Mou69]. View [Smi68]. Vision [AC61]. Visual [Lim69]. Vocoder [FG66]. Voice [Bus60, CG61, McD66, Sch60]. Voice-Frequency [Sch60]. Voice-Switched [Bus60, CG61]. Voiding [Ell66]. Voltage [SR67]. Volterra [Nar67]. Volumes [Ado63].

W [Ano60g, Ano64l, Ano64j, Ano66l]. W. [Ano70]. Wafer [Sha63]. Wall [Mar69d]. Water [Ham67, HS61]. Watt [Kib65b, Kib65a, MM66a]. Wave [Aul65, BLOR63, Bra68a, Che64, CT64, CHLS63, Con66, DKS61, DA63, GS62, HDSD64, HGW^{+}67, Kog69, LS69, LU64, Mar67b, MM69, Mil64b, PR69, Rai67a, Rai67b, Rai68b, SGB69, SD64, Sle64, Sta67, WG68, LP61, LP62, Mac60, SP61, SS65]. Wave-Optics [Mar67b]. Waveform [Oss64]. Waveforms [Lim69]. Wavefronts [Rus66]. Waveguide [Abe67, AG68, CNN67, Con66, Enl63, GW68a, Glo67b, GS69a, GS69b, Goe69, Kib69, KM61, KU62, Mar69b, Mar64b, Mar69c, MD69, McK67, RW62, Sch62b, Ste65, Ste66, Ung60, Ung61a, Ung61b, Ung61c, Ung61d, Ung62, Wan68, WY68, Wu69, You65a]. Waveguides [Bat69, Dal68, Dal69, GW68b, MS64, Mar66b, Mar65b, Mar69d, MM68, NS65, Nag67, WG68]. Wavelength [BG61b, Cha69a]. Wavelength-Dependent [Cha69a]. Wavelengths [HS61]. Waves [AMS65, Ben65a, Hog63, Mil68, Mil69b, Ric69, Gor60]. Wavy [Ber65b]. Way [Bra69]. Weakly [CG64]. Weather [CG65b]. weights [Mac63]. Welding [Coy63]. Which [McM69, San69a]. White [Bur66, LPSL66, Wyn69a]. White-Light [LPSL66]. who [Lap13b]. Wide [DH63]. Wide-Angle [DH63]. Wideband [BR63b, Dar64, GK64a, TY65]. Width [Hub66]. Wind [Coy65]. Winding [Ung61d]. Wire [Bro67a, DGS68, GPZ63, GS65b, Sla64, Var66]. Wires [Coy63]. without [Guc62]. Wollaston [Ano64a, Tab67]. Word [GP64]. Word-Frame [GP64]. Work [Kru69]. Work-Scheduling [Kru69]. world [Lap13a].

X [CLW69, GK62, MM66a]. X-Band [MM66a]. X-Ray [CLW69, GK62].

young [Lap13a]. Yttrium [GWES64].

Zero [SK68, Hsu62, SS65]. Zero-Crossing [SK68]. zero-memory [Hsu62]. Zeros [Rai68a]. Zinc [Dew60].

References

Aaron:1962:PTE

[Aar62] M. R. Aaron. PCM transmission in the exchange plant. The Bell System Technical Journal, 41(1):99–141, January 1962. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol41/bstj41-1-
99.pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-1-99.pdf.

Abele:1968:SBR

[AART68] T. A. Abele, A. J. Alberts, C. L. Ren, and G. A. Tuchen. Schottky barrier receiver modulator. The Bell System Technical Journal, 47(7):1257–1287, September 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-7-1257.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-7-1257.pdf.

Anderson:1965:DDB

[ABDS65] R. R. Anderson, W. R. Bennett, J. R. Davey, and J. Salz. Differential detection of binary FM. The Bell System Technical Journal, 44(1):111–159, January 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-1-111.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-1-111.pdf.

Abele:1967:HQW

[Abe67] T. A. Abele. A high-quality waveguide directional filter. The Bell System Technical Journal, 46(1):81–104, January 1967. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol46/bstj46-1-
81.pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-1-81.pdf.

Averbach:1961:STM

[AC61] E. Averbach and A. S. Coriell.

Short-term memory in vision. The Bell System Technical Journal, 40(1):309–328, January 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-1-309.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-1-309.pdf.

Anderson:1969:ERC

[AC69] R. E. Anderson and C. R. Crue.

The equalizer remote-control system. The Bell System Technical Journal, 48(4):953–981, April 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-4-953.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-4-953.pdf.

Andrews:1969:SFC

[ADH^{+}69] Robert J. Andrews, John J.

Driscoll, John A. Herndon, Philip C. Richards, and L. Ralph Roberts. Service features and call processing plan. The Bell System Technical Journal, 48(8):2713–2764, October 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-8-2713.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-8-2713.pdf.

Adoo:1963:SVB

[Ado63] Kathryn L. Mc Adoo. Speech volumes on Bell System message circuits – 1960 survey. The Bell System Technical Journal, 42(5):1999–2012, September 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-5-1999.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-5-1999.pdf.

Albert:1963:CST

[AEGH63] W. G. Albert, J. B. Evans, Jr., J. J. Ginty, and J. B. Harley. Carrier supplies for L-type multiplex. The Bell System Technical Journal, 42(2):279–317, March 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-2-279.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-2-279.pdf.

Albert:1969:TA

[AEH^{+}69] W. G. Albert, J. B. Evans, Jr., T. J. Haley, T. B. Merrick, and T. H. Simmonds, Jr. Terminal arrangements. The Bell System Technical Journal, 48(4):993–1040, April 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-4-993.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-4-993.pdf.

Aaron:1969:RDM

[AFMP69] M. R. Aaron, J. S. Fleischman, R. A. McDonald, and E. N. Protonotarios. Response of delta modulation to Gaussian signals. The Bell System Technical Journal, 48(5):1167–1195, May/ June 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-5-1167.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-5-1167.pdf.

Aaron:1962:PDP

[AG62] M. R. Aaron and J. R.

Gray. Probability distribution for the phase jitter in selftimed reconstructive repeaters for PCM. The Bell System Technical Journal, 41(2):503–558, March 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol41/bstj41-2-503.
pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-2-503.pdf.

Amitay:1968:ACW

[AG68] N. Amitay and Victor Galindo.

The analysis of circular waveguide phased arrays. The Bell System Technical Journal, 47(9):1903–1932, November 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol47/bstj47-9-1903.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-9-1903.pdf.

Ault:1964:NEP

[AGGK64] C. F. Ault, L. E. Gallaher, T. S. Greenwood, and D. C. Koehler. No. 1 ESS program store. The Bell System Technical Journal, 43(5):2097–2146, September 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-5-2097.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-5-2097.pdf.

Alexander:1960:CTN

[AGN60] A. A. Alexander, R. M. Gryb, and D. W. Nast. Capabilities of the telephone network for data transmission. The Bell System Technical Journal, 39(3):431–476, May 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol39/bstj39-3-431.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol39-1960/
                                        articles/bstj39-3-431.pdf.

Anders:1963:PT

[AHMS63] J. V. Anders, E. F. Higgins, Jr., J. L. Murray, and F. J. Schaefer, Jr. The precision tracker. The Bell System Technical Journal, 42(4):1309–1356, July 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-4-1309.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-4-1309.pdf.

Aiken:1967:EPB

[Aik67] R. T. Aiken. Error probability for binary signaling through an multipath channel. The Bell System Technical Journal, 46(7):1601–1631, September 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-7-1601.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-7-1601.pdf.

Aiken:1969:PPS

[Aik69]  R. T. Aiken.  Propagation

from a point source in a randomly refracting medium. The Bell System Technical Journal, 48(5):1129–1165, May/ June 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-5-1129.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-5-1129.pdf.

Arthurs:1965:BJB

[AK65] E. Arthurs and J. L. Kelly, Jr. B.S.T.J. brief: On the simultaneous measurement of a pair of conjugate observables. The Bell System Technical Journal, 44(4):725–729, April 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol44/bstj44-4-725.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol44-1965/
                                        articles/bstj44-4-725.pdf.

Aikens:1960:EMC

[AL60] A. J. Aikens and D. A. Lewinski. Evaluation of message circuit noise. The Bell System Technical Journal, 39(4):879–909, July 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol39/bstj39-4-879.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol39-1960/
                                        articles/bstj39-4-879.pdf.

Abele:1968:MG

[AL68] T. A. Abele and D. J. Leonard. Microwave generator. The Bell System Technical Journal, 47(7):1301–1322, September 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol47/bstj47-7-1301.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-7-1301.pdf.

Anderson:1965:SCA

[AMS65] R. R. Anderson, J. E. Mazo, and J. Salz. Spectra for a class of asynchronous FM waves. The Bell System Technical Journal, 44(9):2149–2163, November 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-9-2149.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-9-2149.pdf.

Andrews:1963:TSE

[And63]  E. G. Andrews.  Telephone

switching and the early Bell Laboratories computers. The Bell System Technical Journal, 42(2):341–353, March 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-2-341.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-2-341.pdf.

Anderson:1969:OSP

[And69]  Cleo D. Anderson.  Over-

load stability problem in submarine cable systems. The Bell System Technical Journal, 48(6):1853–1864, July/ August 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-6-1853.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-6-1853.pdf.

Anonymous:1960:CIa

[Ano60a] Anonymous. Contributors to this issue. The Bell System Technical Journal, 39(1):261–264, January 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol39/bstj39-1-261.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol39-1960/
                                        articles/bstj39-1-261.pdf.

Anonymous:1960:CIb

[Ano60b] Anonymous. Contributors to this issue. The Bell System Technical Journal, 39(2):428–430, March 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol39/bstj39-2-428.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol39-1960/
                                        articles/bstj39-2-428.pdf.

Anonymous:1960:CIc

[Ano60c] Anonymous. Contributors to this issue. The Bell System Technical Journal, 39(3):742–744, May 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol39/bstj39-3-742.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol39-1960/
                                        articles/bstj39-3-742.pdf.

Anonymous:1960:CId

[Ano60d] Anonymous. Contributors to this issue. The Bell System Technical Journal, 39(4):1121–1124, July 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-4-1121.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-4-1121.pdf.

Anonymous:1960:CIe

[Ano60e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 39(5): 1379–1380, September 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-5-1379.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-5-1379.pdf.

Anonymous:1960:CIf

[Ano60f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 39(6): 1670–1673, November 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-6-1670.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-6-1670.pdf.

Anonymous:1960:EJR

[Ano60g] Anonymous. Erratum: J. R. Klauder and A. C. Price and S. Darlington and W. J. Albersheim, The Theory and Design of Chirp Radars, BSTJ 39(7) 745–808 (1960). The Bell System Technical Journal, 39(5):1192, September 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-5-1192.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-5-1192.pdf.

See [KPDA60].

Anonymous:1960:RMBa

[Ano60h] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 39(1):257–260, January 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-1-257.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-1-257.pdf.

Anonymous:1960:RMBb

[Ano60i] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 39(2):423–427, March 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-2-423.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-2-423.pdf.

Anonymous:1960:RMBc

[Ano60j] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 39(3):733–741, May 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-3-733.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-3-733.pdf.

Anonymous:1960:RMBd

[Ano60k] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 39(4):1113–1120, July 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-4-1113.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-4-1113.pdf.

Anonymous:1960:RMBe

[Ano60l] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 39(5):1375–1378, September 1960. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol39/bstj39-5-1375.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol39-1960/
                                        articles/bstj39-5-1375.pdf.

Anonymous:1960:RMBf

[Ano60m] Anonymous. Recent monographs of Bell System technical papers not published in this journal. The Bell System Technical Journal, 39(6):1665–1669, November 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol39/bstj39-6-1665.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol39-1960/
                                        articles/bstj39-6-1665.pdf.

Anonymous:1961:CIe

[Ano61a] Anonymous. Contributors to the issue. The Bell System Technical Journal, 40(5): 1455–1457, September 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol40/bstj40-5-1455.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol40-1961/
                                        articles/bstj40-5-1455.pdf.

Anonymous:1961:CIa

[Ano61b] Anonymous. Contributors to this issue. The Bell System Technical Journal, 40(1):349–351, January 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-1-349.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-1-349.pdf.

Anonymous:1961:CIb

[Ano61c] Anonymous. Contributors to this issue. The Bell System Technical Journal, 40(2):645–647, March 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-2-645.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-2-645.pdf.

Anonymous:1961:CIc

[Ano61d] Anonymous. Contributors to this issue. The Bell System Technical Journal, 40(3):971–974, May 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-3-971.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-3-971.pdf.

Anonymous:1961:CId

[Ano61e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 40(4):1235–1238, July 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-4-1235.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol40-1961/
                                        articles/bstj40-4-1235.pdf.

Anonymous:1961:CIf

[Ano61f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 40(6): 1745–1748, November 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol40/bstj40-6-1745.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol40-1961/
                                        articles/bstj40-6-1745.pdf.

Anonymous:1961:GSJ

[Ano61g] Anonymous. G. E. Schindler, Jr., new editor of B.S.T.J. The Bell System Technical Journal, 40(2):452, March 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol40/bstj40-2-452.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol40-1961/
                                        articles/bstj40-2-452.pdf.

Anonymous:1961:SPA

[Ano61h] Anonymous. Synthesis of Nport active RC networks. The Bell System Technical Journal, 40(1):329–347, January 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol40/bstj40-1-329.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol40-1961/
articles/bstj40-1-329.pdf.

Anonymous:1962:CQD

[Ano62a] Anonymous. Comment on ‘Discrimination against unwanted orders in the Fabry–Perot resonator’. The Bell System Technical Journal, 41(4):1475–1476, July 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol41/bstj41-4-1475.
pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-4-1475.pdf.

Anonymous:1962:CIa

[Ano62b] Anonymous. Contributors to this issue. The Bell System Technical Journal, 41(1):383–386, January 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol41/bstj41-1-383.
pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-1-383.pdf.

Anonymous:1962:CIb

[Ano62c] Anonymous. Contributors to this issue. The Bell System Technical Journal, 41(2):797–801, March 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol41/bstj41-2-797.
pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-2-797.pdf.

Anonymous:1962:CIc

[Ano62d] Anonymous. Contributors to this issue. The Bell System Technical Journal, 41(3):1171–1173, May 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol41/bstj41-3-1171.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol41-1962/
                                        articles/bstj41-3-1171.pdf.

Anonymous:1962:CId

[Ano62e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 41(4):1477–1480, July 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol41/bstj41-4-1477.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol41-1962/
                                        articles/bstj41-4-1477.pdf.

Anonymous:1962:CIe

[Ano62f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 41(5): 1691–1693, September 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol41/bstj41-5-1691.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol41-1962/
                                        articles/bstj41-5-1691.pdf.

Anonymous:1962:CIf

[Ano62g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 41(6): 1875–1877, November 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol41/bstj41-6-1875.
pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-6-1875.pdf.

Anonymous:1962:F

[Ano62h] Anonymous. Foreword. The Bell System Technical Journal, 41(4):1175, July 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol41/bstj41-4-1175.
pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-4-1175.pdf.

Anonymous:1963:CIa

[Ano63a] Anonymous. Contributors to this issue. The Bell System Technical Journal, 42(1):203–206, January 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-1-203.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-1-203.pdf.

Anonymous:1963:CIb

[Ano63b] Anonymous. Contributors to this issue. The Bell System Technical Journal, 42(2):521–525, March 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-2-521.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-2-521.pdf.

Anonymous:1963:CIc

[Ano63c] Anonymous. Contributors to this issue. The Bell System Technical Journal, 42(3):737–738, May 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-3-737.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-3-737.pdf.

Anonymous:1963:CId

[Ano63d] Anonymous. Contributors to this issue. The Bell System Technical Journal, 42(4):1909–1934, July 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-4-1909.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-4-1909.pdf.

Anonymous:1963:CIe

[Ano63e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 42(5): 2485–2491, September 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-5-2485.

pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-5-2485.pdf.

Anonymous:1963:CIf

[Ano63f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 42(6): 2961–2966, November 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-6-2961.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-6-2961.pdf.

Anonymous:1964:BJB

[Ano64a] Anonymous. B.S.T.J. briefs: The use of wollaston prisms for a high-capacity digital light deflector. The Bell System Technical Journal, 43(3):1153–1154, May 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-3-1153.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-3-1153.pdf.

Anonymous:1964:CIa

[Ano64b] Anonymous. Contributors to this issue. The Bell System Technical Journal, 43(1):329–332, January 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-1-329.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-1-329.pdf.

Anonymous:1964:CIb

[Ano64c] Anonymous. Contributors to this issue. The Bell System Technical Journal, 43(1):561–563, January 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-1-561.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-1-561.pdf.

Anonymous:1964:CIc

[Ano64d] Anonymous. Contributors to this issue. The Bell System Technical Journal, 43(2):817–820, March 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-2-817.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-2-817.pdf.

Anonymous:1964:CId

[Ano64e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 43(3):1149–1151, May 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-3-1149.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-3-1149.pdf.

Anonymous:1964:CIe

[Ano64f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 43(4):1461–1467, July 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-4-1461.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-4-1461.pdf.

Anonymous:1964:CIf

[Ano64g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 43(4):1811–1814, July 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-4-1811.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-4-1811.pdf.

Anonymous:1964:CIg

[Ano64h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 43(5): 2593–2605, September 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-5-2593.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-5-2593.pdf.

Anonymous:1964:CIh

[Ano64i] Anonymous. Contributors to this issue. The Bell System Technical Journal, 43(6): 3059–3063, November 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-6-3059.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-6-3059.pdf.

Anonymous:1964:EWS

[Ano64j] Anonymous. Errata: I. W. Sandberg, On the Theory of Linear Multi-Loop Feedback Systems, BSTJ 42(2) 355–382 (1963). The Bell System Technical Journal, 43(4):1600, July 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-4-1600.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-4-1600.pdf.

See [San63b].

Anonymous:1964:EDM

[Ano64k] Anonymous. Erratum: D. Marcuse and S. E. Miller, Analysis of a Tubular Gas Lens, BSTJ 43(4) 1759–1782 (1964). The Bell System Technical Journal, 43(6):2766, November 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-6-2766.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/

articles/bstj43-6-2766.pdf.

See [MM64].

Anonymous:1964:EMW

[Ano64l] Anonymous. Erratum: M.

W. Bowker and W. G. Nutt and R. M. Riley, Design of Armorless Ocean Cable, BSTJ 43(4) 1185–1208 (1964). The Bell System Technical Journal, 43(6):2766, November 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-6-2766.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-6-2766.pdf.

See [BNR64].

Anonymous:1965:CIa

[Ano65a] Anonymous. Contributors to this issue. The Bell System Technical Journal, 44(1):199–201, January 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-1-199.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-1-199.pdf.

Anonymous:1965:CIb

[Ano65b] Anonymous. Contributors to this issue. The Bell System Technical Journal, 44(2):365–367, February 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-2-365.

                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol44-1965/
                                        articles/bstj44-2-365.pdf.

Anonymous:1965:CIc

[Ano65c] Anonymous. Contributors to this issue. The Bell System Technical Journal, 44(3):539–541, March 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol44/bstj44-3-539.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol44-1965/
                                        articles/bstj44-3-539.pdf.

Anonymous:1965:CId

[Ano65d] Anonymous. Contributors to this issue. The Bell System Technical Journal, 44(4):723–724, April 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol44/bstj44-4-723.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol44-1965/
                                        articles/bstj44-4-723.pdf.

Anonymous:1965:CIe

[Ano65e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 44(5):933–935, May/ June 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol44/bstj44-5-933.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol44-1965/
                                        articles/bstj44-5-933.pdf.

Anonymous:1965:CIf

[Ano65f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 44(6):1215–1217, July/ August 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-6-1215.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-6-1215.pdf.

Anonymous:1965:CIg

[Ano65g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 44(7): 1521–1524, September 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-7-1521.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-7-1521.pdf.

Anonymous:1965:CIh

[Ano65h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 44(8):1805–1807, October 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-8-1805.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-8-1805.pdf.

Anonymous:1965:CIi

[Ano65i] Anonymous. Contributors to this issue. The Bell System Technical Journal, 44(9): 2199–2202, November 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-9-2199.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-9-2199.pdf.

Anonymous:1965:CIj

[Ano65j] Anonymous. Contributors to this issue. The Bell System Technical Journal, 44(10):2447– 2449, December 1965. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol44/bstj44-10-2447.pdf;
http://www.alcatel-lucent.
com/bstj/vol44-1965/articles/
bstj44-10-2447.pdf.

Anonymous:1966:BJR

[Ano66a] Anonymous. B.S.T.J. recommendation for adopting the international system (SI) of units. The Bell System Technical Journal, 45(6):765–766, July/ August 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-6-765.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-6-765.pdf.

Anonymous:1966:CIa

[Ano66b] Anonymous. Contributors to this issue. The Bell System Technical Journal, 45(1):229–231, January 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-1-229.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-1-229.pdf.

Anonymous:1966:CIb

[Ano66c] Anonymous. Contributors to this issue. The Bell System Technical Journal, 45(2):333–334, February 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-2-333.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-2-333.pdf.

Anonymous:1966:CIc

[Ano66d] Anonymous. Contributors to this issue. The Bell System Technical Journal, 45(3):491–492, March 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-3-491.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-3-491.pdf.

Anonymous:1966:CId

[Ano66e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 45(4):657–658, April 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol45/bstj45-4-657.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol45-1966/
                                        articles/bstj45-4-657.pdf.

Anonymous:1966:CIe

[Ano66f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 45(5):753–754, May/ June 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol45/bstj45-5-753.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol45-1966/
                                        articles/bstj45-5-753.pdf.

Anonymous:1966:CIf

[Ano66g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 45(6):985–987, July/ August 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol45/bstj45-6-985.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol45-1966/
                                        articles/bstj45-6-985.pdf.

Anonymous:1966:CIg

[Ano66h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 45(7): 1193–1195, September 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol45/bstj45-7-1193.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol45-1966/
articles/bstj45-7-1193.pdf.

Anonymous:1966:CIh

[Ano66i] Anonymous. Contributors to this issue. The Bell System Technical Journal, 45(8):1369–1371, October 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-8-1369.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-8-1369.pdf.

Anonymous:1966:CIi

[Ano66j] Anonymous. Contributors to this issue. The Bell System Technical Journal, 45(9): 1663–1665, November 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-9-1663.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-9-1663.pdf.

Anonymous:1966:CIj

[Ano66k] Anonymous. Contributors to this issue. The Bell System Technical Journal, 45(10):1837– 1839, December 1966. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol45/bstj45-10-1837.pdf;
http://www.alcatel-lucent.
com/bstj/vol45-1966/articles/
bstj45-10-1837.pdf.

Anonymous:1966:EWS

[Ano66l] Anonymous. Errata: I. W. Sandberg, A Note on a Type of Optimization Problem that Arises in Communication Theory, BSTJ 45(5) 761–764 (1966). The Bell System Technical Journal, 45(6):992, July/August 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-6-992.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-6-992.pdf.

See [San66a].

Anonymous:1967:CIa

[Ano67a] Anonymous. Contributors to this issue. The Bell System Technical Journal, 46(1):273–275, January 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-1-273.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-1-273.pdf.

Anonymous:1967:CIb

[Ano67b] Anonymous. Contributors to this issue. The Bell System Technical Journal, 46(2):489–490, February 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-2-489.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-2-489.pdf.

Anonymous:1967:CIc

[Ano67c] Anonymous. Contributors to this issue. The Bell System Technical Journal, 46(3):673–675, March 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-3-673.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-3-673.pdf.

Anonymous:1967:CId

[Ano67d] Anonymous. Contributors to this issue. The Bell System Technical Journal, 46(4):835–836, April 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-4-835.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-4-835.pdf.

Anonymous:1967:CIe

[Ano67e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 46(5):1043–1045, May/ June 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-5-1043.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-5-1043.pdf.

Anonymous:1967:CIf

[Ano67f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 46(6):1279–1281, July/ August 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-6-1279.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-6-1279.pdf.

Anonymous:1967:CIg

[Ano67g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 46(7): 1651–1653, September 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-7-1651.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-7-1651.pdf.

Anonymous:1967:CIh

[Ano67h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 46(8):1927–1929, October 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-8-1927.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-8-1927.pdf.

Anonymous:1967:CIi

[Ano67i] Anonymous. Contributors to this issue. The Bell System Technical Journal, 46(9): 2229–2233, November 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-9-2229.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-9-2229.pdf.

Anonymous:1967:CIj

[Ano67j] Anonymous. Contributors to this issue. The Bell System Technical Journal, 46(10):2485– 2488, December 1967. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol46/bstj46-10-2485.pdf;
http://www.alcatel-lucent.
com/bstj/vol46-1967/articles/
bstj46-10-2485.pdf.

Anonymous:1968:CIa

[Ano68a] Anonymous. Contributors to this issue. The Bell System Technical Journal, 47(1):159–160, January 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-1-159.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-1-159.pdf.

Anonymous:1968:CIb

[Ano68b] Anonymous. Contributors to this issue. The Bell System Technical Journal, 47(2):261–263, February 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/

                                        images/Vol47/bstj47-2-261.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-2-261.pdf.

Anonymous:1968:CIc

[Ano68c] Anonymous. Contributors to this issue. The Bell System Technical Journal, 47(3):459–460, March 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol47/bstj47-3-459.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-3-459.pdf.

Anonymous:1968:CId

[Ano68d] Anonymous. Contributors to this issue. The Bell System Technical Journal, 47(4):603–604, April 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol47/bstj47-4-603.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-4-603.pdf.

Anonymous:1968:CIe

[Ano68e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 47(5):823–825, May/ June 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol47/bstj47-5-823.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-5-823.pdf.

Anonymous:1968:CIf

[Ano68f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 47(6):1139–1142, July/ August 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-6-1139.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-6-1139.pdf.

Anonymous:1968:CIg

[Ano68g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 47(7): 1531–1537, September 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-7-1531.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-7-1531.pdf.

Anonymous:1968:CIh

[Ano68h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 47(8):1823–1826, October 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-8-1823.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-8-1823.pdf.

Anonymous:1968:CIi

[Ano68i] Anonymous. Contributors to this issue. The Bell System Technical Journal, 47(9): 2071–2074, November 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-9-2071.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-9-2071.pdf.

Anonymous:1968:CIj

[Ano68j] Anonymous. Contributors to this issue. The Bell System Technical Journal, 47(10):2435– 2438, December 1968. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol47/bstj47-10-2435.pdf;
http://www.alcatel-lucent.
com/bstj/vol47-1968/articles/
bstj47-10-2435.pdf.

Anonymous:1968:EMP

[Ano68k] Anonymous. Erratum: M. P. Lepselter and S. M. Sze, Silicon Schottky Barrier Diode with Near-Ideal I–V Characteristics, BSTJ 47(2) 195– 208 (1968). The Bell System Technical Journal, 47(3):464, March 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-3-464.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-3-464.pdf.

See [LS68].

Anonymous:1969:CIa

[Ano69a] Anonymous. Contributors to this issue. The Bell System Technical Journal, 48(1):317–320, January 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-1-317.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-1-317.pdf.

Anonymous:1969:CIb

[Ano69b] Anonymous. Contributors to this issue. The Bell System Technical Journal, 48(2):497–499, February 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-2-497.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-2-497.pdf.

Anonymous:1969:CIc

[Ano69c] Anonymous. Contributors to this issue. The Bell System Technical Journal, 48(3):813–816, March 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-3-813.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-3-813.pdf.

Anonymous:1969:CId

[Ano69d] Anonymous. Contributors to this issue. The Bell System Technical Journal, 48(4):1095–1099, April 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-4-1095.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-4-1095.pdf.

Anonymous:1969:CIe

[Ano69e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 48(5):1555–1561, May/ June 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-5-1555.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-5-1555.pdf.

Anonymous:1969:CIf

[Ano69f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 48(6):2053–2058, July/ August 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-6-2053.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-6-2053.pdf.

Anonymous:1969:CIg

[Ano69g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 48(7): 2601–2605, September 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-7-2601.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-7-2601.pdf.

Anonymous:1969:CIh

[Ano69h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 48(8):2897–2903, October 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-8-2897.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-8-2897.pdf.

Anonymous:1969:CIi

[Ano69i] Anonymous. Contributors to this issue. The Bell System Technical Journal, 48(9): 3131–3134, November 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-9-3131.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-9-3131.pdf.

Anonymous:1969:CIj

[Ano69j] Anonymous. Contributors to this issue. The Bell System Technical Journal, 48(10):3441– 3443, December 1969. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol48/bstj48-10-3441.pdf;

                                        http://www.alcatel-lucent.
                                        com/bstj/vol48-1969/articles/
                                        bstj48-10-3441.pdf.

Anonymous:1969:FCS

[Ano69k] Anonymous. Foreword: The L-4 coaxial system. The Bell System Technical Journal, 48(4):819–820, April 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-4-819.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-4-819.pdf.

Anonymous:1970:EWC

[Ano70] Anonymous. Erratum: W. C. Y. Lee, An extended correlation function of two random variables applied to mobile radio transmission, BSTJ 48(10) 3423–3440 (1969). The Bell System Technical Journal, 49(2):320, February 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-2-320.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-2-320.pdf.

See [Lee69b].

Armstrong:1961:GMA

[Arm61] D. B. Armstrong. A general method of applying error correction to synchronous digital systems. The Bell System Technical Journal, 40(2):577–593, March 1961. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-2-577.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-2-577.pdf.

Anderson:1965:SDF

[AS65] R. R. Anderson and J. Salz.

Spectra of digital FM. The Bell System Technical Journal, 44(6):1165–1189, July/ August 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-6-1165.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-6-1165.pdf.

Aaron:1966:BJB

[AS66] M. R. Aaron and M. K. Simon.

B.S.T.J. briefs: Approximation of the error probability in a regenerative repeater with quantized feedback. The Bell System Technical Journal, 45(10): 1845–1847, December 1966. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol45/bstj45-10-1845.pdf;
http://www.alcatel-lucent.
com/bstj/vol45-1966/articles/
bstj45-10-1845.pdf.

Atalla:1960:IRJ

[AT60] M. M. Atalla and E. Tannenbaum. Impurity redistribution and junction formation in silicon by thermal oxidation. The Bell System Technical Journal, 39(4):933–946, July 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-4-933.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-4-933.pdf.

Auld:1965:GOM

[Aul65] B. A. Auld. Geometrical optics of magnetoelastic wave propagation in a nonuniform magnetic field. The Bell System Technical Journal, 44(3):495–507, March 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-3-495.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-3-495.pdf.

Butler:1964:DPC

[BAMS64] J. H. Butler, C. J. Altenburg, R. J. McSweeney, and L. E. Sutton. Design and powering of cable ship ‘long lines’. The Bell System Technical Journal, 43(4):1435–1459, July 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-4-1435.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-4-1435.pdf.

Barnard:1962:VTA

[Bar62] R. D. Barnard. Variational techniques applied to capture in phase-controlled oscillators. The Bell System Technical Journal, 41(1):227–256, January 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol41/bstj41-1-227.
pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-1-227.pdf.

Barnard:1964:DSD

[Bar64a] R. D. Barnard. On the discrete spectral densities of Markov pulse trains. The Bell System Technical Journal, 43(1):233–259, January 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-1-233.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-1-233.pdf.

Barnard:1964:SPS

[Bar64b] R. D. Barnard. On the spectral properties of single-sideband angle-modulated signals. The Bell System Technical Journal, 43(6):2811–2838, November 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-6-2811.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-6-2811.pdf.

Barnard:1965:BJB

[Bar65]  R. D. Barnard.  B.S.T.J.

brief: A note on a special class of one-sided distribution sums. The Bell System Technical Journal, 44(1):203–206, January 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol44/bstj44-1-203.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol44-1965/
                                        articles/bstj44-1-203.pdf.

Bates:1969:ICJ

[Bat69] C. P. Bates. Intermodal coupling at the junction between straight and curved waveguides. The Bell System Technical Journal, 48(7):2259–2280, September 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-7-2259.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-7-2259.pdf.

Bayliss:1968:DMA

[Bay68]  E. T. Bayliss.  Design of

monopulse antenna difference patterns with low sidelobes. The Bell System Technical Journal, 47(5):623–650, May/June 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol47/bstj47-5-623.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol47-1968/
articles/bstj47-5-623.pdf.

Biddulph:1964:LTJ

[BBC^{+}64] R. Biddulph, A. H. Budlong, R. C. Casterline, D. L. Funk, and L. F. Goeller, Jr. Line, trunk, junctor, and service circuits for no. 1 ESS. The Bell System Technical Journal, 43(5):2321–2353, September 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-5-2321.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-5-2321.pdf.

Blecher:1969:CS

[BBHH69] F. H. Blecher, R. C. Boyd, F. J. Hallenbeck, and F. J. Herr. The L-4 coaxial system. The Bell System Technical Journal, 48(4):821–839, April 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-4-821.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-4-821.pdf.

Brown:1963:SRE

[BBM^{+}63] W. L. Brown, T. M. Buck, L. V. Medford, E. W. Thomas, H. K. Gummel, G. L. Miller, and F. M. Smits. The Spacecraft Radiation Experiments. The Bell System Technical Journal, 42(4):899–941, July 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-4-899.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-4-899.pdf.

Bosworth:1969:COB

[BC69] R. H. Bosworth and J. C. Candy. A companded one-bit coder for television transmission. The Bell System Technical Journal, 48(5):1459–1479, May/ June 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-5-1459.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-5-1459.pdf.

Balde:1964:DTF

[BCD64] J. W. Balde, S. S. Charschan, and J. J. Dineen. Deposition of tantalum films with an open-ended vacuum system. The Bell System Technical Journal, 43(1):127–142, January 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-1-127.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-1-127.pdf.

Buck:1968:IBS

[BCDY68] T. M. Buck, H. C. Casey, Jr., J. V. Dalton, and M. Yamin.

Influence of bulk and surface properties on image sensing silicon diode array. The Bell System Technical Journal, 47(9):1827–1854, November 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-9-1827.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-9-1827.pdf.

Breen:1960:SSC

[BD60] C. Breen and C. A. Dahlbom.

Signaling systems for control of telephone switching. The Bell System Technical Journal, 39(6):1381–1444, November 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-6-1381.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-6-1381.pdf.

Balch:1966:EMS

[BDEL66] H. T. Balch, J. C. Dale, T. W.

Eddy, and R. M. Lauver. Estimation of the mean of a stationary random process by periodic sampling. The Bell System Technical Journal, 45(5):733–741, May/ June 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-5-733.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol45-1966/
                                        articles/bstj45-5-733.pdf.

Brewer:1964:RES

[BDV64] S. T. Brewer, F. R. Dickinson, and C. A. Von Roesgen. Repeaters and equalizers for the SD submarine cable system. The Bell System Technical Journal, 43(4):1243–1273, July 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-4-1243.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-4-1243.pdf.

Burckhardt:1969:TTH

[BE69] C. B. Burckhardt and L. H. Enloe. Television transmission of holograms with reduced resolution requirements on the camera tube. The Bell System Technical Journal, 48(5):1529–1535, May/ June 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-5-1529.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-5-1529.pdf.

Beck:1964:BJBb

[Bec64a] A. C. Beck. B.S.T.J. briefs: Gas mixture lens measurements. The Bell System Technical Journal, 43(4):1821–1825, July 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-4-1821.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-4-1821.pdf.

Beck:1964:BJBa

[Bec64b] A. C. Beck. B.S.T.J. briefs:

Thermal gas lens measurements. The Bell System Technical Journal, 43(4):1818–1820, July 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-4-1818.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-4-1818.pdf.

Bangert:1963:SA

[BEH^{+}63] J. T. Bangert, R. S. Engelbrecht, E. T. Harkless, R. V. Sperry, and E. J. Walsh. The Spacecraft Antennas. The Bell System Technical Journal, 42(4):869–897, July 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-4-869.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-4-869.pdf.

Benes:1960:GSP

[Ben60a] V. E. Beneš. General stochastic processes in traffic systems with one server. The Bell System Technical Journal, 39(1):127–160, January 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol39/bstj39-1-127.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol39-1960/
                                        articles/bstj39-1-127.pdf.

Benes:1960:TPT

[Ben60b] V. E. Beneš. Transition probabilities for telephone traffic. The Bell System Technical Journal, 39(5):1297–1320, September 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol39/bstj39-5-1297.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol39-1960/
                                        articles/bstj39-5-1297.pdf.

Benes:1961:ABG

[Ben61a] V. E. Beneš. Asymptotic behavior of general queues with one server. The Bell System Technical Journal, 40(5): 1281–1307, September 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol40/bstj40-5-1281.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol40-1961/
                                        articles/bstj40-5-1281.pdf.

Benes:1961:CFS

[Ben61b] V. E. Beneš. The covariance function of a simple trunk group, with applications to traffic measurement. The Bell System Technical Journal, 40(1):117–148, January 1961. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-1-117.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-1-117.pdf.

Benes:1961:NIE

[Ben61c] V. E. Beneš. A nonlinear integral equation from the theory of servomechanisms. The Bell System Technical Journal, 40(5):1309–1321, September 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-5-1309.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-5-1309.pdf.

Benes:1962:ATP

[Ben62a] V. E. Beneš. Algebraic and topological properties of connecting networks. The Bell System Technical Journal, 41(4):1249–1274, July 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol41/bstj41-4-1249.
pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-4-1249.pdf.

Benes:1962:HRM

[Ben62b] V. E. Beneš. Heuristic remarks and mathematical problems regarding the theory of connecting systems. The Bell System Technical Journal, 41(4):1201–1247, July 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol41/bstj41-4-1201.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol41-1962/
                                        articles/bstj41-4-1201.pdf.

Benes:1962:RTS

[Ben62c] V. E. Beneš. On rearrangeable three-stage connecting networks. The Bell System Technical Journal, 41(5):1481–1492, September 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol41/bstj41-5-1481.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol41-1962/
                                        articles/bstj41-5-1481.pdf.

Benes:1962:UPS

[Ben62d] V. E. Beneš. Ultimately periodic solutions to a nonlinear integro-differential equation. The Bell System Technical Journal, 41(1):257–268, January 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol41/bstj41-1-257.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol41-1962/
                                        articles/bstj41-1-257.pdf.

Benes:1963:MPR

[Ben63a] V. E. Beneš. Markov processes representing traffic in connecting networks. The Bell System Technical Journal, 42(6):2795–2837, November 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-6-2795.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-6-2795.pdf.

Benes:1963:TTT

[Ben63b] V. E. Beneš. A “thermodynamic”

theory of traffic in connecting networks. The Bell System Technical Journal, 42(3):567–607, May 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-3-567.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-3-567.pdf.

Benes:1964:ORM

[Ben64a] V. E. Beneš. Optimal rearrangeable multistage connecting networks. The Bell System Technical Journal, 43(4):1641–1656, July 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-4-1641.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-4-1641.pdf.

Benes:1964:PGC

[Ben64b] V. E. Beneš. Permutation groups, complexes, and rearrangeable connecting networks. The Bell System Technical Journal, 43(4):1619–1640, July 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-4-1619.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-4-1619.pdf.

Benes:1965:IRF

[Ben65a] V. E. Beneš. Index reduction of FM waves by feedback and power-law nonlinearities. The Bell System Technical Journal, 44(4):589–601, April 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol44/bstj44-4-589.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol44-1965/
                                        articles/bstj44-4-589.pdf.

Benes:1965:PRT

[Ben65b] V. E. Beneš. Properties of random traffic in nonblocking telephone connecting networks. The Bell System Technical Journal, 44(3):509–525, March 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol44/bstj44-3-509.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol44-1965/
                                        articles/bstj44-3-509.pdf.

Benes:1965:SIT

[Ben65c] V. E. Beneš. Some inequalities in the theory of telephone traffic.

The Bell System Technical Journal, 44(9):1941–1975, November 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-9-1941.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-9-1941.pdf.

Benes:1966:PCP

[Ben66a] V. E. Beneš. Programming and control problems arising from optimal routing in telephone networks. The Bell System Technical Journal, 45(9):1373–1438, November 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-9-1373.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-9-1373.pdf.

Benes:1966:SEC

[Ben66b] V. E. Beneš. Some examples of comparisons of connecting networks. The Bell System Technical Journal, 45(10): 1829–1835, December 1966. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol45/bstj45-10-1829.pdf;
http://www.alcatel-lucent.
com/bstj/vol45-1966/articles/
bstj45-10-1829.pdf.

Benes:1967:CSP

[Ben67a] V. E. Beneš. Combinatorial solution to the problem of optimal routing in progressive gradings. The Bell System Technical Journal, 46(5):865–881, May/ June 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol46/bstj46-5-865.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-5-865.pdf.

Benes:1967:SPM

[Ben67b] V. E. Beneš. On some proposed models for traffic in connecting networks. The Bell System Technical Journal, 46(1):105–116, January 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol46/bstj46-1-105.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-1-105.pdf.

Benes:1967:ORC

[Ben67c] V. E. Beneš. Optimal routing in connecting networks over finite time intervals. The Bell System Technical Journal, 46(10): 2341–2352, December 1967. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol46/bstj46-10-2341.pdf;
                                        http://www.alcatel-lucent.

com/bstj/vol46-1967/articles/
bstj46-10-2341.pdf.

Berreman:1964:BJBb

[Ber64a] D. W. Berreman. B.S.T.J.

briefs: A gas lens using unlike, counter-flowing gases. The Bell System Technical Journal, 43(4):1476–1479, July 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-4-1476.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-4-1476.pdf.

Berreman:1964:BJBa

[Ber64b] D. W. Berreman. B.S.T.J.

briefs: A lens or light guide using convectively distorted thermal gradients in gases. The Bell System Technical Journal, 43(4):1469–1475, July 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-4-1469.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-4-1469.pdf.

Bergh:1965:AHR

[Ber65a] A. A. Bergh. Atomic hydrogen as a reducing agent. The Bell System Technical Journal, 44(2):261–271, February 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/

                                        images/Vol44/bstj44-2-261.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol44-1965/
                                        articles/bstj44-2-261.pdf.

Berreman:1965:GOR

[Ber65b] D. W. Berreman. Growth of oscillations of a ray about the irregularly wavy axis of a lens light guide. The Bell System Technical Journal, 44(9):2117–2132, November 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol44/bstj44-9-2117.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol44-1965/
                                        articles/bstj44-9-2117.pdf.

Berlekamp:1967:EIS

[Ber67a] E. R. Berlekamp. The enumeration of information symbols in BCH codes. The Bell System Technical Journal, 46(8):1861–1880, October 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol46/bstj46-8-1861.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-8-1861.pdf.

Berlekamp:1967:FPF

[Ber67b] E. R. Berlekamp. Factoring polynomials over finite fields. The Bell System Technical Journal, 46(8):1853–1859, October 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-8-1853.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-8-1853.pdf.

Bakanowski:1960:EPG

[BF60] A. E. Bakanowski and J. H.

Forster. Electrical properties of gold-doped diffused silicon computer diodes. The Bell System Technical Journal, 39(1):87–104, January 1960. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol39/bstj39-1-
87.pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-1-87.pdf.

Beuscher:1969:AMP

[BFH^{+}69] Hugh J. Beuscher, George E.

Fessler, D. Wayne Huffman, Peter J. Kennedy, and Eric Nussbaum. Administration and maintenance plan. The Bell System Technical Journal, 48(8):2765–2815, October 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-8-2765.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-8-2765.pdf.

Bomberger:1963:SPS

[BFT^{+}63] D. C. Bomberger, D. Feldman, D. E. Trucksess, S. J. Brolin, and P. W. Ussery. The Spacecraft Power Supply System. The Bell System Technical Journal, 42(4):943–972, July 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-4-943.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-4-943.pdf.

Bodle:1961:LSP

[BG61a] D. W. Bodle and P. A. Gresh. Lightning surges in paired telephone cable facilities. The Bell System Technical Journal, 40(2):547–576, March 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-2-547.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-2-547.pdf.

Boyd:1961:CMR

[BG61b] G. D. Boyd and J. P. Gordon. Confocal multimode resonator for millimeter through optical wavelength masers. The Bell System Technical Journal, 40(2):489–508, March 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-2-489.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-2-489.pdf.

Bodtmann:1969:BMI

[BG69] W. F. Bodtmann and F. E.

Guilfoyle. Broadband 300 MHz IF amplifier design. The Bell System Technical Journal, 48(6):1665–1686, July/ August 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-6-1665.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-6-1665.pdf.

Bonner:1960:ENI

[BGK60] A. L. Bonner, J. L. Garrison, and W. J. Kopp. The E6 negative impedance repeater. The Bell System Technical Journal, 39(6):1445–1504, November 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-6-1445.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-6-1445.pdf.

Barry:1968:FTT

[BGLS68] J. F. Barry, J. Gammie, N. E. Lentz, and R. C. Salvage. 3A FM terminal transmitter and receiver. The Bell System Technical Journal, 47(7):1423–1458, September 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-7-1423.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-7-1423.pdf.

Brown:1963:RTR

[BGR63] W. L. Brown, J. D. Gabbe, and W. Rosenzweig. Results of the Telstar radiation experiments. The Bell System Technical Journal, 42(4):1505–1560, July 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-4-1505.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-4-1505.pdf.

Blecher:1962:TAC

[BH62a] F. H. Blecher and F. J. Hallenbeck. The transistorized A5 channel bank for broadband systems. The Bell System Technical Journal, 41(1):321–359, January 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol41/bstj41-1-321.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol41-1962/
                                        articles/bstj41-1-321.pdf.

Bowyer:1962:AID

[BH62b] L. R. Bowyer and W. H. Highleyman. An analysis of inherent distortion in asynchronous frequency-shift modulators. The Bell System Technical Journal, 41(6):1695–1736, November 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol41/bstj41-6-1695.
pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-6-1695.pdf.

Brady:1963:ESD

[BH63] P. T. Brady and G. K. Helder.

Echo suppressor design in telephone communications. The Bell System Technical Journal, 42(6):2893–2917, November 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-6-2893.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-6-2893.pdf.

Boyd:1965:NCT

[BH65] R. C. Boyd and F. J. Herr.

The N2 carrier terminal – objectives and analysis. The Bell System Technical Journal, 44(5):731–759, May/June 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-5-731.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-5-731.pdf.

Barton:1969:SP

[BHR69] Marshall E. Barton, Neil M.

Haller, and Guy W. Ricker. Service programs. The Bell System Technical Journal, 48 (8):2865–2896, October 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-8-2865.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-8-2865.pdf.

Brown:1964:ASN

[BHT64] W. S. Brown, J. P. Hyde, and B. A. Tague. The ALPAK system for nonnumerical algebra on a digital computer — II: Rational functions of several variables and truncated power series with rational-function coefficients. The Bell System Technical Journal, 43(2):785–804, March 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-2-785.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-2-785.pdf.

Bhuta:1960:PII

[Bhu60] P. G. Bhuta. Pressure at the interface of inner conductor and dielectric of armorless ocean cable. The Bell System Technical Journal, 39(4):963–972, July 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-4-963.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-4-963.pdf.

Bleisch:1966:NCS

[BI66] G. W. Bleisch and C. W. Irby.

The N3 carrier system: Objectives and transmission features. The Bell System Technical Journal, 45(6):767–799, July/ August 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-6-767.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-6-767.pdf.

Boyd:1962:GCR

[BK62] G. D. Boyd and H. Kogelnik.

Generalized confocal resonator theory. The Bell System Technical Journal, 41(4):1347–1369, July 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol41/bstj41-4-1347.
pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-4-1347.pdf.

Byrne:1963:SJC

[BKR63] C. J. Byrne, B. J. Karafin, and D. B. Robinson, Jr. Systematic jitter in a chain of digital regenerators. The Bell System Technical Journal, 42(6):2679–2714, November 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-6-2679.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-6-2679.pdf.

Bosworth:1968:DSI

[BKRS68] R. H. Bosworth, F. W. Kammerer, D. E. Rowlinson, and J. V. Scattaglia. Design of a simulator for investigating organic synchronization systems. The Bell System Technical Journal, 47(2):209–226, February 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol47/bstj47-2-209.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-2-209.pdf.

Becker:1966:BJBa

[BKS66] F. K. Becker, E. R. Kretzmer, and J. R. Sheehan. B.S.T.J. briefs: A new signal format for efficient data transmission. The Bell System Technical Journal, 45(5):755–758, May/ June 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol45/bstj45-5-755.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol45-1966/
                                        articles/bstj45-5-755.pdf.

Blackman:1964:MOR

[Bla64]  R. B. Blackman.  Methods

of orbit refinement. The Bell System Technical Journal, 43(3):885–909, May 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-3-885.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-3-885.pdf.

Bleickardt:1968:MST

[Ble68]  W. Bleickardt.  Multimod-

ing and its suppression in twisted ring counters. The Bell System Technical Journal, 47(9):2029–2050, November 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-9-2029.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-9-2029.pdf.

Bodmer:1963:STW

[BLOR63] M. G. Bodmer, J. P. Laico, E. G. Olsen, and A. T. Ross. The satellite traveling-wave tube. The Bell System Technical Journal, 42(4):1703–1748, July 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-4-1703.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-4-1703.pdf.

Blue:1969:ALS

[Blu69] J. L. Blue. Approximate largesignal analysis of IMPATT oscillators. The Bell System Technical Journal, 48(2):383–396, February 1969. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-2-383.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-2-383.pdf.

Bomberger:1963:NCC

[BM63] D. C. Bomberger and L. F. Moose. Nickel–cadmium cells for the Spacecraft battery. The Bell System Technical Journal, 42(4):1687–1702, July 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-4-1687.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-4-1687.pdf.

Bishop:1964:CPF

[BM64] J. D. Bishop and S. Mottel. Cable power facility. The Bell System Technical Journal, 43(4):1339–1366, July 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-4-1339.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-4-1339.pdf.

Bishop:1969:PS

[BM69] J. D. Bishop and S. Mottel. Power supplies. The Bell System Technical Journal, 48(4):1041–1064, April 1969. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-4-1041.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-4-1041.pdf.

Brainard:1967:LRTb

[BMP67] R. C. Brainard, F. W. Mounts, and B. Prasada. Low resolution TV: Subjective effects of frame repetition and picture replenishment. The Bell System Technical Journal, 46(1):261–271, January 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-1-261.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-1-261.pdf.

Bowker:1964:DAO

[BNR64] M. W. Bowker, W. G. Nutt, and R. M. Riley. Design of armorless ocean cable. The Bell System Technical Journal, 43(4):1185–1208, July 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-4-1185.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-4-1185.pdf.

See erratum [Ano64l].

Bobeck:1967:PDA

[Bob67] A. H. Bobeck. Properties and device applications of magnetic domains in orthoferrites. The Bell System Technical Journal, 46(8):1901–1925, October 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol46/bstj46-8-1901.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-8-1901.pdf.

Bodtmann:1969:DEB

[Bod69] W. F. Bodtmann. Design of efficient broadband variolossers. The Bell System Technical Journal, 48(6):1687–1702, July/ August 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-6-1687.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-6-1687.pdf.

Bogner:1969:FSF

[Bog69] R. E. Bogner. Frequency sampling filters — Hilbert transformers and resonators. The Bell System Technical Journal, 48(3):501–510, March 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-3-501.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-3-501.pdf.

Borenstein:1963:SCD

[Bor63]  D. P. Borenstein.  Spec-

tral characteristics of digitstimulating speech sounds. The Bell System Technical Journal, 42(6):2839–2847, November 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-6-2839.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-6-2839.pdf.

Bostwick:1962:MTR

[Bos62]  L. G. Bostwick.  A minia-

ture tuned reed selector of high sensitivity and stability. The Bell System Technical Journal, 41(2):411–424, March 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol41/bstj41-2-411.
pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-2-411.pdf.

Browne:1969:CUS

[BQTY69] Thomas E. Browne, Thomas M.

Quinn, Wing N. Toy, and John E. Yates. Control unit system. The Bell System Technical Journal, 48(8):2619–2668, October 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-8-2619.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-8-2619.pdf.

Bennett:1963:SDA

[BR63a] W. R. Bennett and S. O. Rice. Spectral density and autocorrelation functions associated with binary frequency-shift keying. The Bell System Technical Journal, 42(5):2355–2385, September 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-5-2355.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-5-2355.pdf.

Bodtmann:1963:WTI

[BR63b] W. F. Bodtmann and C. L. Ruthroff. A wideband transistor IF amplifier for Space and terrestrial repeaters using grounded-base transformercoupled stages. The Bell System Technical Journal, 42(1):37– 54, January 1963. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol42/bstj42-1-
                                        37.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-1-37.pdf.

Becker:1966:BJBb

[BR66] F. K. Becker and H. R. Rudin. B.S.T.J. briefs: Application of automatic transversal filters to the problem of echo suppression. The Bell System Technical Journal, 45(10):1847–1850, December 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

bell-labs.com/BSTJ/images/
Vol45/bstj45-10-1847.pdf;
http://www.alcatel-lucent.
com/bstj/vol45-1966/articles/
bstj45-10-1847.pdf.

Brady:1965:SBO

[Bra65a] Paul T. Brady. A statistical basis for objective measurement of speech levels. The Bell System Technical Journal, 44(7):1453–1486, September 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-7-1453.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-7-1453.pdf.

Brady:1965:TIP

[Bra65b] Paul T. Brady. A technique for investigating on-off patterns of speech. The Bell System Technical Journal, 44(1):1–22, January 1965. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol44/bstj44-1-
1.pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-1-1.pdf.

Brainard:1967:LRTa

[Bra67]  R. C. Brainard.  Low res-

olution TV: Subjective effects of noise added to a signal. The Bell System Technical Journal, 46(1):233–260, January 1967. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol46/bstj46-1-233.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-1-233.pdf.

Bradford:1968:TWT

[Bra68a] C. E. Bradford. The travelingwave tube amplifier for the microwave transmitter. The Bell System Technical Journal, 47(7):1379–1395, September 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol47/bstj47-7-1379.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-7-1379.pdf.

Brady:1968:SAP

[Bra68b] Paul T. Brady. A statistical analysis of on-off patterns in sixteen conversations. The Bell System Technical Journal, 47(1):73–91, January 1968. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol47/bstj47-1-
                                        73.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-1-73.pdf.

Brady:1969:MGS

[Bra69] Paul T. Brady. A model for generating on-off speech patterns in two-way conversation. The Bell System Technical Journal, 48(7):2445–2472, September 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-7-2445.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-7-2445.pdf.

Brilliant:1966:DFS

[Bri66] M. B. Brilliant. The determination of frequency in systems of mutually synchronized oscillators. The Bell System Technical Journal, 45(10):1737– 1748, December 1966. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol45/bstj45-10-1737.pdf;
http://www.alcatel-lucent.
com/bstj/vol45-1966/articles/
bstj45-10-1737.pdf.

Brilliant:1967:DRS

[Bri67] M. B. Brilliant. Dynamic response of systems of mutually synchronized oscillators. The Bell System Technical Journal, 46(2):319–356, February 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-2-319.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-2-319.pdf.

Briley:1968:SHC

[Bri68]  Bruce E. Briley.  A self-

healing control. The Bell System Technical Journal, 47(10): 2367–2378, December 1968. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol47/bstj47-10-2367.pdf;
http://www.alcatel-lucent.
com/bstj/vol47-1968/articles/
bstj47-10-2367.pdf.

Brown:1963:ASN

[Bro63] W. S. Brown. The ALPAK system for nonnumerical algebra on a digital computer — I: Polynomials in several variables and truncated power series with polynomial coefficients. The Bell System Technical Journal, 42(5):2081–2119, September 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-5-2081.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-5-2081.pdf.

Brown:1967:SMW

[Bro67a] C. B. Brown. A series of miniature wire spring relays. The Bell System Technical Journal, 46(1):117–147, January 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-1-117.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-1-117.pdf.

Brown:1967:LRT

[Bro67b] Earl F. Brown. Low-resolution TV: Subjective comparison of interlaced and noninterlaced pictures. The Bell System Technical Journal, 46(1):199–232, January 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-1-199.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-1-199.pdf.

Brown:1969:SSD

[Bro69]  Earl F. Brown.  A sliding-

scale direct-feedback PCM coder for television. The Bell System Technical Journal, 48(5): 1537–1553, May/June 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-5-1537.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-5-1537.pdf.

Bennett:1963:BDT

[BS63] W. R. Bennett and J. Salz.

Binary data transmission by FM over a real channel. The Bell System Technical Journal, 42(5):2387–2426, September 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-5-2387.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-5-2387.pdf.

Benes:1964:ATD

[BS64] V. E. Beneš and I. W. Sandberg. Applications of a theorem of Dubrovskii to the periodic responses of nonlinear systems. The Bell System Technical Journal, 43(6):2855–2872, November 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-6-2855.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-6-2855.pdf.

Bell:1965:EAP

[BST65] D. T. Bell, L. H. Steiff, and E. R. Taylor. Equipment aspects of packaged N2 carrier system terminals. The Bell System Technical Journal, 44(5):787–822, May/ June 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol44/bstj44-5-787.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol44-1965/
                                        articles/bstj44-5-787.pdf.

Ballentine:1962:SSR

[BSW62] W. E. Ballentine, V. R. Saari, and F. J. Witt. The solid-state receiver in the TL radio system. The Bell System Technical Journal, 41(6):1831–1863, November 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol41/bstj41-6-1831.
pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-6-1831.pdf.

Beach:1963:MPI

[BT63] C. D. Beach and J. M. Trecker.

A method for predicting interchannel modulation due to multipath propagation in FM and PM tropospheric radio systems. The Bell System Technical Journal, 42(1):1–36, January 1963. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol42/bstj42-1-
1.pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-1-1.pdf.

Buchner:1966:CSB

[Buc66] M. M. Buchner, Jr. Computing the spectrum of a binary group code. The Bell System Technical Journal, 45(3):441–449, March 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-3-441.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-3-441.pdf.

Buchner:1967:CND

[Buc67] M. M. Buchner, Jr. Coding for numerical data transmission. The Bell System Technical Journal, 46(5):1025–1041, May/ June 1967. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol46/bstj46-5-1025.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-5-1025.pdf.

Buchner:1969:SAQ

[Buc69a] M. M. Buchner, Jr. A system approach to quantization and transmission error. The Bell System Technical Journal, 48(5):1219–1247, May/ June 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-5-1219.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-5-1219.pdf.

Buchner:1969:ECH

[Buc69b] Morgan M. Buchner, Jr. The equivalence of certain Harper codes. The Bell System Technical Journal, 48(9):3113– 3130, November 1969. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol48/bstj48-9-3113.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol48-1969/articles/
                                        bstj48-9-3113.pdf.

Bugnolo:1965:SEL

[Bug65] D. S. Bugnolo. Some effects of laminar and turbulent flow on breakdown in gases. The Bell System Technical Journal, 44 (10):2393–2408, December 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol44/bstj44-10-2393.pdf;
http://www.alcatel-lucent.
com/bstj/vol44-1965/articles/
bstj44-10-2393.pdf.

Bullington:1963:REA

[Bul63] K. Bullington. Reflections from an exponential atmosphere. The Bell System Technical Journal, 42(6):2849–2867, November 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-6-2849.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-6-2849.pdf.

Bunin:1969:RDF

[Bun69] Barry J. Bunin. Rate-distortion function for Gaussian Markov processes. The Bell System Technical Journal, 48(9): 3059–3074, November 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-9-3059.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-9-3059.pdf.

Burbank:1960:MSC

[Bur60]  R. D. Burbank.  Molecular

structure in crystal aggregates of linear polyethylene. The Bell System Technical Journal, 39(6):1627–1663, November 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-6-1627.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-6-1627.pdf.

Burckhardt:1965:AVF

[Bur65]  C. B. Burckhardt.  Analy-

sis of varactor frequency multipliers for arbitrary capacitance variation and drive level. The Bell System Technical Journal, 44(4):675–692, April 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-4-675.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-4-675.pdf.

Burckhardt:1966:BJB

[Bur66] C. B. Burckhardt. B.S.T.J. briefs: Display of holograms in white light. The Bell System Technical Journal, 45(10): 1841–1844, December 1966. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol45/bstj45-10-1841.pdf;
http://www.alcatel-lucent.
com/bstj/vol45-1966/articles/
bstj45-10-1841.pdf.

Busala:1960:FCD

[Bus60]  A. Busala.  Fundamental

considerations in the design of a voice-switched speakerphone. The Bell System Technical Journal, 39(2):265–294, March 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-2-265.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-2-265.pdf.

Browne:1969:NTD

[BWY69] T. E. Browne, D. J. Wadsworth, and R. K. York. New time division switch units for no. 101 ESS. The Bell System Technical Journal, 48(2):443–476, February 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-2-443.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-2-443.pdf.

Byrne:1962:PDP

[Byr62] C. J. Byrne. Properties and design of the phase-controlled oscillator with a sawtooth comparator. The Bell System Technical Journal, 41(2):559–602, March 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol41/bstj41-2-559.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol41-1962/
                                        articles/bstj41-2-559.pdf.

Claus:1963:ODP

[CBHR63] A. J. Claus, R. B. Blackman, E. G. Halline, and W. C. Ridgway III. Orbit determination and prediction, and computer programs. The Bell System Technical Journal, 42(4):1357–1382, July 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-4-1357.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-4-1357.pdf.

Crowell:1967:BJB

[CBL^{+}67] M. H. Crowell, T. M. Buck, E. F. Labuda, J. V. Dalton, and E. J. Walsh. B.S.T.J. briefs: A camera tube with a silicon diode array target. The Bell System Technical Journal, 46(2):491–495, February 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol46/bstj46-2-491.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-2-491.pdf.

Cravis:1963:ETC

[CC63] H. Cravis and T. V. Crater. Engineering of T1 carrier system repeatered lines. The Bell System Technical Journal, 42(2):431–486, March 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-2-431.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-2-431.pdf.

Curtis:1960:ICD

[CCJ60] H. E. Curtis, T. R. D. Collins, and B. C. Jamison. Interstitial channels for doubling TD-2 radio system capacity. The Bell System Technical Journal, 39(6):1505–1527, November 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-6-1505.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-6-1505.pdf.

Crawford:1963:RBT

[CCKT63] A. B. Crawford, C. C. Cutler, R. Kompfner, and L. C. Tillotson. The research background of the Telstar Experiment. The Bell System Technical Journal, 42(4):747–764, July 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-4-747.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-4-747.pdf.

Chapman:1963:CTS

[CCM63] R. C. Chapman, Jr., G. F.

Critchlow, and H. Mann. Command and telemetry systems. The Bell System Technical Journal, 42(4):1027–1062, July 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-4-1027.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-4-1027.pdf.

Carbaugh:1964:NEC

[CDGH64] D. H. Carbaugh, G. G. Drew, H. Ghiron, and Mrs. E. S. Hoover. No. 1 ESS call processing. The Bell System Technical Journal, 43(5): 2483–2531, September 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-5-2483.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-5-2483.pdf.

Chaplin:1969:SSD

[CDJ69] Norman J. Chaplin, Graydon A. Dodson, and Richard M. Jacobs. Solid state devices. The Bell System Technical Journal, 48(4):983–992, April 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-4-983.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-4-983.pdf.

Clark:1963:PLP

[CDW63] O. P. Clark, E. J. Drazy, and D. C. Weller. A phase-locked primary frequency supply for the L multiplex. The Bell System Technical Journal, 42(2):319–340, March 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-2-319.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-2-319.pdf.

Chevalier:1964:NEC

[CE64] J. G. Chevalier and R. K.

Eisenhart. No. 1 ESS circuit packs and connectors. The Bell System Technical Journal, 43(5):2441–2456, September 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-5-2441.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-5-2441.pdf.

Cook:1965:OCA

[CEZ65] J. S. Cook, E. M. Elam, and H. Zucker. The open Cassegrain antenna: Part I. electromagnetic design and analysis. The Bell System Technical Journal, 44(7):1255–1300, September 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-7-1255.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol44-1965/
                                        articles/bstj44-7-1255.pdf.

Chang:1968:HDT

[CF68] Robert W. Chang and S. L. Freeney. Hybrid digital transmission systems Part I: Joint optimization of analog and digital repeaters. The Bell System Technical Journal, 47(8):1663–1686, October 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol47/bstj47-8-1663.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-8-1663.pdf.

Clemency:1961:FDV

[CG61] W. F. Clemency and W. D. Goodale, Jr. Functional design of a voice-switched speakerphone. The Bell System Technical Journal, 40(3):649–668, May 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol40/bstj40-3-649.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol40-1961/
                                        articles/bstj40-3-649.pdf.

Cohen:1964:BJB

[CG64] M. G. Cohen and E. I. Gordon. B.S.T.J. briefs: Detection of weakly modulated light at microwave frequencies. The Bell System Technical Journal, 43(6):3068–3070, November 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-6-3068.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-6-3068.pdf.

Cohen:1965:ABP

[CG65a] M. G. Cohen and E. I. Gordon. Acoustic beam probing using optical techniques. The Bell System Technical Journal, 44(4):693–721, April 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-4-693.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-4-693.pdf.

Cook:1965:AWE

[CG65b] J. S. Cook and A. J. Giger.

All-weather Earth station satellite communication antennas. The Bell System Technical Journal, 44(7):1225–1228, September 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-7-1225.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-7-1225.pdf.

Cohen:1966:ASL

[CG66] M. G. Cohen and E. I. Gordon. Acoustic scattering of light in a Fabry–Perot resonator. The Bell System Technical Journal, 45(6):945–966, July/ August 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol45/bstj45-6-945.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol45-1966/
                                        articles/bstj45-6-945.pdf.

Chu:1966:ALR

[CH66] T. S. Chu and D. C. Hogg. The attenuation of 3.392 He– Ne laser radiation by methane in the atmosphere. The Bell System Technical Journal, 45(2):301–306, February 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol45/bstj45-2-301.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol45-1966/
                                        articles/bstj45-2-301.pdf.

Chu:1968:EPP

[CH68] T. S. Chu and D. C. Hogg. Effects of precipitation on propagation at 0.63, 3.5, and 10.6 microns. The Bell System Technical Journal, 47(5):723–759, May/ June 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol47/bstj47-5-723.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-5-723.pdf.

Chang:1966:SBL

[Cha66] Robert W. Chang. Synthesis of band-limited orthogonal signals for multichannel data transmission. The Bell System Technical Journal, 45(10): 1775–1796, December 1966. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol45/bstj45-10-1775.pdf;
http://www.alcatel-lucent.
com/bstj/vol45-1966/articles/
bstj45-10-1775.pdf.

Chang:1967:PMS

[Cha67] Robert W. Chang. Precoding for multiple-speed data transmission. The Bell System Technical Journal, 46(7): 1633–1649, September 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-7-1633.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-7-1633.pdf.

Chai:1969:ISE

[Cha69a] S. Y. Chai. An imaging system exhibiting wavelengthdependent resolution. The Bell System Technical Journal, 48(6):1885–1907, July/ August 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-6-1885.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-6-1885.pdf.

Chang:1969:MDS

[Cha69b] Herbert Y. Chang. A method for digitally simulating shorted input diode failures. The Bell System Technical Journal, 48(6):1957–1966, July/ August 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-6-1957.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-6-1957.pdf.

Chen:1964:CTS

[Che64] F. S. Chen. The comb-type slowwave structure for TWM applications. The Bell System Technical Journal, 43(3):1035–1066, May 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-3-1035.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-3-1035.pdf.

Crawford:1961:PEH

[CHH61] A. B. Crawford, D. C. Hogg, and L. E. Hunt. Project Echo: A horn-reflector antenna for space communication. The Bell System Technical Journal, 40(4):1095–1116, July 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol40/bstj40-4-1095.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol40-1961/
articles/bstj40-4-1095.pdf.

Connell:1964:NEB

[CHK64] J. B. Connell, L. W. Hussey, and R. W. Ketchledge. No. 1 ESS bus system. The Bell System Technical Journal, 43(5):2021–2054, September 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-5-2021.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-5-2021.pdf.

Collier:1963:GSH

[CHLS63] R. J. Collier, G. D. Helm, J. P. Laico, and K. M. Striny. The ground station high-power traveling-wave tube. The Bell System Technical Journal, 42(4):1829–1861, July 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-4-1829.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-4-1829.pdf.

Chu:1966:GRG

[Chu66] T. S. Chu. Geometrical representation of Gaussian beam propagation. The Bell System Technical Journal, 45 (2):287–299, February 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol45/bstj45-2-287.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol45-1966/
                                        articles/bstj45-2-287.pdf.

Candy:1968:OSD

[CK68] J. C. Candy and M. Karnaugh. Organic synchronization: Design of the controls and some simulation results. The Bell System Technical Journal, 47(2):227–259, February 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol47/bstj47-2-227.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-2-227.pdf.

Cooney:1968:TE

[CKS68] R. T. Cooney, F. M. Klisch, and C. P. Susen. Test equipment. The Bell System Technical Journal, 47(7):1459–1485, September 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol47/bstj47-7-1459.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-7-1459.pdf.

Cutler:1963:SSA

[CKT63] C. C. Cutler, R. Kompfner, and L. C. Tillotson. A selfsteering array repeater. The Bell System Technical Journal, 42(5):2013–2032, September 1963. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-5-2013.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-5-2013.pdf.

Cochran:1960:NMS

[CL60] W. T. Cochran and D. A.

Lewinski. A new measuring set for message circuit noise. The Bell System Technical Journal, 39(4):911–931, July 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-4-911.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-4-911.pdf.

Cook:1963:AS

[CL63] J. S. Cook and R. Lowell.

The autotrack system. The Bell System Technical Journal, 42(4):1283–1307, July 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-4-1283.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-4-1283.pdf.

Crowell:1969:SDA

[CL69] Merton H. Crowell and Edward F. Labuda. The silicon diode array camera tube. The Bell System Technical Journal, 48(5):1481–1528, May/ June 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-5-1481.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-5-1481.pdf.

Clarke:1968:STM

[Cla68] R. H. Clarke. A statistical theory of mobile-radio reception. The Bell System Technical Journal, 47(6):957–1000, July/ August 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol47/bstj47-6-957.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-6-957.pdf.

Chester:1969:DAC

[CLW69] Arthur N. Chester, Thomas C. Loomis, and M. M. Weiss. Diode array camera tubes and X-ray imaging. The Bell System Technical Journal, 48(2):345–382, February 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-2-345.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-2-345.pdf.

Cooper:1969:QSC

[CM69] R. B. Cooper and G. Murray. Queues served in cyclic order. The Bell System Technical Journal, 48(3):675–689, March 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-3-675.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-3-675.pdf.

Cagle:1964:NEL

[CMS^{+}64] W. B. Cagle, R. S. Menne, R. S. Skinner, R. E. Staehler, and M. D. Underwood. No. 1 ESS logic circuits and their application to the design of the central control. The Bell System Technical Journal, 43(5):2055–2095, September 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-5-2055.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-5-2055.pdf.

Cole:1967:ISW

[CNN67] W. J. Cole, E. R. Nagelberg, and C. M. Nagel. Iterative solution of waveguide discontinuity problems. The Bell System Technical Journal, 46(3):649–672, March 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-3-649.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-3-649.pdf.

Cochran:1965:EEI

[Coc65] J. Alan Cochran. The existence of eigenvalues for the integral equations of laser theory. The Bell System Technical Journal, 44(1):77–88, January 1965. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol44/bstj44-1-
                                        77.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol44-1965/
                                        articles/bstj44-1-77.pdf.

Collinson:1965:SSF

[Col65]  J. A. Collinson.  A sta-

ble, single-frequency RF-excited gas laser at 6328Å. The Bell System Technical Journal, 44(7):1511–1519, September 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol44/bstj44-7-1511.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol44-1965/
                                        articles/bstj44-7-1511.pdf.

Conklin:1966:OGA

[Con66] Glenn E. Conklin. Observed 50 to 60 Gc/s attenuation for the circular electric wave in dielectriccoated cylindrical waveguide bends. The Bell System Technical Journal, 45(5):723–732, May/ June 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol45/bstj45-5-723.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol45-1966/
articles/bstj45-5-723.pdf.

Cooper:1961:DHR

[Coo61] H. G. Cooper. Design of a highresolution electrostatic cathode ray tube for the flying spot store. The Bell System Technical Journal, 40(3):723–759, May 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-3-723.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-3-723.pdf.

Copeland:1967:BJB

[Cop67] John A. Copeland. B.S.T.J.

briefs: CW operation of LSA oscillator diodes — 44 to 88 GHz. The Bell System Technical Journal, 46(1):284–287, January 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-1-284.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-1-284.pdf.

Coyne:1963:MPW

[Coy63]  J. C. Coyne.  Monitoring

the percussive welding process for attaching wires to terminals. The Bell System Technical Journal, 42(1):55–78, January 1963. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/

                                        BSTJ/images/Vol42/bstj42-1-
                                        55.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-1-55.pdf.

Coyne:1965:HTS

[Coy65] K. N. Coyne. Hydrodynamic techniques for study of wind effects on antenna structures. The Bell System Technical Journal, 44(7):1339–1365, September 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol44/bstj44-7-1339.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol44-1965/
                                        articles/bstj44-7-1339.pdf.

Cross:1966:INF

[Cro66] T. G. Cross. Intermodulation noise in FM systems due to transmission deviations and AM/ PM conversation. The Bell System Technical Journal, 45(10): 1749–1773, December 1966. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol45/bstj45-10-1749.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol45-1966/articles/
                                        bstj45-10-1749.pdf.

Cross:1967:PDS

[Cro67] T. G. Cross. Power density spectrum of the sum of two correlated intermodulation noise contributors in FM systems. The Bell System Technical Journal, 46 (10):2437–2452, December 1967.

CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol46/bstj46-10-2437.pdf;
http://www.alcatel-lucent.
com/bstj/vol46-1967/articles/
bstj46-10-2437.pdf.

Chu:1965:GEH

[CS65] T. S. Chu and R. A. Semplak. Gain of electromagnetic horns. The Bell System Technical Journal, 44(3):527–537, March 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-3-527.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-3-527.pdf.

Chen:1964:FFI

[CT64] F. S. Chen and W. J. Tabor.

Filling factor and isolator performance of the traveling-wave maser. The Bell System Technical Journal, 43(3):1005–1033, May 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-3-1005.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-3-1005.pdf.

Chang:1967:MID

[CT67] H. Y. Chang and W. Thomis.

Methods of interpreting diagnostic data for locating faults in digital machines. The Bell System Technical Journal, 46(2):289–317, February 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol46/bstj46-2-289.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-2-289.pdf.

Crawford:1969:PAS

[CT69] A. B. Crawford and R. H. Turrin. A packaged antenna for short-hop microwave radio systems. The Bell System Technical Journal, 48(6):1605–1622, July/ August 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-6-1605.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-6-1605.pdf.

Cohen:1968:LMT

[CUM68] M. I. Cohen, B. A. Unger, and J. F. Milkosky. Laser machining of thin films and integrated circuits. The Bell System Technical Journal, 47(3):385–405, March 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol47/bstj47-3-385.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-3-385.pdf.

Curtis:1960:RFI

[Cur60]  H. E. Curtis.  Radio fre-

quency interference considerations in the TD-2 radio relay system. The Bell System Technical Journal, 39(2):369–387, March 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-2-369.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-2-369.pdf.

Curtis:1962:IBS

[Cur62]  Harold E. Curtis.  Interfer-

ence between satellite communication systems and common carrier surface systems. The Bell System Technical Journal, 41(3):921–943, May 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol41/bstj41-3-921.
pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-3-921.pdf.

Curtis:1963:SSI

[Cur63]  Harold E. Curtis.  Satel-

lite system interference tests at Andover, Maine. The Bell System Technical Journal, 42(6):2715–2739, November 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-6-2715.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-6-2715.pdf.

Cuttriss:1961:RBS

[Cut61]  D. B. Cuttriss.  Relation

between surface concentration and average conductivity in diffused layers in germanium. The Bell System Technical Journal, 40(2):509–521, March 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol40/bstj40-2-509.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol40-1961/
                                        articles/bstj40-2-509.pdf.

DiDomenico:1963:BEO

[DA63] M. DiDomenico, Jr. and L. K. Anderson. Broadband electrooptic traveling-wave light modulators. The Bell System Technical Journal, 42(6):2621– 2678, November 1963. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol42/bstj42-6-2621.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol42-1963/articles/
                                        bstj42-6-2621.pdf.

Daiser:1968:MBD

[Dai68]  P. Daiser.  Measured beam

deformations in a guide made of tubular gas lenses. The Bell System Technical Journal, 47(2):179–194, February 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol47/bstj47-2-179.

pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-2-179.pdf.

Dale:1967:BJB

[Dal67] J. C. Dale. B.S.T.J. briefs:

Estimation of the variance of a stationary Gaussian random process by periodic sampling. The Bell System Technical Journal, 46(6):1283–1287, July/ August 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-6-1283.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-6-1283.pdf.

Daly:1968:LBP

[Dal68] James C. Daly. Linear beam position control in optical waveguides. The Bell System Technical Journal, 47(5):783–799, May/ June 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-5-783.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-5-783.pdf.

Daly:1969:SCM

[Dal69]  J. C. Daly.  Step correc-

tion of misaligned beam waveguides. The Bell System Technical Journal, 48(6):1909–1918, July/ August 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/

                                        images/Vol48/bstj48-6-1909.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-6-1909.pdf.

Darlington:1963:BJB

[Dar63a] Sidney Darlington. B.S.T.J. briefs: A relation between the basis functions of periodically varying nondissipative circuits. The Bell System Technical Journal, 42(6):2970–2971, November 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-6-2970.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-6-2970.pdf.

Darlington:1963:LTV

[Dar63b] Sidney Darlington. Linear timevarying circuits — matrix manipulations, power relations, and some bounds on stability. The Bell System Technical Journal, 42(6):2575–2608, November 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-6-2575.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-6-2575.pdf.

Darlington:1964:DWL

[Dar64] Sidney Darlington. Demodulation of wideband, low-power FM signals. The Bell System Technical Journal, 43(1):339–374, January 1964. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-1-339.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-1-339.pdf.

Davis:1962:EPC

[Dav62] C. G. Davis. An experimental pulse code modulation system for short-haul trunks. The Bell System Technical Journal, 41(1):1– 24, January 1962. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol41/bstj41-1-
1.pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-1-1.pdf.

Davey:1964:DDS

[Dav64]  J. R. Davey.  Digital data

signal space diagrams. The Bell System Technical Journal, 43(6):2973–2983, November 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-6-2973.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-6-2973.pdf.

Dolling:1963:MDH

[DBKW63] J. C. Dolling, R. W. Blackmore, W. J. Kindermann, and K. B. Woodard. The mechanical design of the horn-reflector antenna and radome. The Bell System Technical Journal, 42(4):1137–1186, July 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-4-1137.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-4-1137.pdf.

DeLange:1968:OHE

[DD68] O. E. DeLange and A. F. Dietrich. Optical heterodyne experiments with enclosed transmission paths. The Bell System Technical Journal, 47(2):161–178, February 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol47/bstj47-2-161.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-2-161.pdf.

Danielsen:1964:NES

[DDH64] D. Danielsen, K. S. Dunlap, and H. R. Hofmann. No. 1 ESS switching network framers and circuits. The Bell System Technical Journal, 43(5):2221–2253, September 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-5-2221.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-5-2221.pdf.

DeLange:1961:PES

[De 61]  O. E. De Lange.  Project

Echo: Satellite tracking radar.

The Bell System Technical Journal, 40(4):1157–1182, July 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-4-1157.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-4-1157.pdf.

Deininger:1960:HFE

[Dei60] R. L. Deininger. Human factors engineering studies of the design and use of pushbutton telephone sets. The Bell System Technical Journal, 39(4):995–1012, July 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-4-995.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-4-995.pdf.

DeLange:1965:LSC

[DeL65] O. E. DeLange. Losses suffered by coherent light redirected and refocused many times in an enclosed medium. The Bell System Technical Journal, 44(2):283–302, February 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-2-283.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-2-283.pdf.

Descloux:1963:OPT

[Des63] A. Descloux. On overflow processes of trunk groups with Poisson inputs and exponential service times. The Bell System Technical Journal, 42(2):383–397, March 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-2-383.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-2-383.pdf.

Descloux:1965:ALE

[Des65] A. Descloux. On the accuracy of loss estimates. The Bell System Technical Journal, 44(6):1139–1164, July/ August 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol44/bstj44-6-1139.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol44-1965/
                                        articles/bstj44-6-1139.pdf.

Dewald:1960:CPD

[Dew60]  J. F. Dewald.  The charge

and potential distributions at the zinc oxide electrode. The Bell System Technical Journal, 39(3):615–639, May 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol39/bstj39-3-615.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol39-1960/
                                        articles/bstj39-3-615.pdf.

Digrindakis:1969:PS

[DFHT69] John Digrindakis, Laimons Freimanis, H. Robert Hofmann, and Robert G. Taylor. Peripheral system. The Bell System Technical Journal, 48(8):2669–2712, October 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-8-2669.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-8-2669.pdf.

Dixon:1967:ALM

[DG67] R. W. Dixon and E. I. Gordon. Acoustic light modulators using optical heterodyne mixing. The Bell System Technical Journal, 46(2):367–389, February 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-2-367.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-2-367.pdf.

Denkmann:1965:OCA

[DGPS65] W. J. Denkmann, F. T. Geyling, D. L. Pope, and A. O. Schwarz. The open Cassegrain antenna: Part II. structural and mechanical evaluation. The Bell System Technical Journal, 44(7):1301–1319, September 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-7-1301.

                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol44-1965/
                                        articles/bstj44-7-1301.pdf.

Danylchuk:1968:SSA

[DGS68] I. Danylchuk, U. F. Gianola, and J. T. Sibilia. Step stress aging of plated wire memories. The Bell System Technical Journal, 47(8):1539–1559, October 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol47/bstj47-8-1539.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-8-1539.pdf.

Dragone:1963:WAR

[DH63] C. Dragone and D. C. Hogg. Wide-angle radiation due to rough phase fronts. The Bell System Technical Journal, 42(5):2285–2296, September 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-5-2285.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-5-2285.pdf.

Davis:1963:SCR

[DHWM63] C. G. Davis, P. T. Hutchison, F. J. Witt, and H. I. Maunsell. The Spacecraft Communications Repeater. The Bell System Technical Journal, 42(4):831–867, July 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-4-831.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-4-831.pdf.

Dickieson:1963:TE

[Dic63] A. C. Dickieson. The Telstar Experiment. The Bell System Technical Journal, 42(4):739–746, July 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-4-739.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-4-739.pdf.

Delchamps:1963:STE

[DJS63] T. B. Delchamps, G. C. Jonasson, and R. A. Swift. The Spacecraft Test and Evaluation Program. The Bell System Technical Journal, 42(4):1007–1025, July 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-4-1007.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-4-1007.pdf.

Desoer:1965:NRN

[DK65] Charles A. Desoer and Jacob Katzenelson. Nonlinear RLC networks. The Bell System Technical Journal, 44(1):161–198, January 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/

                                        images/Vol44/bstj44-1-161.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol44-1965/
                                        articles/bstj44-1-161.pdf.

DeGrasse:1961:PED

[DKS61] R. W. De Grasse, J. J. Kostelnick, and H. E. D. Scovil. Project Echo: The dual channel 2390-mc traveling-wave maser. The Bell System Technical Journal, 40(4):1117–1127, July 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol40/bstj40-4-1117.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol40-1961/
                                        articles/bstj40-4-1117.pdf.

Drazy:1968:N

[DMS68] E. J. Drazy, R. C. MacLean, and R. E. Sheehey. Networks. The Bell System Technical Journal, 47(7):1397–1422, September 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol47/bstj47-7-1397.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-7-1397.pdf.

Downing:1964:NEM

[DNT64] R. W. Downing, J. S. Nowak, and L. S. Tuomenoksa. No. 1 ESS maintenance plan. The Bell System Technical Journal, 43(5):1961–2019, September 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-5-1961.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-5-1961.pdf.

Dostis:1965:EDE

[Dos65] I. Dostis. The effects of digital errors on PCM transmission of compandored speech. The Bell System Technical Journal, 44 (10):2227–2243, December 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol44/bstj44-10-2227.pdf;
http://www.alcatel-lucent.
com/bstj/vol44-1965/articles/
bstj44-10-2227.pdf.

Dove:1967:DFT

[Dov67] D. B. Dove. Demagnetizing fields in thin magnetic films. The Bell System Technical Journal, 46(7):1527–1559, September 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-7-1527.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-7-1527.pdf.

Dragone:1967:SRL

[DP67] C. Dragone and V. K. Prabhu.

Scattering relations in lossless varactor frequency multipliers. The Bell System Technical Journal, 46(8):1699–1731, October 1967. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol46/bstj46-8-1699.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-8-1699.pdf.

Dragone:1968:SCS

[DP68] C. Dragone and V. K. Prabhu. Some considerations of stability in lossy varactor harmonic generators. The Bell System Technical Journal, 47(6):887–896, July/ August 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol47/bstj47-6-887.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-6-887.pdf.

Duvall:1969:CCA

[DR69] G. H. Duvall and L. M. Rackson. Coaxial cable and apparatus. The Bell System Technical Journal, 48(4):1065–1093, April 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-4-1065.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-4-1065.pdf.

Dragone:1967:PAMa

[Dra67a] C. Dragone. Phase and amplitude modulation in high efficiency varactor frequency multipliers — general scattering properties. The Bell System Technical Journal, 46(4):775–796, April 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-4-775.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-4-775.pdf.

Dragone:1967:PAMb

[Dra67b] C. Dragone. Phase and amplitude modulation in highefficiency varactor frequency multipliers of order N = 2^{n} — stability and noise. The Bell System Technical Journal, 46(4):797–834, April 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-4-797.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-4-797.pdf.

Dragone:1968:ATS

[Dra68]  Corrado Dragone.  Analysis

of thermal and shot noise in pumped resistive diodes. The Bell System Technical Journal, 47(9):1883–1902, November 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-9-1883.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-9-1883.pdf.

Dragone:1969:APM

[Dra69] C. Dragone. Amplitude and phase modulation in resistive diode mixers. The Bell System Technical Journal, 48(6): 1967–1998, July/August 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-6-1967.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-6-1967.pdf.

Dougherty:1964:NEM

[DRRS64] H. J. Dougherty, H. Raag, P. G. Ridinger, and A. A. Stockert. No. 1 ESS master control center. The Bell System Technical Journal, 43(5):2283–2319, September 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-5-2283.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-5-2283.pdf.

Dorros:1966:EMD

[DSW66] I. Dorros, J. M. Sipress, and F. D. Waldhauer. An experimental 224 Mb/s digital repeatered line. The Bell System Technical Journal, 45(7):993–1043, September 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol45/bstj45-7-993.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol45-1966/
articles/bstj45-7-993.pdf.

Eckler:1960:CMG

[Eck60] A. R. Eckler. The construction of missile guidance codes resistant to random interference. The Bell System Technical Journal, 39(4):973–994, July 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-4-973.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-4-973.pdf.

Ehrbar:1964:SSC

[EFK^{+}64] R. D. Ehrbar, J. M. Fraser, R. A. Kelley, L. H. Morris, E. T. Mottram, and P. W. Rounds. The SD submarine cable system. The Bell System Technical Journal, 43(4):1155–1184, July 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-4-1155.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-4-1155.pdf.

Edson:1965:BCE

[EH65a] J. O. Edson and H. H. Henning.

Broadband codecs for an experimental 224 Mb/s PCM terminal. The Bell System Technical Journal, 44(9):1887–1940, November 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol44/bstj44-9-1887.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol44-1965/
                                        articles/bstj44-9-1887.pdf.

Edwards:1965:NCE

[EH65b] H. S. Edwards and H. Z. Hardaway. New concepts in exchange outside plant engineering. The Bell System Technical Journal, 44(3):373–399, March 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol44/bstj44-3-373.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol44-1965/
                                        articles/bstj44-3-373.pdf.

Ehrbar:1964:CLF

[Ehr64] R. D. Ehrbar. A cable laying facility. The Bell System Technical Journal, 43(4):1367–1372, July 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-4-1367.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-4-1367.pdf.

Enloe:1968:HHS

[EJR68] L. H. Enloe, W. C. Jakes, Jr., and C. B. Rubinstein. Hologram heterodyne scanners. The Bell System Technical Journal, 47(9):1875–1882, November 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-9-1875.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-9-1875.pdf.

Elder:1968:ASS

[Eld68]  H. E. Elder.  Active solid-

state devices. The Bell System Technical Journal, 47(7): 1323–1377, September 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-7-1323.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-7-1323.pdf.

Elliott:1962:HPD

[Ell62] J. S. Elliott. A high-precision direct-reading loss and phase measuring set for carrier frequencies. The Bell System Technical Journal, 41(5): 1493–1517, September 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol41/bstj41-5-1493.
pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-5-1493.pdf.

Elliott:1963:EER

[Ell63]  E. O. Elliott.  Estimates

of error rates for codes on burst-noise channels. The Bell System Technical Journal, 42(5):1977–1997, September 1963. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-5-1977.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-5-1977.pdf.

Elliott:1965:MST

[Ell65] E. O. Elliott. A model of the switched telephone network for data communications. The Bell System Technical Journal, 44(1):89–109, January 1965. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol44/bstj44-1-
                                        89.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol44-1965/
                                        articles/bstj44-1-89.pdf.

Elliott:1966:UPE

[Ell66] E. O. Elliott. On the use and performance of error-voiding and error-marking codes. The Bell System Technical Journal, 45(8):1273–1283, October 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol45/bstj45-8-1273.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol45-1966/
                                        articles/bstj45-8-1273.pdf.

Emling:1963:ETD

[EM63] J. W. Emling and D. Mitchell. The effects of time delay and echoes on telephone conversations. The Bell System Technical Journal, 42(6): 2869–2891, November 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-6-2869.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-6-2869.pdf.

Enloe:1966:BJB

[EMR66] L. H. Enloe, J. A. Murphy, and C. B. Rubinstein. B.S.T.J. briefs: Hologram transmission via television. The Bell System Technical Journal, 45(2):335–339, February 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-2-335.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-2-335.pdf.

Evens:1963:CGT

[EMT63] M. J. Evens, G. H. Myers, and J. W. Timko. Command guidance of Telstar launch vehicle. The Bell System Technical Journal, 42(5):2153–2168, September 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-5-2153.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-5-2153.pdf.

Ebert:1969:OOD

[EMT69] P. M. Ebert, James E. Maxo, and Michael G. Taylor. Overflow oscillations in digital filters. The Bell System Technical Journal, 48(9):2999–3020, November 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-9-2999.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-9-2999.pdf.

Engel:1965:DTP

[Eng65] J. S. Engel. Digital transmission in the presence of impulsive noise. The Bell System Technical Journal, 44(8):1699–1743, October 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-8-1699.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-8-1699.pdf.

Enloe:1963:IIF

[Enl63]  L. H. Enloe.  Interchannel

interference in FM systems produced by continuous random mode conversion in circular electric waveguide. The Bell System Technical Journal, 42(6):2765–2786, November 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-6-2765.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-6-2765.pdf.

Enloe:1967:NLS

[Enl67] L. H. Enloe. Noise-like structure in the image of diffusely reflecting objects in coherent illumination. The Bell System Technical Journal, 46(7):1479–1489, September 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-7-1479.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-7-1479.pdf.

Enloe:1969:BJB

[Enl69] L. H. Enloe. B.S.T.J. briefs:

Correction concerning reflecting objects in coherent illumination. The Bell System Technical Journal, 48(3):817, March 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-3-817.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-3-817.pdf.

Ebert:1969:CRS

[ET69] P. M. Ebert and S. Y. Tong.

Convolutional Reed–Solomon codes. The Bell System Technical Journal, 48(3):729–742, March 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-3-729.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-3-729.pdf.

Evans:1969:MFC

[Eva69] James G. Evans. Measuring frequency characteristics of linear two-port networks automatically. The Bell System Technical Journal, 48(5):1313–1338, May/ June 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-5-1313.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-5-1313.pdf.

Froehlich:1964:DTS

[FA64] F. E. Froehlich and R. R. Anderson. Data transmission over a self-contained error detection and retransmission channel. The Bell System Technical Journal, 43(1):375–398, January 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-1-375.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-1-375.pdf.

Falconer:1969:HCS

[Fal69] D. D. Falconer. A hybrid coding scheme for discrete memoryless channels. The Bell System Technical Journal, 48(3):691–728, March 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-3-691.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol48-1969/
articles/bstj48-3-691.pdf.

Fraser:1962:ACT

[FBL62] J. M. Fraser, D. B. Bullock, and N. G. Long. Over-all characteristics of a TASI system. The Bell System Technical Journal, 41(4):1439–1454, July 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol41/bstj41-4-1439.
pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-4-1439.pdf.

Freeny:1968:HDT

[FC68] S. L. Freeny and Robert W.

Chang. Hybrid digital transmission systems Part II: Information rate of hybrid coaxial cable systems. The Bell System Technical Journal, 47(8):1687–1711, October 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-8-1687.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-8-1687.pdf.

Feiner:1964:F

[Fei64] A. Feiner. The ferreed. The Bell System Technical Journal, 43(1):1–14, January 1964. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol43/bstj43-1-1.pdf;   http:

                                        //www.alcatel-lucent.com/bstj/
                                        vol43-1964/articles/bstj43-
                                        1-1.pdf.

Fennick:1969:ADT

[Fen69] J. H. Fennick. Amplitude distributions of television channel noise and a model for impulse noise. The Bell System Technical Journal, 48(10):3243– 3263, December 1969. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol48/bstj48-10-3243.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol48-1969/articles/
                                        bstj48-10-3243.pdf.

Flanagan:1966:PV

[FG66] J. L. Flanagan and R. M. Golden. Phase vocoder. The Bell System Technical Journal, 45(9):1493–1509, November 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol45/bstj45-9-1493.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol45-1966/
                                        articles/bstj45-9-1493.pdf.

Freeny:1969:SDI

[FG69] Mrs. A. E. Freeny and J. D. Gabbe. A statistical description of intense rainfall. The Bell System Technical Journal, 48(6):1789–1851, July/ August 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-6-1789.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-6-1789.pdf.

Ferguson:1964:NEA

[FGK^{+}64] J. G. Ferguson, W. E. Grutzner, D. C. Koehler, R. S. Skinner, M. T. Skubiak, and D. H. Wetherell. No. 1 ESS apparatus and equipment. The Bell System Technical Journal, 43(5):2355–2439, September 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-5-2355.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-5-2355.pdf.

Freimanis:1964:NES

[FGM64] L. Freimanis, A. M. Guercio, and H. F. May. No. 1 ESS scanner, signal distributor, and central pulse distributor. The Bell System Technical Journal, 43(5):2255–2282, September 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-5-2255.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-5-2255.pdf.

Farber:1966:SPL

[FGP66] D. J. Farber, R. E. Griswold, and I. P. Polonsky. The SNOBOL3 programming language. The Bell System Technical Journal, 45(6):895–944, July/ August 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-6-895.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-6-895.pdf.

Frisch:1962:MCS

[FGVH62] H. L. Frisch, S. B. Gordon, V. A. Vyssotsky, and J. M. Hammersley. Monte Carlo solution of bond percolation processes in various crystal lattices. The Bell System Technical Journal, 41(3):909–920, May 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol41/bstj41-3-909.
pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-3-909.pdf.

Feiner:1964:NES

[FH64] A. Feiner and W. S. Hayward. No. 1 ESS switching network plan. The Bell System Technical Journal, 43 (5):2193–2220, September 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-5-2193.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-5-2193.pdf.

Fraser:1967:IHF

[FHS67] J. M. Fraser, H. H. Haas, and M. G. Schachtman. An improved high-frequency radiotelephone system featuring constant net loss operation. The Bell System Technical Journal, 46(4):677–720, April 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-4-677.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-4-677.pdf.

Friis:1966:TTS

[FJJK66] R. W. Friis, J. J. Jansen, R. M.

Jensen, and H. T. King. The TM-1/TL-2 short haul microwave systems. The Bell System Technical Journal, 45(1):1–95, January 1966. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol45/bstj45-1-
1.pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-1-1.pdf.

Fenderson:1968:AIU

[FJL68] G. L. Fenderson, J. J. Jansen, and S. H. Lee. Active IF units for the tranmitter and receiver. The Bell System Technical Journal, 47(7):1227–1256, September 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-7-1227.

                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-7-1227.pdf.

Fox:1961:RMM

[FL61] A. G. Fox and Tingye Li. Resonant modes in a maser interferometer. The Bell System Technical Journal, 40(2):453–488, March 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol40/bstj40-2-453.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol40-1961/
                                        articles/bstj40-2-453.pdf.

Foxhall:1964:RIS

[FL64] G. F. Foxhall and J. A. Lewis. The resistance of an infinite slab with a disk electrode. The Bell System Technical Journal, 43(4):1609–1617, July 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-4-1609.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-4-1609.pdf.

Flanagan:1960:MAB

[Fla60] J. L. Flanagan. Models for approximating basilar membrane displacement. The Bell System Technical Journal, 39(5):1163–1191, September 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/

images/Vol39/bstj39-5-1163.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-5-1163.pdf.

Flanagan:1962:MAB

[Fla62] James L. Flanagan. Models for approximating basilar membrane displacement — Part II. effects of middle-ear transmission and some relations between subjective and physiological behavior. The Bell System Technical Journal, 41(3):959–1009, May 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol41/bstj41-3-959.
pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-3-959.pdf.

Feiner:1960:FNS

[FLLR60] A. Feiner, C. A. Lovell, T. N.

Lowry, and P. G. Ridinger. The ferreed — A new switching device. The Bell System Technical Journal, 39(1):1–30, January 1960. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol39/bstj39-1-
1.pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-1-1.pdf.

Forys:1969:DCN

[For69] Leonard J. Forys. The determinability of classes of noisy channels. The Bell System Technical Journal, 48(10):3265– 3285, December 1969. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol48/bstj48-10-3265.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol48-1969/articles/
                                        bstj48-10-3265.pdf.

Fultz:1965:TCS

[FP65] K. E. Fultz and D. B. Penick. The T1 carrier system. The Bell System Technical Journal, 44(7):1405–1451, September 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol44/bstj44-7-1405.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol44-1965/
                                        articles/bstj44-7-1405.pdf.

Fritsch:1969:QCG

[FP69] C. A. Fritsch and D. J. Prager. A quadratic conduction gas lens. The Bell System Technical Journal, 48(7):2281–2303, September 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-7-2281.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-7-2281.pdf.

Franks:1966:MRV

[Fra66]  L. E. Franks.  A model

for the random video process. The Bell System Technical Journal, 45(4):609–630, April 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-4-609.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-4-609.pdf.

Franaszek:1968:SSC

[Fra68] P. A. Franaszek. Sequencestate coding for digital transmission. The Bell System Technical Journal, 47(1):143–157, January 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-1-143.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-1-143.pdf.

Freeny:1969:STR

[Fre69] Mrs. A. E. Freeny. Statistical treatment of rain gauge calibration data. The Bell System Technical Journal, 48(6): 1757–1766, July/August 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-6-1757.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-6-1757.pdf.

Fritsch:1968:EAT

[Fri68] C. A. Fritsch. Effects associated with the thermal response of the T1 telephone transmitter. The Bell System Technical Journal, 47(8):1615–1636, October 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-8-1615.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-8-1615.pdf.

Fletcher:1963:DAT

[FRY63] H. J. Fletcher, L. Rongved, and E. Y. Yu. Dynamics analysis of a two-body gravitationally oriented satellite. The Bell System Technical Journal, 42(5):2239–2266, September 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-5-2239.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-5-2239.pdf.

Franks:1960:AAR

[FS60] L. E. Franks and I. W. Sandberg. An alternative approach to the realization of network transfer functions: The N-path filter. The Bell System Technical Journal, 39(5):1321–1350, September 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-5-1321.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-5-1321.pdf.

Graham:1963:NGS

[GAPB63] R. S. Graham, W. E. Adams, R. E. Powers, and F. R. Bies. New group and supergroup terminals for L multiplex. The Bell System Technical Journal, 42(2):223–278, March 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-2-223.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-2-223.pdf.

Gardner:1960:STD

[Gar60] Mark B. Gardner. A study of talking distance and related parameters in hands-free telephony. The Bell System Technical Journal, 39(6):1529–1551, November 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-6-1529.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-6-1529.pdf.

Giguere:1966:EDN

[GBW66] W. J. Giguere, L. A. Bowie, and J. A. Watters. Equipment design of N3 carrier terminals and carrier supply. The Bell System Technical Journal, 45(6):845–894, July/August 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-6-845.

                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol45-1966/
                                        articles/bstj45-6-845.pdf.

Giger:1963:FDN

[GC63] A. J. Giger and J. C. Chaffee. The FM demodulator with negative feedback. The Bell System Technical Journal, 42(4):1109–1135, July 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-4-1109.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-4-1109.pdf.

Gordon:1968:CST

[GC68] Eugene I. Gordon and Merton H. Crowell. A charge storage target for electron image sensing. The Bell System Technical Journal, 47(9):1855–1873, November 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol47/bstj47-9-1855.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-9-1855.pdf.

Gloge:1969:SDM

[GCE69] D. Gloge, E. L. Chinnock, and H. E. Earl. Scattering from dielectric mirrors. The Bell System Technical Journal, 48(3):511–526, March 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-3-511.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-3-511.pdf.

Gentleman:1968:MMF

[Gen68] W. Morven Gentleman. Matrix multiplication and Fast Fourier Transforms. The Bell System Technical Journal, 47(6):1099–1103, July/ August 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-6-1099.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-6-1099.pdf.

Gershenzon:1966:SAG

[Ger66]  M. Gershenzon.  State of

the art in GaP electroluminescent junctions. The Bell System Technical Journal, 45(9):1599–1609, November 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-9-1599.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-9-1599.pdf.

Gersho:1969:AEH

[Ger69a] Allen Gersho. Adaptive equalization of highly dispersive channels for data transmission. The Bell System Technical Journal, 48(1):55–70, January 1969. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol48/bstj48-1-
                                        55.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-1-55.pdf.

Gersho:1969:BJB

[Ger69b] Allen Gersho. B.S.T.J. briefs: Solving nonlinear network equations using optimization techniques. The Bell System Technical Journal, 48(9):3135– 3138, November 1969. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol48/bstj48-9-3135.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol48-1969/articles/
                                        bstj48-9-3135.pdf.

Geyling:1964:PMS

[Gey64]  F. T. Geyling.  Perturba-

tion methods for satellite orbits. The Bell System Technical Journal, 43(3):847–884, May 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-3-847.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-3-847.pdf.

Geyling:1960:SST

[GF60] F. T. Geyling and J. J. Forst. Semiconductor strain transducers. The Bell System Technical Journal, 39(3):705–731, May 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-3-705.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-3-705.pdf.

Gammie:1960:TRR

[GH60] J. Gammie and S. D. Hathaway. The TJ radio relay system. The Bell System Technical Journal, 39(4):821–877, July 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-4-821.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-4-821.pdf.

Gay:1961:PST

[GHS61] R. R. Gay, B. H. Hamilton, and H. H. Spencer. Power systems for the TH radio system. The Bell System Technical Journal, 40(6):1627–1646, November 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-6-1627.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-6-1627.pdf.

Genke:1964:NEC

[GHS64] R. M. Genke, P. A. Harding, and R. E. Staehler. No. 1 ESS call store — A 0.2megabit ferrite sheet memory.

The Bell System Technical Journal, 43(5):2147–2191, September 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-5-2147.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-5-2147.pdf.

Geldart:1969:HMC

[GHS69] W. J. Geldart, G. D. Haynie, and R. G. Schleich. A 50 Hz–250 MHz computer-operated transmission measuring set. The Bell System Technical Journal, 48(5):1339–1381, May/ June 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-5-1339.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-5-1339.pdf.

Gianola:1960:IMC

[Gia60] U. F. Gianola. Integrated magnetic circuits for synchronous sequential logic machines. The Bell System Technical Journal, 39(2):295–332, March 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-2-295.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-2-295.pdf.

Gibby:1960:EDS

[Gib60]  R. A. Gibby.  An evalua-

tion of AM data system performance by computer simulation. The Bell System Technical Journal, 39(3):675–704, May 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-3-675.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-3-675.pdf.

Giger:1965:BBG

[Gig65]  A. J. Giger.  BSTJ brief:

4-gc transmission degradation due to rain at the Andover, Maine, satellite station. The Bell System Technical Journal, 44(7):1528–1533, September 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-7-1528.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-7-1528.pdf.

Gilbert:1960:CBN

[Gil60] E. N. Gilbert. Capacity of a burst-noise channel. The Bell System Technical Journal, 39(5):1253–1265, September 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-5-1253.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol39-1960/
                                        articles/bstj39-5-1253.pdf.

Gilbert:1965:ERM

[Gil65] E. N. Gilbert. Energy reception for mobile radio. The Bell System Technical Journal, 44(8):1779–1803, October 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol44/bstj44-8-1779.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol44-1965/
                                        articles/bstj44-8-1779.pdf.

Gilbert:1967:MCC

[Gil67] E. N. Gilbert. Minimum cost communication networks. The Bell System Technical Journal, 46(9):2209–2227, November 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol46/bstj46-9-2209.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-9-2209.pdf.

Gilbert:1969:MRD

[Gil69]  E. N. Gilbert.  Mobile ra-

dio diversity reception. The Bell System Technical Journal, 48(7):2473–2492, September 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-7-2473.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol48-1969/
articles/bstj48-7-2473.pdf.

Geller:1962:XRD

[GK62] S. Geller and H. Katz. An X-ray diffraction study of the structure of guanidinium aluminum sulfate hexahydrate. The Bell System Technical Journal, 41(2):425–452, March 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol41/bstj41-2-425.
pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-2-425.pdf.

Golden:1964:DWS

[GK64a] R. M. Golden and J. F. Kaiser.

Design of wideband sampled-data filters. The Bell System Technical Journal, 43(4):1533–1546, July 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-4-1533.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-4-1533.pdf.

Gordon:1964:ERA

[GK64b] J. P. Gordon and H. Kogelnik. Equivalence relations among spherical mirror optical resonators. The Bell System Technical Journal, 43(6):2873–2886, November 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-6-2873.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-6-2873.pdf.

Gersho:1966:MSG

[GK66] A. Gersho and B. J. Karafin. Mutual synchronization of geographically separated oscillators. The Bell System Technical Journal, 45(10):1689–1704, December 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

                                        bell-labs.com/BSTJ/images/
                                        Vol45/bstj45-10-1689.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol45-1966/articles/
                                        bstj45-10-1689.pdf.

Githens:1963:APS

[GKLL63] J. A. Githens, H. P. Kelly, J. C. Lozier, and A. A. Lundstrom. Antenna pointing system: Organization and performance. The Bell System Technical Journal, 42(4):1213–1221, July 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-4-1213.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-4-1213.pdf.

Giger:1961:APS

[GL61] A. J. Giger and F. K. Low. The automatic protection switching system of TH radio. The Bell System Technical Journal, 40(6):1665–1715, November 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-6-1665.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-6-1665.pdf.

Gordon:1964:BJBb

[GL64] E. I. Gordon and E. F. Labuda.

B.S.T.J. briefs: Gas pumping in continuously operated ion lasers. The Bell System Technical Journal, 43(4):1827–1829, July 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-4-1827.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-4-1827.pdf.

Glass:1967:PDM

[Gla67] M. S. Glass. Principles of design of magnetic devices for attitude control of satellites. The Bell System Technical Journal, 46(5):893–912, May/June 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-5-893.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-5-893.pdf.

Gloge:1967:DGL

[Glo67a] D. Gloge. Deformation of gas lenses by gravity. The Bell System Technical Journal, 46(2):357–365, February 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol46/bstj46-2-357.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-2-357.pdf.

Gloge:1967:EUL

[Glo67b] D. Gloge. Experiments with an underground lens waveguide. The Bell System Technical Journal, 46(4):721–735, April 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol46/bstj46-4-721.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-4-721.pdf.

Gloge:1967:MCL

[Glo67c] D. Gloge. Mode conversion in lens guides with imperfect lenses. The Bell System Technical Journal, 46(10):2467–2484, December 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

                                        bell-labs.com/BSTJ/images/
                                        Vol46/bstj46-10-2467.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol46-1967/articles/
                                        bstj46-10-2467.pdf.

Garrison:1969:BRR

[GLR69] J. L. Garrison, L. P. Labbe, and C. C. Rock. Basic and regulating repeaters. The Bell System Technical Journal, 48(4):841–888, April 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-4-841.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-4-841.pdf.

Graham:1966:NIS

[GM66] R. L. Graham and Jessie MacWilliams. On the number of information symbols in difference-set cyclic codes. The Bell System Technical Journal, 45(7):1057–1070, September 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-7-1057.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-7-1057.pdf.

Gewartowski:1967:LSC

[GM67] J. W. Gewartowski and R. H.

Minetti. Large-signal calculations for the overdriven varactor upper-sideband upconverter operating on maximum power output. The Bell System Technical Journal, 46(6):1223–1242, July/ August 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-6-1223.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-6-1223.pdf.

Griffiths:1965:PSS

[GN65] H. D. Griffiths and J. Nedelka. 100Å protection switching system. The Bell System Technical Journal, 44(10):2295– 2336, December 1965. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol44/bstj44-10-2295.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol44-1965/articles/
                                        bstj44-10-2295.pdf.

Goetzberger:1967:TDI

[GN67] A. Goetzberger and E. H. Nicollian. Temperature dependence of inversion-layer frequency response in silicon. The Bell System Technical Journal, 46(3):513–522, March 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol46/bstj46-3-513.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-3-513.pdf.

Goetzberger:1966:IMC

[Goe66] A. Goetzberger. Ideal MOS curves for silicon. The Bell System Technical Journal, 45 (7):1097–1122, September 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol45/bstj45-7-1097.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol45-1966/
articles/bstj45-7-1097.pdf.

Goell:1969:CHC

[Goe69] J. E. Goell. A circular-harmonic computer analysis of rectangular dielectric waveguide. The Bell System Technical Journal, 48(7):2133–2160, September 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-7-2133.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-7-2133.pdf.

Goldstein:1962:APC

[Gol62]  A. Jay Goldstein.  Anal-

ysis of the phase-controlled loop with a sawtooth comparator. The Bell System Technical Journal, 41(2):603–633, March 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol41/bstj41-2-603.
pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-2-603.pdf.

Golden:1966:DCS

[Gol66] Roger M. Golden. Digital computer simulation of sampleddata communication systems using the block diagram compiler: BLODIB. The Bell System Technical Journal, 45(3):345–358, March 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol45/bstj45-3-345.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol45-1966/
                                        articles/bstj45-3-345.pdf.

Goodman:1969:ADM

[Goo69a] David J. Goodman. The application of delta modulation to analog-to-PCM encoding. The Bell System Technical Journal, 48(2):321–343, February 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-2-321.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-2-321.pdf.

Goodman:1969:DMG

[Goo69b] David J. Goodman. Delta modulation granular quantizing noise. The Bell System Technical Journal, 48(5):1197–1218, May/ June 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-5-1197.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-5-1197.pdf.

Gopinath:1969:ILT

[Gop69] B. Gopinath. On the identification of linear time-invariant systems from input-output data. The Bell System Technical Journal, 48(5):1101–1113, May/ June 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-5-1101.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-5-1101.pdf.

Gordon:1960:TEB

[Gor60]  E. I. Gordon.  Transverse

electron beam waves in varying magnetic fields. The Bell System Technical Journal, 39(6):1603–1616, November 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-6-1603.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-6-1603.pdf.

Gordon:1964:OMO

[Gor64a] Eugene I. Gordon. Optical maser oscillators and noise. The Bell System Technical Journal, 43(1):507–539, January 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-1-507.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-1-507.pdf.

Gordon:1964:BJBa

[Gor64b] J. P. Gordon. B.S.T.J. briefs:

A circle diagram for optical resonators. The Bell System Technical Journal, 43(4):1826–1827, July 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-4-1826.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-4-1826.pdf.

Gordon:1966:OGG

[Gor66]  J. P. Gordon.  Optics of

general guilding media. The Bell System Technical Journal, 45(2):321–332, February 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol45/bstj45-2-321.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol45-1966/
                                        articles/bstj45-2-321.pdf.

Gilbert:1960:ADS

[GP60] E. N. Gilbert and H. O. Pollak. Amplitude distribution of shot noise. The Bell System Technical Journal, 39(2):333–350, March 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol39/bstj39-2-333.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol39-1960/
                                        articles/bstj39-2-333.pdf.

Githens:1963:DEA

[GP63] J. A. Githens and T. R. Peters. Digital equipment for the antenna pointing system. The Bell System Technical Journal, 42(4):1223–1252, July 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-4-1223.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-4-1223.pdf.

Gray:1964:UDS

[GP64] J. R. Gray and J. W. Pan.

Using digit statistics to wordframe PCM signals. The Bell System Technical Journal, 43(6):2985–3007, November 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-6-2985.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-6-2985.pdf.

Giger:1963:GTR

[GPW63] A. J. Giger, S. Pardee, Jr., and P. R. Wickliffe, Jr. The ground transmitter and receiver. The Bell System Technical Journal, 42(4):1063–1107, July 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-4-1063.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-4-1063.pdf.

Graff:1963:DSW

[GPZ63] H. J. Graff, J. M. Peacock, and J. J. Zalmans. Development of solderless wire connector for splicing multipair cable. The Bell System Technical Journal, 42(1):131–153, January 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-1-131.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-1-131.pdf.

Gordon:1963:FPE

[GR63] E. I. Gordon and J. D. Rigden. The Fabry–Perot electrooptic modulator. The Bell System Technical Journal, 42(1):155–179, January 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-1-155.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-1-155.pdf.

Gardner:1967:GDS

[GR67] K. R. Gardner and T. R. Robillard. Gold doped silicon compandor diodes for N2 and N3 carrier systems. The Bell System Technical Journal, 46(7):1451–1477, September 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol46/bstj46-7-1451.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-7-1451.pdf.

Graham:1965:DLP

[Gra65a] R. L. Graham. On the decomposition of lattice-periodic functions. The Bell System Technical Journal, 44(6):1191–1214, July/ August 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-6-1191.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-6-1191.pdf.

Grant:1965:CLSa

[Gra65b] W. J. C. Grant. Computation of lattice sums: Generalization of the Ewald method. The Bell System Technical Journal, 44(3):427–437, March 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-3-427.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-3-427.pdf.

Grant:1965:CLSb

[Gra65c] Walter J. C. Grant. Computation of lattice sums: Generalization of the Ewald method II. The Bell System Technical Journal, 44 (10):2433–2446, December 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol44/bstj44-10-2433.pdf;
http://www.alcatel-lucent.
com/bstj/vol44-1965/articles/
bstj44-10-2433.pdf.

Graham:1966:BCM

[Gra66] R. L. Graham. Bounds for certain multiprocessing anomalies. The Bell System Technical Journal, 45(9):1563–1581, November 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol45/bstj45-9-1563.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol45-1966/
                                        articles/bstj45-9-1563.pdf.

Gretter:1964:CPS

[Gre64] R. W. Gretter. Cable payout system. The Bell System Technical Journal, 43(4):1395–1434, July 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-4-1395.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-4-1395.pdf.

Gresh:1969:PTC

[Gre69] Philip A. Gresh. Physical and transmission characteristics of customer loop plant. The Bell System Technical Journal, 48 (10):3337–3385, December 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol48/bstj48-10-3337.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol48-1969/articles/
                                        bstj48-10-3337.pdf.

Grismore:1964:CRH

[Gri64]  O. D. Grismore.  Cable

and repeater handling system. The Bell System Technical Journal, 43(4):1373–1394, July 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-4-1373.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-4-1373.pdf.

Gummel:1961:MCE

[GS61] H. K. Gummel and F. M.

Smits. Margin considerations for an esaki diode resistor OR gate. The Bell System Technical Journal, 40(1):213–232, January 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-1-213.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-1-213.pdf.

Geusic:1962:UTW

[GS62] J. E. Geusic and H. E. D. Scovil.

A unidirectional traveling-wave optical maser. The Bell System Technical Journal, 41(4): 1371–1397, July 1962. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol41/bstj41-4-1371.pdf;
http://www.alcatel-lucent.
com/bstj/vol41-1962/articles/
bstj41-4-1371.pdf.

Grantges:1964:NGP

[GS64a] R. F. Grantges and N. R. Sinowitz. NEASIM: A generalpurpose computer simulation program for load-loss analysis of multistage central office switching networks. The Bell System Technical Journal, 43(3):965–1004, May 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-3-965.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-3-965.pdf.

Gummel:1964:ESC

[GS64b] H. K. Gummel and F. M. Smits. Evaluation of solar cells by means of spectral analysis. The Bell System Technical Journal, 43(3):1103–1121, May 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-3-1103.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-3-1103.pdf.

Gibby:1965:SEN

[GS65a] R. A. Gibby and J. W. Smith. Some extensions of Nyquist’s telegraph transmission theory. The Bell System Technical Journal, 44(7):1487–1510, September 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/

images/Vol44/bstj44-7-1487.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-7-1487.pdf.

Grau:1965:CMM

[GS65b] T. G. Grau and A. K. Spiegler.

Core materials for magnetic latching wire spring relays. The Bell System Technical Journal, 44(6):999–1017, July/August 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-6-999.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-6-999.pdf.

Gummel:1966:ARI

[GS66] H. K. Gummel and D. L. Scharfetter. Avalanche region of IM-PATT diodes. The Bell System Technical Journal, 45(10): 1797–1827, December 1966. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol45/bstj45-10-1797.pdf;
http://www.alcatel-lucent.
com/bstj/vol45-1966/articles/
bstj45-10-1797.pdf.

Gloge:1968:PSH

[GS68] D. Gloge and W. H. Steier.

Pulse shuttling in a half-mile optical lens guideI. The Bell System Technical Journal, 47(5):767–782, May/June 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol47/bstj47-5-767.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-5-767.pdf.

Gloge:1969:ESM

[GS69a] D. Gloge and W. H. Steier. Experimental simulation of a multiple beam optical waveguide. The Bell System Technical Journal, 48(5):1445–1457, May/ June 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-5-1445.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-5-1445.pdf.

Goell:1969:BJB

[GS69b] J. E. Goell and R. D. Standley. B.S.T.J. briefs: Sputtered glass waveguide for integrated optical circuits. The Bell System Technical Journal, 48(10): 3445–3448, December 1969. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol48/bstj48-10-3445.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol48-1969/articles/
                                        bstj48-10-3445.pdf.

Giger:1965:TFH

[GT65] A. J. Giger and R. H. Turrin. The triply-folded horn reflector: A compact ground station antenna design for satellite communications. The Bell System Technical Journal, 44(7):1229–1253, September 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-7-1229.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-7-1229.pdf.

Gucker:1962:LTF

[Guc62]  George B. Gucker.  Long-

term frequency stability for a reflex klystron without the use of external cavities. The Bell System Technical Journal, 41(3):945–958, May 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol41/bstj41-3-945.
pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-3-945.pdf.

Galindo:1968:DLC

[GW68a] V. Galindo and C. P. Wu. Dielectric loaded and covered rectangular waveguide phased arrays. The Bell System Technical Journal, 47(1):93–116, January 1968. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol47/bstj47-1-
93.pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-1-93.pdf.

Gloge:1968:CMB

[GW68b] D. Gloge and D. Weiner. The capacity of multiple beam waveguides and optical delay lines. The Bell System Technical Journal, 47(10):2095–2109, December 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

                                        bell-labs.com/BSTJ/images/
                                        Vol47/bstj47-10-2095.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol47-1968/articles/
                                        bstj47-10-2095.pdf.

Gabbe:1967:SAM

[GWB67] J. D. Gabbe, M. B. Wilk, and W. L. Brown. Statistical analysis and modeling of the high-energy proton data from the Telstar I satellite. The Bell System Technical Journal, 46(7):1301–1450, September 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol46/bstj46-7-1301.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-7-1301.pdf.

Geller:1964:IIM

[GWES64] S. Geller, H. J. Williams, G. P. Espinosa, and R. C. Sherwood. Importance of intrasublattice magnetic interactions and of substitutional ion type in the behavior of substituted yttrium iron garnets. The Bell System Technical Journal, 43(2):565–623, March 1964. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-2-565.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-2-565.pdf.

Haenschke:1963:ADM

[Hae63]  D. G. Haenschke.  Analy-

sis of delay in mathematical switching models for data systems. The Bell System Technical Journal, 42(3):709–736, May 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-3-709.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-3-709.pdf.

Hall:1969:PSH

[Hal69] Harry M. Hall. Power spectrum of hard-limited Gaussian processes. The Bell System Technical Journal, 48(9): 3031–3057, November 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-9-3031.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-9-3031.pdf.

Hamasaki:1964:TUP

[Ham64] J. Hamasaki. A theory of a unilateral parametric amplifier using two diodes. The Bell System Technical Journal, 43(3):1123–1147, May 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-3-1123.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-3-1123.pdf.

Hamilton:1967:WVP

[Ham67] R. L. Hamilton. Water vapor permeability of polyethylene and other plastic materials. The Bell System Technical Journal, 46(2):391–415, February 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol46/bstj46-2-391.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-2-391.pdf.

Hatch:1963:RTS

[HBK63] R. W. Hatch, S. D. Bennett, and J. P. Kinzer. Results of the Telstar system communication tests. The Bell System Technical Journal, 42(4):1561–1629, July 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-4-1561.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-4-1561.pdf.

Houghton:1961:TET

[HD61] E. W. Houghton and E. J. Drazy. Test equipment for the TH radio system. The Bell System Technical Journal, 40(6):1717–1743, November 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-6-1717.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-6-1717.pdf.

Harris:1964:CFD

[HDSD64] S. E. Harris, R. W. DeGrasse, and E. O. Schulz-Dubois. Cutoff frequencies of the dielectrically loaded comb structure as used in traveling-wave masers. The Bell System Technical Journal, 43(1):437–484, January 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-1-437.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-1-437.pdf.

Heidenreich:1966:EPC

[Hei66] R. D. Heidenreich. Electron phase contrast images of molecular detail. The Bell System Technical Journal, 45(4):651–656, April 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-4-651.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-4-651.pdf.

Hein:1967:CCC

[Hei67] V. L. Hein. Convection and conduction cooling of substrates containing multiple heat sources. The Bell System Technical Journal, 46(8):1659–1678, October 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol46/bstj46-8-1659.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-8-1659.pdf.

Heidenreich:1968:BJB

[Hei68] R. D. Heidenreich. B.S.T.J. briefs: AC focusing of an electron microscope objective lens. The Bell System Technical Journal, 47(2):265–268, February 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol47/bstj47-2-265.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-2-265.pdf.

Helms:1965:MRE

[Hel65] H. D. Helms. Maximally reliable exponential prediction equations for data-rate-limited tracking servomechanisms. The Bell System Technical Journal, 44 (10):2337–2362, December 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol44/bstj44-10-2337.pdf;

http://www.alcatel-lucent.
com/bstj/vol44-1965/articles/
bstj44-10-2337.pdf.

Helder:1966:CET

[Hel66]  George K. Helder.  Cus-

tomer evaluation of telephone circuits with delay. The Bell System Technical Journal, 45(7):1157–1191, September 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-7-1157.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-7-1157.pdf.

Haury:1961:TRS

[HF61] P. T. Haury and W. O.

Fullerton. TH radio system equipment aspects. The Bell System Technical Journal, 40(6):1495–1520, November 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-6-1495.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-6-1495.pdf.

Holt:1962:ENR

[HGG62] V. E. Holt, R. J. Grosh, and R. Geynet. Evaluation of the net radiant heat transfer between specularly reflecting plates. The Bell System Technical Journal, 41(6):1865–1874, November 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol41/bstj41-6-1865.
pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-6-1865.pdf.

Hubbard:1967:SSR

[HGW^{+}67] W. M. Hubbard, J. E. Goell, W. D. Warters, R. D. Standley, G. D. Mandeville, T. P. Lee, R. C. Shaw, and P. L. Clouser. A solid-state regenerative repeater for guided millimeterwave communication systems. The Bell System Technical Journal, 46(9):1977–2018, November 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-9-1977.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-9-1977.pdf.

Houghton:1961:FTT

[HH61] E. W. Houghton and R. W. Hatch. FM terminal transmitter and receiver for the TH radio system. The Bell System Technical Journal, 40 (6):1587–1626, November 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-6-1587.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-6-1587.pdf.

Heidenreich:1965:NEE

[HH65] R. D. Heidenreich and R. W.

Hamming. Numerical evaluation of electron image phase contrast. The Bell System Technical Journal, 44(2):207–233, February 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-2-207.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-2-207.pdf.

Hathaway:1968:TMR

[HHJP68] S. D. Hathaway, W. G. Hensel, D. R. Jordan, and R. C. Prime. TD-3 microwave radio relay system. The Bell System Technical Journal, 47 (7):1143–1188, September 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-7-1143.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-7-1143.pdf.

Harr:1964:ONEb

[HHS64] J. A. Harr, Mrs.E. S. Hoover, and R. B. Smith. Organization of no. 1 ESS stored program. The Bell System Technical Journal, 43(5):1923–1959, September 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-5-1923.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-5-1923.pdf.

Highleyman:1962:DAP

[Hig62] W. H. Highleyman. The design and analysis of pattern recognition experiments. The Bell System Technical Journal, 41(2):723–744, March 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol41/bstj41-2-723.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol41-1962/
                                        articles/bstj41-2-723.pdf.

Higgins:1965:SBS

[Hig65] W. H. C. Higgins. A survey of Bell System progress in electronic switching. The Bell System Technical Journal, 44(6):937–997, July/August 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol44/bstj44-6-937.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol44-1965/
                                        articles/bstj44-6-937.pdf.

Hill:1963:CSA

[Hil63] D. W. Hill. Calculation of the spin-axis orientation of the TELSTAR satellites from optical data. The Bell System Technical Journal, 42(6): 2943–2960, November 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-6-2943.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-6-2943.pdf.

Hines:1960:HFN

[Hin60] M. E. Hines. High-frequency negative-resistance circuit principles for esaki diode applications. The Bell System Technical Journal, 39(3):477–513, May 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-3-477.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-3-477.pdf.

Hittinger:1963:CTP

[Hit63]  W. C. Hittinger.  Compo-

nents for the Telstar Project. The Bell System Technical Journal, 42(4):1659–1664, July 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-4-1659.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-4-1659.pdf.

Hrycak:1963:SST

[HKM^{+}63] P. Hrycak, D. E. Koontz, C. Maggs, J. W. Stafford, B. A. Unger, and A. M. Wittenberg. The Spacecraft Structure and thermal design considerations. The Bell System Technical Journal, 42(4):973–1005, July 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-4-973.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-4-973.pdf.

Hines:1963:ECC

[HLT63] J. N. Hines, Tingye Li, and R. H. Turrin. The electrical characteristics of the conical horn-reflector antenna. The Bell System Technical Journal, 42(4):1187–1211, July 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-4-1187.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-4-1187.pdf.

Hallenbeck:1963:NMS

[HM63] F. J. Hallenbeck and J. J. Mahoney, Jr. The new L multiplex — system description and design objectives. The Bell System Technical Journal, 42(2):207–221, March 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-2-207.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-2-207.pdf.

Hubbard:1967:EVE

[HM67] W. M. Hubbard and G. D. Mandeville. Experimental verification of the error-rate performance of two types of regenerative repeaters for differentially coherent phase-shift-keyed signals. The Bell System Technical Journal, 46(6):1173–1202, July/ August 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-6-1173.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-6-1173.pdf.

Hogg:1963:BJB

[Hog63] D. C. Hogg. B.S.T.J. briefs:

On the spectrum of optical waves propagated through the atmosphere. The Bell System Technical Journal, 42(6): 2967–2969, November 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-6-2967.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-6-2967.pdf.

Hogg:1969:SAM

[Hog69] D. C. Hogg. Statistics an attenuation of microwaves by intense rain. The Bell System Technical Journal, 48(9): 2949–2962, November 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-9-2949.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-9-2949.pdf.

Holtzman:1966:SNR

[Hol66] J. M. Holtzman. Signal-noise ratio maximization using the Pontryagin Maximum Principle. The Bell System Technical Journal, 45(3):473–489, March 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol45/bstj45-3-473.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol45-1966/
                                        articles/bstj45-3-473.pdf.

Holtzman:1967:CME

[Hol67]  J. M. Holtzman.  Contrac-

tion maps and equivalent linearization. The Bell System Technical Journal, 46(10):2405– 2435, December 1967. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol46/bstj46-10-2405.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol46-1967/articles/
                                        bstj46-10-2405.pdf.

Holtzman:1968:NDF

[Hol68]  J. M. Holtzman.  Nonlin-

ear distortion in feedback systems. The Bell System Technical Journal, 47(4):503–509, April 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol47/bstj47-4-503.

pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-4-503.pdf.

Hoth:1963:TSS

[HOW63] D. F. Hoth, E. F. O’Neill, and I. Welber. The Telstar satellite system. The Bell System Technical Journal, 42(4):765–799, July 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-4-765.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-4-765.pdf.

Husta:1960:MLR

[HP60] P. Husta and G. E. Perreault.

Magnetic latching relays using glass-sealed contacts. The Bell System Technical Journal, 39(6):1553–1571, November 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-6-1553.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-6-1553.pdf.

Hamori:1968:TMR

[HP68] Andras Hamori and Paul L.

Penney. Transmitter modulator and receiver shift modulator. The Bell System Technical Journal, 47(7):1289–1299, September 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/

                                        images/Vol47/bstj47-7-1289.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-7-1289.pdf.

Haynie:1963:ACT

[HR63] G. D. Haynie and P. E. Rosenfeld. An automated 20-20,000-cps transmission measuring set for laboratory use. The Bell System Technical Journal, 42(6):2501–2531, November 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-6-2501.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-6-2501.pdf.

Hogg:1961:ERW

[HS61] D. C. Hogg and R. A. Semplak. The effect of rain and water vapor on sky noise at centimeter wavelengths. The Bell System Technical Journal, 40(5):1331–1348, September 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol40/bstj40-5-1331.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol40-1961/
                                        articles/bstj40-5-1331.pdf.

Hutchison:1963:RTS

[HS63] P. T. Hutchison and R. A. Swift. Results of the Telstar satellite space experiments. The Bell System Technical Journal, 42(4):1475–1504, July 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-4-1475.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-4-1475.pdf.

Hogg:1964:ESN

[HS64] D. C. Hogg and R. A. Semplak.

An experimental study of nearfield Cassegrainian antennas. The Bell System Technical Journal, 43(6):2677–2704, November 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-6-2677.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-6-2677.pdf.

Heffes:1969:UAL

[HS69] Harry Heffes and Philip E.

Sarachik. Uniform approximation of linear systems. The Bell System Technical Journal, 48(1):209–231, January 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-1-209.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-1-209.pdf.

Hsiao:1967:CSS

[Hsi67] J. K. Hsiao. Characteristics of superconductor strip transmission lines with periodic load.

The Bell System Technical Journal, 46(8):1679–1697, October 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol46/bstj46-8-1679.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-8-1679.pdf.

Hsu:1962:IRZ

[Hsu62]  J. C. Hsu.  Integral repre-

sentation of zero-memory nonlinear functions. The Bell System Technical Journal, 41 (6):1813–1830, November 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol41/bstj41-6-1813.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol41-1962/
                                        articles/bstj41-6-1813.pdf.

Hathaway:1963:TRR

[HSW63] S. D. Hathaway, D. D. Sagaser, and J. A. Word. The TL radio relay system. The Bell System Technical Journal, 42(5):2297–2353, September 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-5-2297.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-5-2297.pdf.

Harr:1964:ONEa

[HTU64] J. A. Harr, F. F. Taylor, and W. Ulrich. Organization of no. 1 ESS central processor. The Bell System Technical Journal, 43(5):1845–1922, September 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-5-1845.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-5-1845.pdf.

Haugk:1964:STN

[HTZ64] G. Haugk, S. H. Tsiang, and L. Zimmerman. System testing of the no. 1 Electronic Switching System. The Bell System Technical Journal, 43(5):2575–2592, September 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-5-2575.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-5-2575.pdf.

Hopcroft:1967:AUT

[HU67] J. E. Hopcroft and J. D. Ullman. An approach to a unified theory of automata. The Bell System Technical Journal, 46(8):1793–1829, October 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-8-1793.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-8-1793.pdf.

Hubbard:1966:EFW

[Hub66]  W. M. Hubbard.  The ef-

fect of a finite-width decision threshold on binary differentially coherent PSK systems. The Bell System Technical Journal, 45(2):307–319, February 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol45/bstj45-2-307.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol45-1966/
                                        articles/bstj45-2-307.pdf.

Hubbard:1967:BJB

[Hub67a] W. M. Hubbard. B.S.T.J. briefs: The effect of noise correlation on binary differentially coherent PSK communication systems. The Bell System Technical Journal, 46(1):277–280, January 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol46/bstj46-1-277.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-1-277.pdf.

Hubbard:1967:EII

[Hub67b] W. M. Hubbard. The effect of intersymbol interference on error rate in binary differentiallycoherent phase-shifted-keyed systems. The Bell System Technical Journal, 46(6):1149–1172, July/ August 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/

images/Vol46/bstj46-6-1149.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-6-1149.pdf.

Holdaway:1964:ETS

[HVW64] V. L. Holdaway, W. Van Haste, and E. J. Walsh. Electron tubes for the SD submarine cable system. The Bell System Technical Journal, 43(4):1311–1338, July 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-4-1311.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-4-1311.pdf.

Hatch:1961:ARC

[HW61] R. W. Hatch and P. R. Wickliffe. Auxiliary radio channels for the TH radio relay system. The Bell System Technical Journal, 40(6):1647–1663, November 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-6-1647.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-6-1647.pdf.

Hogg:1965:PMG

[HW65] D. C. Hogg and R. W. Wilson.

A precise measurement of the gain of a large horn-reflector antenna. The Bell System Technical Journal, 44(6):1019–1030, July/ August 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol44/bstj44-6-1019.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol44-1965/
                                        articles/bstj44-6-1019.pdf.

Haner:1966:CDN

[HW66] R. L. Haner and I. E. Wood. Circuit design of the N3 carrier terminal. The Bell System Technical Journal, 45(6):801–844, July/ August 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol45/bstj45-6-801.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol45-1966/
                                        articles/bstj45-6-801.pdf.

Hyde:1964:ASN

[Hyd64]  J. P. Hyde.  The ALPAK

system for nonnumerical algebra on a digital computer — III: Systems of linear equations and a class of side relations. The Bell System Technical Journal, 43(4):1547–1562, July 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-4-1547.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-4-1547.pdf.

Ingle:1967:ATM

[IKM67] J. F. Ingle, J. J. Kokinda, and G. E. McLaughlin. An automatic transmission measuring system for telephone trunks.

The Bell System Technical Journal, 46(9):1935–1975, November 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-9-1935.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-9-1935.pdf.

Irvin:1962:RBS

[Irv62] John C. Irvin. Resistivity of bulk silicon and of diffused layers in silicon. The Bell System Technical Journal, 41(2):387–410, March 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol41/bstj41-2-387.
pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-2-387.pdf.

Iwersen:1969:CQN

[Iwe69] J. E. Iwersen. Calculated quantizing noise of single-integration delta-modulation coders. The Bell System Technical Journal, 48(7):2359–2389, September 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-7-2359.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-7-2359.pdf.

Jacobs:1967:CAM

[Jac67] Ira Jacobs. Comparison of Mary modulation systems. The Bell System Technical Journal, 46(5):843–864, May/June 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-5-843.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-5-843.pdf.

Jagerman:1965:CSA

[Jag65]  D. Jagerman.  Cosine sum

approximation and synthesis of array antennas. The Bell System Technical Journal, 44(8):1761–1777, October 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-8-1761.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-8-1761.pdf.

Jakes:1961:PBT

[Jak61] William C. Jakes, Jr. Participation of Bell Telephone Laboratories in Project Echo and experimental results. The Bell System Technical Journal, 40(4):975–1028, July 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-4-975.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-4-975.pdf.

Jakes:1963:PHS

[Jak63]  W. C. Jakes, Jr.  Partic-

ipation of the Holmdel Station in the Telstar Project. The Bell System Technical Journal, 42(4):1421–1447, July 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-4-1421.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-4-1421.pdf.

Johnston:1965:BJB

[JDC65] R. L. Johnston, B. C. De Loach, Jr., and B. G. Cohen. B.S.T.J. brief: A silicon diode microwave oscillator. The Bell System Technical Journal, 44(2):369–372, February 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-2-369.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-2-369.pdf.

James:1960:RLC

[JJ60] D. B. James and J. D. Johannesen. A remote line concentrator for a time-separation switching experiment. The Bell System Technical Journal, 39(1):31– 57, January 1960. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol39/bstj39-1-
31.pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol39-1960/
                                        articles/bstj39-1-31.pdf.

Julesz:1968:QMH

[JJ68] Bela Julesz and Stephen C. Johnson. ‘mental holography’: Stereograms portraying ambiguously perceivable surfaces. The Bell System Technical Journal, 47 (10):2075–2093, December 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol47/bstj47-10-2075.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol47-1968/articles/
                                        bstj47-10-2075.pdf.

Julesz:1962:ASP

[JM62] Bela Julesz and Joan E. Miller. Automatic stereoscopic presentation of functions of two variables. The Bell System Technical Journal, 41(2):663–676, March 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol41/bstj41-2-663.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol41-1962/
                                        articles/bstj41-2-663.pdf.

Jewett:1968:PS

[JM68] W. E. Jewett and S. Mottle. Power system. The Bell System Technical Journal, 47(7):1487–1509, September 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/

images/Vol47/bstj47-7-1487.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-7-1487.pdf.

Joel:1968:PSN

[Joe68] A. E. Joel, Jr. On permutation switching networks. The Bell System Technical Journal, 47(5):813–822, May/June 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-5-813.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-5-813.pdf.

Johansson:1964:MRR

[Joh64] S. G. Johansson. Manufacture of rigid repeaters and ocean-block equalizers. The Bell System Technical Journal, 43(4):1275–1310, July 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-4-1275.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-4-1275.pdf.

Jensen:1968:MTR

[JRS68] R. M. Jensen, R. E. Rowe, and R. E. Sherman. Microwave transmitter and receiver. The Bell System Technical Journal, 47(7):1189–1225, September 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol47/bstj47-7-1189.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-7-1189.pdf.

Julesz:1960:BDP

[Jul60] Bela Julesz. Binocular depth perception of computer-generated patterns. The Bell System Technical Journal, 39(5): 1125–1162, September 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol39/bstj39-5-1125.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol39-1960/
                                        articles/bstj39-5-1125.pdf.

Julesz:1967:SMP

[Jul67]  Bela Julesz.  The suppres-

sion of monocularly perceivable symmetry during binocular fusion. The Bell System Technical Journal, 46(6):1203–1221, July/ August 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol46/bstj46-6-1203.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-6-1203.pdf.

Kadota:1964:ORB

[Kad64] T. T. Kadota. Optimum reception of binary Gaussian signals. The Bell System Technical Journal, 43(6):2767–2810, November 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-6-2767.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-6-2767.pdf.

Kadota:1965:ORB

[Kad65a] T. T. Kadota. Optimum reception of binary sure and Gaussian signals. The Bell System Technical Journal, 44(8):1621–1658, October 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-8-1621.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-8-1621.pdf.

Kadota:1965:ORA

[Kad65b] T. T. Kadota. Optimum reception of M-ary Gaussian signals in Gaussian noise. The Bell System Technical Journal, 44(9):2187–2197, November 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-9-2187.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-9-2187.pdf.

Kadota:1966:SOE

[Kad66]  T. T. Kadota.  Simultane-

ously orthogonal expansion of two stationary Gaussian processes — examples. The Bell System Technical Journal, 45(7):1071–1096, September 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol45/bstj45-7-1071.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol45-1966/
                                        articles/bstj45-7-1071.pdf.

Kadota:1967:GOR

[Kad67a] T. T. Kadota. Generalized optimum receivers of Gaussian signals. The Bell System Technical Journal, 46(3):577–591, March 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol46/bstj46-3-577.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-3-577.pdf.

Kadota:1967:IES

[Kad67b] T. T. Kadota. Integral equation for simultaneous diagonalization of two covariance kernels. The Bell System Technical Journal, 46(5):883–892, May/ June 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol46/bstj46-5-883.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-5-883.pdf.

Kahng:1964:ATG

[Kah64] D. Kahng. Au-n-type GaAs Schottky barrier and its varactor application. The Bell System Technical Journal, 43 (1):215–224, January 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-1-215.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-1-215.pdf.

Kahng:1967:BJBb

[Kah67]  D. Kahng.  B.S.T.J. briefs:

Semipermanent memory using capacitor charge storage and IGFET read-out. The Bell System Technical Journal, 46(6):1296–1300, July/ August 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-6-1296.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-6-1296.pdf.

Kaiser:1968:SGS

[Kai68] P. Kaiser. The stream guide, a simple, low-loss optical guiding medium. The Bell System Technical Journal, 47 (5):761–765, May/June 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-5-761.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-5-761.pdf.

Karnaugh:1966:MOS

[Kar66] M. Karnaugh. A model for the organic synchronization of communications systems. The Bell System Technical Journal, 45 (10):1705–1735, December 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol45/bstj45-10-1705.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol45-1966/articles/
                                        bstj45-10-1705.pdf.

Katzenelson:1965:ASN

[Kat65] Jacob Katzenelson. An algorithm for solving nonlinear resistor networks. The Bell System Technical Journal, 44(8):1605–1620, October 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol44/bstj44-8-1605.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol44-1965/
                                        articles/bstj44-8-1605.pdf.

Kahng:1964:GES

[KD64] D. Kahng and L. A. D’Asaro. Gold-epitaxial silicon high-frequency diodes. The Bell System Technical Journal, 43(1):225–232, January 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-1-225.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-1-225.pdf.

Keller:1964:RDB

[Kel64] A. C. Keller. Recent developments in Bell System relays — particularly sealed contact and miniature relays. The Bell System Technical Journal, 43(1):15– 44, January 1964. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol43/bstj43-1-
15.pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-1-15.pdf.

Kester:1968:NRF

[Kes68] W. A. Kester. A new reference frequency standard for the L multiplex system. The Bell System Technical Journal, 47 (10):2183–2205, December 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol47/bstj47-10-2183.pdf;
http://www.alcatel-lucent.
com/bstj/vol47-1968/articles/
bstj47-10-2183.pdf.

Kinariwala:1966:SCN

[KG66] B. K. Kinariwala and A. Gersho. A stability criterion for nonuniformly sampled and distributed parameter systems. The Bell System Technical Journal, 45(7):1153–1155, September 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-7-1153.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-7-1153.pdf.

Kruskal:1966:GID

[KH66] J. B. Kruskal and R. E. Hart. A geometric interpretation of diagnostic data from a digital machine: Based on a study of the Morris, Illinois Electronic Central Office. The Bell System Technical Journal, 45(8):1299–1338, October 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol45/bstj45-8-1299.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol45-1966/
                                        articles/bstj45-8-1299.pdf.

Kibler:1961:PES

[Kib61] L. U. Kibler. Project Echo: Standby receiver system. The Bell System Technical Journal, 40(4):1129–1147, July 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol40/bstj40-4-1129.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol40-1961/
                                        articles/bstj40-4-1129.pdf.

Kibler:1965:WTC

[Kib65a] L. U. Kibler. A 14-watt transistor CW amplifier with a 50-mc bandwidth. The Bell System Technical Journal, 44(9):2003–2015, November 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol44/bstj44-9-2003.

pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-9-2003.pdf.

Kibler:1965:MWT

[Kib65b] L. U. Kibler. An 80-megabit 15watt transistor pulse amplifier. The Bell System Technical Journal, 44(9):1977–2001, November 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-9-1977.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-9-1977.pdf.

Kibler:1969:CRR

[Kib69] L. U. Kibler. The cutoff region of a rectangular waveguide with losses, its properties and uses. The Bell System Technical Journal, 48(7):2221–2258, September 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-7-2221.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-7-2221.pdf.

Kinariwala:1962:TECa

[Kin62a] B. K. Kinariwala. Timing errors in a chain of regenerative repeaters, I. The Bell System Technical Journal, 41(6):1769–1780, November 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/

                                        images/Vol41/bstj41-6-1769.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol41-1962/
                                        articles/bstj41-6-1769.pdf.

Kinariwala:1962:TECb

[Kin62b] B. K. Kinariwala. Timing errors in a chain of regenerative repeaters, II. The Bell System Technical Journal, 41(6):1781–1797, November 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol41/bstj41-6-1781.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol41-1962/
                                        articles/bstj41-6-1781.pdf.

Kinariwala:1964:TEC

[Kin64]  B. K. Kinariwala.  Timing

errors in a chain of regenerative repeaters, III. The Bell System Technical Journal, 43(4):1481–1504, July 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-4-1481.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-4-1481.pdf.

Kinariwala:1966:TCS

[Kin66]  B. K. Kinariwala.  Theory

of cascaded structures: Lossless transmission lines. The Bell System Technical Journal, 45(4):631–649, April 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-4-631.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-4-631.pdf.

Kleinman:1962:DAU

[KK62] D. A. Kleinman and P. P. Kisliuk.

Discrimination against unwanted orders in the Fabry–Perot resonator. The Bell System Technical Journal, 41(2):453–462, March 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol41/bstj41-2-453.
pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-2-453.pdf.

Keister:1964:NES

[KKV64] W. Keister, R. W. Ketchledge, and H. E. Vaughan. No: 1 ESS system organization and objectives. The Bell System Technical Journal, 43(5):1831–1844, September 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-5-1831.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-5-1831.pdf.

Kinzer:1961:EAT

[KL61] J. P. Kinzer and J. F. Laidig.

Engineering aspects of the TH microwave radio relay system. The Bell System Technical Journal, 40(6):1459–1494, November 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol40/bstj40-6-1459.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol40-1961/
                                        articles/bstj40-6-1459.pdf.

Kahng:1965:BBP

[KL65] D. Kahng and M. P. Lepselter. BSTJ brief: Planar epitaxial silicon Schottky barrier diodes. The Bell System Technical Journal, 44(7):1525–1528, September 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol44/bstj44-7-1525.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol44-1965/
                                        articles/bstj44-7-1525.pdf.

Klauder:1960:DRS

[Kla60]  J. R. Klauder.  The de-

sign of radar signals having both high range resolution and high velocity resolution. The Bell System Technical Journal, 39(4):809–820, July 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol39/bstj39-4-809.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol39-1960/
                                        articles/bstj39-4-809.pdf.

Klauder:1961:CMH

[Kla61] John R. Klauder. The calculations of metallic hall constants: Topological aspects and applications to copper. The Bell System Technical Journal, 40(5):1349–1367, September 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-5-1349.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-5-1349.pdf.

Kleinman:1961:CSC

[Kle61] D. A. Kleinman. Considerations on the solar cell. The Bell System Technical Journal, 40(1):85–115, January 1961. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol40/bstj40-1-
85.pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-1-85.pdf.

Kleinman:1964:MRE

[Kle64] D. A. Kleinman. The maser rate equations and spiking. The Bell System Technical Journal, 43(4):1505–1532, July 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-4-1505.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-4-1505.pdf.

Klemmer:1967:SET

[Kle67] E. T. Klemmer. Subjective evaluation of transmission delay in telephone conversations. The Bell System Technical Journal, 46(6):1141–1147, July/ August 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-6-1141.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-6-1141.pdf.

Kelly:1961:BDC

[KLV61] John L. Kelly, Jr., Carol Lochbaum, and V. A. Vyssotsky. A block diagram compiler. The Bell System Technical Journal, 40(3):669–676, May 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-3-669.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-3-669.pdf.

King:1961:OKA

[KM61] A. P. King and G. D. Mandeville. The observed 33 to 90 Kmc attenuation of twoinch improved waveguide. The Bell System Technical Journal, 40(5):1323–1330, September 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-5-1323.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-5-1323.pdf.

Kozemchak:1969:CAC

[KML69] E. B. Kozemchak and M. A.

Murray-Lasso. Computer-aided circuit design by singular imbedding. The Bell System Technical Journal, 48(1):275–315, January 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-1-275.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-1-275.pdf.

Klahne:1961:PEA

[KNG61] R. Klahne, J. A. Norton, and J. A. Githens. Project Echo: Antenna steering system. The Bell System Technical Journal, 40(4):1207–1225, July 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-4-1207.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-4-1207.pdf.

Koehler:1967:CCC

[Koe67] Dankwart Koehler. The charge controlled concept in the form of equivalent circuits, representing a link between classic large signal diode and transistor models. The Bell System Technical Journal, 46(3):523–576, March 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-3-523.

                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-3-523.pdf.

Kogelnik:1964:BJB

[Kog64] H. Kogelnik. B.S.T.J. briefs: Matching of optical modes. The Bell System Technical Journal, 43(1):334–337, January 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-1-334.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-1-334.pdf.

Kogelnik:1965:BJB

[Kog65a] H. Kogelnik. B.S.T.J. brief: Holographic image projection through inhomogeneous media. The Bell System Technical Journal, 44(10):2451–2455, December 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

                                        bell-labs.com/BSTJ/images/
                                        Vol44/bstj44-10-2451.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol44-1965/articles/
                                        bstj44-10-2451.pdf.

Kogelnik:1965:IOM

[Kog65b] Herwig Kogelnik. Imaging of optical modes – resonators with internal lenses. The Bell System Technical Journal, 44(3):455–494, March 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/

images/Vol44/bstj44-3-455.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-3-455.pdf.

Kogelnik:1969:CWT

[Kog69]  Herwig Kogelnik.  Coupled

wave theory for thick hologram gratings. The Bell System Technical Journal, 48(9): 2909–2947, November 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-9-2909.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-9-2909.pdf.

Klauder:1960:TDC

[KPDA60] J. R. Klauder, A. C. Price, S. Darlington, and W. J. Albersheim. The theory and design of chirp radars. The Bell System Technical Journal, 39(4):745–808, July 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-4-745.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-4-745.pdf.

See erratum [Ano60g].

Kennedy:1962:USR

[KR62] J. T. Kennedy and J. W.

Rosson. The use of solar radio emission for the measurement of radar angle errors. The Bell System Technical Journal, 41(6):1799–1812, November 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol41/bstj41-6-1799.
pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-6-1799.pdf.

Kronacher:1967:SLL

[Kro67] G. Kronacher. Scaling laws for large shields in quasi-stationary magnetic fields. The Bell System Technical Journal, 46(10): 2323–2339, December 1967. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol46/bstj46-10-2323.pdf;
http://www.alcatel-lucent.
com/bstj/vol46-1967/articles/
bstj46-10-2323.pdf.

Kruskal:1969:WSA

[Kru69]  Joseph B. Kruskal.  Work-

scheduling algorithms: A nonprobabilistic queuing study (with possible application to no. 1 ESS). The Bell System Technical Journal, 48(9): 2963–2974, November 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-9-2963.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-9-2963.pdf.

Kerwien:1963:DMP

[KS63] A. E. Kerwien and L. H.

Steiff. Design of a 150-megacycle pocket receiver for the BELL-BOY personal signalling system. The Bell System Technical Journal, 42(3):527–565, May 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-3-527.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-3-527.pdf.

Kimber:1966:DCP

[KS66] R. A. Kimber and R. R.

Stokes. Design considerations and proposals for compatible U.S. subsidiary coinage. The Bell System Technical Journal, 45(4):493–522, April 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-4-493.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-4-493.pdf.

Kahng:1967:BJBa

[KS67] D. Kahng and S. M. Sze. B.S.T.J.

briefs: A floating gate and its application to memory devices. The Bell System Technical Journal, 46(6):1288–1295, July/ August 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-6-1288.

                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-6-1288.pdf.

Kelcourse:1969:EMS

[KSW69] F. C. Kelcourse, W. G. Sheerer, and R. J. Wirtz. Equalizing and main station repeaters. The Bell System Technical Journal, 48(4):889–952, April 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-4-889.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-4-889.pdf.

Kurokawa:1961:MNF

[KU61] K. Kurokawa and M. Uenohara. Minimum noise figure of the variable-capacitance amplifier. The Bell System Technical Journal, 40(3):695–722, May 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol40/bstj40-3-695.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol40-1961/
                                        articles/bstj40-3-695.pdf.

Kreipe:1962:ILW

[KU62] H. L. Kreipe and H. G. Unger. Imperfections in lined waveguide. The Bell System Technical Journal, 41(5):1589–1619, September 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/

images/Vol41/bstj41-5-1589.
pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-5-1589.pdf.

Kurokawa:1962:UPC

[Kur62] Kaneyuki Kurokawa. On the use of passive circuit measurements for the adjustment of variable capacitance amplifiers. The Bell System Technical Journal, 41(1):361–381, January 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol41/bstj41-1-361.
pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-1-361.pdf.

Kurokawa:1965:DTB

[Kur65] K. Kurokawa. Design theory of balanced transistor amplifiers. The Bell System Technical Journal, 44(8):1675–1698, October 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-8-1675.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-8-1675.pdf.

Kurtz:1966:DEO

[Kur66] S. K. Kurtz. Design of an electro-optic polarization switch for a high-capacity high-speed digital light deflecting system. The Bell System Technical Journal, 45(8):1209–1246, October 1966. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol45/bstj45-8-1209.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol45-1966/
                                        articles/bstj45-8-1209.pdf.

Kurokawa:1967:DHF

[Kur67] K. Kurokawa. The dynamics of high-field propagating domains in bulk semiconductors. The Bell System Technical Journal, 46 (10):2235–2259, December 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol46/bstj46-10-2235.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol46-1967/articles/
                                        bstj46-10-2235.pdf.

Kurokawa:1969:SBC

[Kur69] K. Kurokawa. Some basic characteristics of broadband negative resistance oscillator circuits. The Bell System Technical Journal, 48(6):1937–1955, July/ August 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-6-1937.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-6-1937.pdf.

Kyong:1969:JCN

[Kyo69] Sang H. Kyong. Jump criteria of nonlinear control systems and the validity of statistical linearization approximation.

The Bell System Technical Journal, 48(7):2529–2543, September 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-7-2529.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-7-2529.pdf.

Landau:1960:RBL

[Lan60] H. J. Landau. On the recovery of a band-limited signal, after instantaneous companding and subsequent band limiting. The Bell System Technical Journal, 39(2):351–364, March 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-2-351.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-2-351.pdf.

Lapsley:2013:PMB

[Lap13a] Phil Lapsley. Phreaking out Ma Bell: How a buccaneering young engineer built the little blue box that broke into the biggest network in the world. IEEE Spectrum, ??(??):??, February

2013. URL http://spectrum.
ieee.org/telecom/standards/
phreaking-out-ma-bell.

Lapsley:2013:EPU

[Lap13b] Philip D. Lapsley. Exploding the phone: the untold story of the teenagers and outlaws who hacked Ma Bell. Grove Press, New York, NY, USA, 2013. ISBN 0-8021-2061-X. xvi + 431 + 16 pp. LCCN ????

Larsen:1969:HSS

[Lar69]  Arthur B. Larsen.  A het-

erodyne scanning system for hologram transmission. The Bell System Technical Journal, 48(7):2507–2527, September 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-7-2507.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-7-2507.pdf.

Lee:1960:AFA

[Lee60]  C. Y. Lee.  Automata and

finite automata. The Bell System Technical Journal, 39 (5):1267–1295, September 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol39/bstj39-5-1267.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol39-1960/
                                        articles/bstj39-5-1267.pdf.

Leed:1966:ILP

[Lee66] D. Leed. An insertion loss, phase and delay measuring set for characterizing transistors and twoport networks between 0.25 and 4.2 gc. The Bell System Technical Journal, 45(3):397–440, March 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-3-397.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-3-397.pdf.

Lee:1967:EDA

[Lee67a] W. C. Y. Lee. An energydensity antenna for independent measurement of the electric and magnetic field. The Bell System Technical Journal, 46(7):1587–1599, September 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-7-1587.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-7-1587.pdf.

Lee:1967:SAL

[Lee67b] W. C. Y. Lee. Statistical analysis of the level crossings and duration of fades of the signal from an energy density mobile radio antenna. The Bell System Technical Journal, 46(2):417–448, February 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-2-417.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-2-417.pdf.

Lee:1969:EBC

[Lee69a] S. Y. Lee. Extension of Bode’s constant resistance lattice synthesis of transfer impedance function. The Bell System Technical Journal, 48(3):637–657, March 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-3-637.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-3-637.pdf.

Lee:1969:ECF

[Lee69b] W. C. Y. Lee. An extended correlation function of two random variables applied to mobile radio transmission. The Bell System Technical Journal, 48(10): 3423–3440, December 1969. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol48/bstj48-10-3423.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol48-1969/articles/
                                        bstj48-10-3423.pdf. See erra-

tum [Ano70].

Legg:1960:AQF

[Leg60] V. E. Legg. Analysis of quality factor of annular core inductors. The Bell System Technical Journal, 39(1):105–126, January 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol39/bstj39-1-105.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol39-1960/
                                        articles/bstj39-1-105.pdf.

Lepselter:1966:BLT

[Lep66]  M. P. Lepselter.  Beam-

lead technology. The Bell System Technical Journal, 45 (2):233–253, February 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-2-233.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-2-233.pdf.

Lepselter:1968:BJB

[Lep68]  M. P. Lepselter.  B.S.T.J.

briefs: Air-insulated beam-lead crossovers for integrated circuits. The Bell System Technical Journal, 47(2):269–271, February 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-2-269.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-2-269.pdf.

Levine:1961:NTI

[Lev61] N. Levine. A new technique for increasing the flexibility of recursive least squares data smoothing. The Bell System Technical Journal, 40(3):821–840, May 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-3-821.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-3-821.pdf.

Lewis:1961:EDE

[Lew61]  J. A. Lewis.  The effect

of driving electrode shape on the electrical properties of piezoelectric crystals. The Bell System Technical Journal, 40(5):1259–1280, September 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol40/bstj40-5-1259.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol40-1961/
                                        articles/bstj40-5-1259.pdf.

Lewinski:1964:NOM

[Lew64]  D. A. Lewinski.  A new

objective for message circuit noise. The Bell System Technical Journal, 43(2):719–740, March 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-2-719.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-2-719.pdf.

Li:1963:MSA

[Li63]  Tingye Li.  Mode selection

in an aperture-limited concentric maser interferometer. The Bell System Technical Journal, 42(6):2609–2620, November 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-6-2609.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol42-1963/
articles/bstj42-6-2609.pdf.

Li:1965:DLS

[Li65] Tingye Li. Diffraction loss and selection of modes in maser resonators with circular mirrors. The Bell System Technical Journal, 44(5):917–932, May/ June 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-5-917.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-5-917.pdf.

Liehr:1960:QIR

[Lie60]  Andrew D. Liehr.  ‘ionic

radii,’ spin-orbit coupling and the geometrical stability of inorganic complexes. The Bell System Technical Journal, 39(6):1617–1626, November 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-6-1617.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-6-1617.pdf.

Limb:1969:DDW

[Lim69]  J. O. Limb.  Design of

dither waveforms for quantized visual signals. The Bell System Technical Journal, 48(7):2555–2582, September 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-7-2555.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-7-2555.pdf.

Lindner:1962:SSV

[Lin62]  R. Lindner.  Semiconduc-

tor surface varactor. The Bell System Technical Journal, 41(3):803–831, May 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol41/bstj41-3-803.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol41-1962/
                                        articles/bstj41-3-803.pdf.

Linhart:1964:OTT

[Lin64] Peter Linhart. Overflow traffic from a trunk group with balking. The Bell System Technical Journal, 43(1):45–89, January 1964. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol43/bstj43-1-
                                        45.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-1-45.pdf.

Lin:1965:CST

[Lin65] Shen Lin. Computer solutions of the traveling salesman problem. The Bell System Technical Journal, 44(10):2245–2269, December 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

                                        bell-labs.com/BSTJ/images/

Vol44/bstj44-10-2245.pdf;
http://www.alcatel-lucent.
com/bstj/vol44-1965/articles/
bstj44-10-2245.pdf.

Liou:1966:NFS

[Lio66]  M. L. Liou.  Noise in an

FM system due to an imperfect linear transducer. The Bell System Technical Journal, 45(9):1537–1561, November 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-9-1537.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-9-1537.pdf.

Liu:1968:SSN

[Liu68a] S. C. Liu. The spectrum of a simple nonlinear system. The Bell System Technical Journal, 47 (10):2259–2272, December 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol47/bstj47-10-2259.pdf;
http://www.alcatel-lucent.
com/bstj/vol47-1968/articles/
bstj47-10-2259.pdf.

Liu:1968:SAS

[Liu68b] S. C. Liu. Statistical analysis and stochastic simulation of ground-motion data. The Bell System Technical Journal, 47 (10):2273–2298, December 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.

                                        bell-labs.com/BSTJ/images/
                                        Vol47/bstj47-10-2273.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol47-1968/articles/
                                        bstj47-10-2273.pdf.

Liu:1969:SFP

[Liu69] S. C. Liu. Solution of Fokker– Planck equation with applications in nonlinear random vibration. The Bell System Technical Journal, 48(6):2031–2051, July/ August 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-6-2031.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-6-2031.pdf.

Leed:1961:LPS

[LK61] D. Leed and O. Kummer. A loss and phase set for measuring transistor parameters and twoport networks between 5 and 250 mc. The Bell System Technical Journal, 40(3):841–884, May 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol40/bstj40-3-841.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol40-1961/
                                        articles/bstj40-3-841.pdf.

Lloyd:1967:EGE

[Llo67] P. Lloyd. Equations governing the electrical behavior of an arbitrary piezoelectric resonator having N electrodes. The Bell System Technical Journal, 46(8):1881–1900, October 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-8-1881.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-8-1881.pdf.

Limb:1969:DDQ

[LM69a] J. O. Limb and F. W. Mounts.

Digital differential quantizer for television. The Bell System Technical Journal, 48(7): 2583–2599, September 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-7-2583.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-7-2583.pdf.

Lundgren:1969:RRA

[LM69b] C. W. Lundgren and A. S. May.

Radio-relay antenna pointing for controlled interference with geostationary satellites. The Bell System Technical Journal, 48 (10):3387–3422, December 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol48/bstj48-10-3387.pdf;
http://www.alcatel-lucent.
com/bstj/vol48-1969/articles/
bstj48-10-3387.pdf.

Lonnquist:1969:AE

[LMS^{+}69] Chester W. Lonnquist, Joseph C.

Manganello, Robert S. Skinner, Michael T. Skubiak, and Donald J. Wadsworth. Apparatus and equipment. The Bell System Technical Journal, 48(8):2817–2863, October 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-8-2817.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-8-2817.pdf.

Lozier:1963:SSA

[LNI63] J. C. Lozier, J. A. Norton, and M. Iwama. The servo system for antenna positioning. The Bell System Technical Journal, 42(4):1253–1281, July 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-4-1253.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-4-1253.pdf.

Logan:1961:ABS

[Log61] M. A. Logan. An AC bridge for semiconductor resistivity measurements using a four-point probe. The Bell System Technical Journal, 40(3):885–919, May 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-3-885.
pdf;   http://www.alcatel-

lucent.com/bstj/vol40-1961/
articles/bstj40-3-885.pdf.

Logan:1967:SRM

[Log67] M. A. Logan. Sheet resistivity measurements on rectangular surfaces — general solution for four point probe conversion factors. The Bell System Technical Journal, 46(10): 2277–2322, December 1967. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol46/bstj46-10-2277.pdf;
http://www.alcatel-lucent.
com/bstj/vol46-1967/articles/
bstj46-10-2277.pdf.

Landau:1961:PSW

[LP61] H. J. Landau and H. O. Pollak. Prolate spheroidal wave functions, Fourier analysis and uncertainty. II. The Bell System Technical Journal, 40(1):65– 84, January 1961. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol40/bstj40-1-
65.pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-1-65.pdf.

Landau:1962:PSW

[LP62] H. J. Landau and H. O. Pollak. Prolate spheroidal wave functions, Fourier analysis and uncertainty. III. The dimension of the space of essentially time- and band-limited signals. The Bell System Technical Journal, 41(4):1295–1336, July 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol41/bstj41-4-1295.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol41-1962/
                                        articles/bstj41-4-1295.pdf.

Lerch:1964:ACM

[LP64] B. W. Lerch and J. W. Phelps. Armorless cable manufacture. The Bell System Technical Journal, 43(4):1209–1242, July 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-4-1209.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-4-1209.pdf.

Lin:1966:BJB

[LPSL66] L. H. Lin, K. S. Pennington, G. W. Stroke, and A. E. Labeyrie. B.S.T.J. briefs: Multicolor holographic image reconstruction with white-light illumination. The Bell System Technical Journal, 45(4):659–660, April 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol45/bstj45-4-659.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol45-1966/
                                        articles/bstj45-4-659.pdf.

Lucky:1967:AEG

[LR67] R. W. Lucky and H. R.

Rudin. An automatic equalizer for general-purpose communication channels. The Bell System Technical Journal, 46(9):2179–2208, November 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-9-2179.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-9-2179.pdf.

Landau:1966:ORS

[LS66] H. J. Landau and David Slepian.

On the optimality of the regular simplex code. The Bell System Technical Journal, 45(8):1247–1272, October 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-8-1247.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-8-1247.pdf.

Lepselter:1968:SSB

[LS68] M. P. Lepselter and S. M. Sze.

Silicon Schottky barrier diode with near-ideal I–V characteristics. The Bell System Technical Journal, 47(2):195–208, February 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-2-195.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/

                                        articles/bstj47-2-195.pdf.

See erratum [Ano68k].

Lee:1969:FMM

[LS69] T. P. Lee and R. D. Standley. Frequency modulation of a millimeter-wave IM-PATT diode oscillator and related harmonic generation effects. The Bell System Technical Journal, 48(1):143–161, January 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-1-143.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-1-143.pdf.

Luderer:1964:CEW

[LU64] G. W. Luderer and H. G. Unger. Circular electric wave propagation in periodic structures. The Bell System Technical Journal, 43(2):755–783, March 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-2-755.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-2-755.pdf.

Lubowe:1968:PLV

[Lub68]  Anthony G. Lubowe.  Path

length variation in a synchronous satellite communications link. The Bell System Technical Journal, 47(10):2139–2144, December 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

bell-labs.com/BSTJ/images/
Vol47/bstj47-10-2139.pdf;
http://www.alcatel-lucent.
com/bstj/vol47-1968/articles/
bstj47-10-2139.pdf.

Lucky:1963:FAR

[Luc63] R. W. Lucky. A functional analysis relating delay variation and intersymbol interference in data transmission. The Bell System Technical Journal, 42(5):2427–2483, September 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-5-2427.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-5-2427.pdf.

Lucky:1965:AED

[Luc65] R. W. Lucky. Automatic equalization for digital communication. The Bell System Technical Journal, 44(4):547–588, April 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-4-547.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-4-547.pdf.

Lucky:1966:TAE

[Luc66] R. W. Lucky. Techniques for adaptive equalization of digital communication systems. The Bell System Technical Journal, 45(2):255–286, February 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol45/bstj45-2-255.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol45-1966/
                                        articles/bstj45-2-255.pdf.

Lucky:1968:ARR

[Luc68] R. W. Lucky. Adaptive redundancy removal in data transmission. The Bell System Technical Journal, 47(4):549–573, April 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol47/bstj47-4-549.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-4-549.pdf.

Lundquist:1969:DPS

[Lun69] Leif Lundquist. Digital PM spectra by transform techniques. The Bell System Technical Journal, 48(2):397–411, February 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-2-397.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-2-397.pdf.

Lawrence:1960:DJD

[LW60] H. Lawrence and R. M. Warner, Jr. Diffused junction depletion layer calculations. The Bell System Technical Journal, 39(2):389–403, March 1960.

CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-2-389.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-2-389.pdf.

Lundry:1965:NCT

[LW65] W. R. Lundry and L. F. Willey.

The N2 carrier terminal – circuit design. The Bell System Technical Journal, 44(5):761–785, May/ June 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-5-761.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-5-761.pdf.

Lytle:1968:CCT

[Lyt68]  D. W. Lytle.  Convergence

criteria for transversal equalizers. The Bell System Technical Journal, 47(8):1775–1800, October 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-8-1775.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-8-1775.pdf.

Lewis:1964:TGG

[LZ64] J. A. Lewis and E. E. Zajac. A two-gyro, gravity-gradient satellite attitude control system. The Bell System Technical Journal, 43(6):2705–2765, November 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-6-2705.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-6-2705.pdf.

MacColl:1960:CMV

[Mac60]  L. A. MacColl.  Certain

mean values in the theory of the traveling-wave amplifier. The Bell System Technical Journal, 39(2):365–367, March 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol39/bstj39-2-365.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol39-1960/
                                        articles/bstj39-2-365.pdf.

MacWilliams:1961:ECC

[Mac61]  Jessie MacWilliams.  Error-

correcting codes for multiplelevel transmission. The Bell System Technical Journal, 40 (1):281–308, January 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol40/bstj40-1-281.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol40-1961/
                                        articles/bstj40-1-281.pdf.

MacWilliams:1963:TDW

[Mac63] Jessie MacWilliams. A theorem on the distribution of weights in a systematic code. The Bell System Technical Journal, 42(1):79– 94, January 1963. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol42/bstj42-1-
79.pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-1-79.pdf.

MacWilliams:1964:PDS

[Mac64] Jessie MacWilliams. Permutation decoding of systematic codes. The Bell System Technical Journal, 43(1):485–505, January 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-1-485.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-1-485.pdf.

MacWilliams:1965:SPB

[Mac65] Jessie MacWilliams. The structure and properties of binary cyclic alphabets. The Bell System Technical Journal, 44(2):303–332, February 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-2-303.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-2-303.pdf.

Magnani:1968:STP

[Mag68] R. Magnani. Some theorems and procedures for loop-free rooting in directed communications networks. The Bell System Technical Journal, 47(4):465–486, April 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-4-465.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-4-465.pdf.

Manley:1963:TCE

[Man63] J. M. Manley. A three-conductor elementary clogston coaxial transmission line — calculation, fabrication and experiment. The Bell System Technical Journal, 42(6):2551–2574, November 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-6-2551.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-6-2551.pdf.

Manley:1969:GAT

[Man69] J. M. Manley. The generation and accumulation of timing noise in PCM systems — an experimental and theoretical study. The Bell System Technical Journal, 48(3):541–613, March 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-3-541.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-3-541.pdf.

Marcatili:1961:CEH

[Mar61a] E. A. Marcatili. A circularelectric hybrid junction and some channel-dropping filters. The Bell System Technical Journal, 40(1):185–196, January 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-1-185.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-1-185.pdf.

Marcatili:1961:CFS

[Mar61b] E. A. Marcatili. Compression, filtering, and signal-to-noise ratio in a pulse-modulated system. The Bell System Technical Journal, 40(5):1421–1453, September 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-5-1421.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-5-1421.pdf.

Marcatili:1961:EDR

[Mar61c] E. A. Marcatili. Errors in detection of RF pulses embedded in time crosstalk, frequency crosstalk, and noise. The Bell System Technical Journal, 40(3):921–950, May 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-3-921.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol40-1961/
                                        articles/bstj40-3-921.pdf.

Marcatili:1961:MCF

[Mar61d] E. A. Marcatili. Mode-conversion filters. The Bell System Technical Journal, 40(1):149–184, January 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol40/bstj40-1-149.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol40-1961/
                                        articles/bstj40-1-149.pdf.

Marcatili:1961:TFC

[Mar61e] E. A. Marcatili. Time and frequency crosstalk in pulsemodulated systems. The Bell System Technical Journal, 40(3):951–970, May 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol40/bstj40-3-951.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol40-1961/
                                        articles/bstj40-3-951.pdf.

Marcuse:1962:SEB

[Mar62] D. Marcuse. Stimulated emission of bremsstrahlung. The Bell System Technical Journal, 41(5):1557–1571, September 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol41/bstj41-5-1557.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol41-1962/
articles/bstj41-5-1557.pdf.

Marcuse:1963:FDS

[Mar63] Dietrich Marcuse. A further discussion of stimulated emission of bremsstrahlung. The Bell System Technical Journal, 42(2):415–430, March 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-2-415.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-2-415.pdf.

Marcatili:1964:MST

[Mar64a] E. A. J. Marcatili. Modes in a sequence of thick astigmatic lens-like focusers. The Bell System Technical Journal, 43(6):2887–2904, November 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-6-2887.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-6-2887.pdf.

Marcuse:1964:PLR

[Mar64b] D. Marcuse. Propagation of light rays through a lenswaveguide with curved axis. The Bell System Technical Journal, 43(2):741–753, March 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/

                                        images/Vol43/bstj43-2-741.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-2-741.pdf.

Marcuse:1965:PPG

[Mar65a] D. Marcuse. Properties of periodic gas lenses. The Bell System Technical Journal, 44(9):2083–2116, November 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol44/bstj44-9-2083.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol44-1965/
                                        articles/bstj44-9-2083.pdf.

Marcuse:1965:STL

[Mar65b] D. Marcuse. Statistical treatment of light-ray propagation in beam-waveguides. The Bell System Technical Journal, 44(9):2065–2081, November 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol44/bstj44-9-2065.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol44-1965/
                                        articles/bstj44-9-2065.pdf.

Marcatili:1966:LTM

[Mar66a] E. A. J. Marcatili. Light transmission in a multiple dielectric (gaseous and solid) guide. The Bell System Technical Journal, 45(1):97–103, January 1966. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol45/bstj45-1-
97.pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-1-97.pdf.

Marcatili:1966:RPB

[Mar66b] E. A. J. Marcatili. Ray propagation in beam-waveguides with redirectors. The Bell System Technical Journal, 45 (1):105–115, January 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-1-105.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-1-105.pdf.

Marcuse:1966:CBG

[Mar66c] D. Marcuse. Comparison between a gas lens and its equivalent thin lens. The Bell System Technical Journal, 45(8):1339–1344, October 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-8-1339.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-8-1339.pdf.

Marcuse:1966:DFP

[Mar66d] D. Marcuse. Deformation of fields propagating through gas lenses. The Bell System Technical Journal, 45(8):1345–1368, October 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol45/bstj45-8-1345.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol45-1966/
                                        articles/bstj45-8-1345.pdf.

Marcuse:1966:PLR

[Mar66e] D. Marcuse. Physical limitations on ray oscillation suppressors. The Bell System Technical Journal, 45(5):743–751, May/ June 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol45/bstj45-5-743.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol45-1966/
                                        articles/bstj45-5-743.pdf.

Marcatili:1967:ERR

[Mar67a] E. A. J. Marcatili. Effect of redirectors, refocusers, and mode filters on light transmission through aberrated and misaligned lenses. The Bell System Technical Journal, 46 (8):1733–1752, October 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol46/bstj46-8-1733.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-8-1733.pdf.

Marcatili:1967:AWO

[Mar67b] E. A. J. Marcatili. Off-axis wave-optics transmission in a lens-like medium with aberration. The Bell System Technical Journal, 46(1):149–166, January 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-1-149.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-1-149.pdf.

Marlow:1967:NLT

[Mar67c] N. A. Marlow. A normal limit theorem for power sums of independent random variables. The Bell System Technical Journal, 46(9):2081–2089, November 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-9-2081.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-9-2081.pdf.

Marino:1968:ISL

[Mar68] P. J. Marino. Identification and synthesis of linear sequential machines. The Bell System Technical Journal, 47(3):343–384, March 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-3-343.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-3-343.pdf.

Marcatili:1969:BOD

[Mar69a] E. A. J. Marcatili. Bends in optical dielectric guides. The Bell System Technical Journal, 48(7):2103–2132, September 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-7-2103.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-7-2103.pdf.

Marcatili:1969:DRW

[Mar69b] E. A. J. Marcatili. Dielectric rectangular waveguide and directional coupler for integrated optics. The Bell System Technical Journal, 48(7): 2071–2102, September 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-7-2071.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-7-2071.pdf.

Marcuse:1969:MCCa

[Mar69c] Dietrich Marcuse. Mode conversation caused by surface imperfections of a dielectric slab waveguide. The Bell System Technical Journal, 48(10):3187– 3215, December 1969. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol48/bstj48-10-3187.pdf;
http://www.alcatel-lucent.
com/bstj/vol48-1969/articles/
bstj48-10-3187.pdf.

Marcuse:1969:RLD

[Mar69d] Dietrich Marcuse. Radiation losses of dielectric waveguides in terms of the power spectrum of the wall distortion function. The Bell System Technical Journal, 48(10):3233–3242, December 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

bell-labs.com/BSTJ/images/
Vol48/bstj48-10-3233.pdf;
http://www.alcatel-lucent.
com/bstj/vol48-1969/articles/
bstj48-10-3233.pdf.

Mathews:1961:ACM

[Mat61] Max V. Mathews. An acoustic compiler for music and psychological stimuli. The Bell System Technical Journal, 40(3):677–694, May 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-3-677.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-3-677.pdf.

Maurer:1967:CAM

[Mau67] D. W. Maurer. Cathode activity measurement: A modification of the dip test. The Bell System Technical Journal, 46(10): 2363–2374, December 1967. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol46/bstj46-10-2363.pdf;

                                        http://www.alcatel-lucent.
                                        com/bstj/vol46-1967/articles/
                                        bstj46-10-2363.pdf.

Mayo:1962:BRP

[May62] J. S. Mayo. A bipolar repeater for pulse code modulation signals. The Bell System Technical Journal, 41(1):25–97, January 1962. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol41/bstj41-1-
                                        25.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol41-1962/
                                        articles/bstj41-1-25.pdf.

Mayo:1965:EMP

[May65] J. S. Mayo. Experimental 224 Mb/s PCM terminals. The Bell System Technical Journal, 44(9):1813–1841, November 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol44/bstj44-9-1813.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol44-1965/
                                        articles/bstj44-9-1813.pdf.

Mazo:1968:QND

[Maz68]  James E. Mazo.  Quantiz-

ing noise and data transmission. The Bell System Technical Journal, 47(8):1737–1753, October 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol47/bstj47-8-1737.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol47-1968/
articles/bstj47-8-1737.pdf.

Marcatili:1961:BSF

[MB61] E. A. Marcatili and D. L.

Bisbee. Band-splitting filter. The Bell System Technical Journal, 40(1):197–212, January 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-1-197.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-1-197.pdf.

McCluskey:1962:MSS

[McC62] E. J. McCluskey, Jr. Minimumstate sequential circuits for a restricted class of incompletely specified flow tables. The Bell System Technical Journal, 41(6):1759–1768, November 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol41/bstj41-6-1759.
pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-6-1759.pdf.

McCumber:1965:ESC

[McC65] D. E. McCumber. Eigenmodes of a symmetric cylindrical confocal laser resonator and their perturbation by output-coupling apertures. The Bell System Technical Journal, 44(2):333–363, February 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol44/bstj44-2-333.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol44-1965/
                                        articles/bstj44-2-333.pdf.

McCullough:1968:BRC

[McC68] Richard H. McCullough. The binary regenerative channel. The Bell System Technical Journal, 47(8):1713–1735, October 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol47/bstj47-8-1713.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-8-1713.pdf.

McCumber:1969:EAC

[McC69] D. E. McCumber. Eigenmodes of an asymmetric cylindrical confocal laser resonator with a single output-coupling aperture. The Bell System Technical Journal, 48(6):1919–1936, July/ August 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-6-1919.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-6-1919.pdf.

McDonald:1966:SNI

[McD66] R. A. McDonald. Signal-to-noise and idle channel performance of differential pulse code modulation systems — particular applications to voice signals. The Bell System Technical Journal, 45(7):1123–1151, September 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-7-1123.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-7-1123.pdf.

McKenna:1967:EPD

[McK67] J. McKenna. The excitation of planar dielectric waveguide at p–n junctions, I. The Bell System Technical Journal, 46(7):1491–1526, September 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-7-1491.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-7-1491.pdf.

McMillan:1969:CSW

[McM69] Brockway McMillan. Communication systems which minimize coding noise. The Bell System Technical Journal, 48(9):3091–3112, November 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-9-3091.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-9-3091.pdf.

Marcuse:1969:MCCb

[MD69] Dietrich Marcuse and Richard M.

Derosier. Mode conversation caused by diameter changes of a round dielectric waveguide. The Bell System Technical Journal, 48(10):3217–3232, December 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

bell-labs.com/BSTJ/images/
Vol48/bstj48-10-3217.pdf;
http://www.alcatel-lucent.
com/bstj/vol48-1969/articles/
bstj48-10-3217.pdf.

Meadows:1962:SSL

[Mea62] H. E. Meadows. Solution of systems of linear ordinary differential equations with periodic coefficients. The Bell System Technical Journal, 41(4):1275–1294, July 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol41/bstj41-4-1275.
pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-4-1275.pdf.

Meacham:1968:NIB

[Mea68]  L. A. Meacham.  Negative

impedance boosting. The Bell System Technical Journal, 47(6):1019–1041, July/ August 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-6-1019.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-6-1019.pdf.

Melzak:1968:CCC

[Mel68] Z. A. Melzak. On a class of configuration and coincidence problems. The Bell System Technical Journal, 47(6):1105–1129, July/ August 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-6-1105.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-6-1105.pdf.

Miller:1963:NSI

[Mil63] S. E. Miller. The nature of and system inferences of delay distortion due to mode conversion in multimode transmission systems. The Bell System Technical Journal, 42(6):2741–2760, November 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-6-2741.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-6-2741.pdf.

Miller:1964:AGF

[Mil64a] S. E. Miller. Alternating-gradient focusing and related properties of conventional convergent lens focusing. The Bell System Technical Journal, 43(4):1741–1758, July 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-4-1741.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-4-1741.pdf.

Miller:1964:DCL

[Mil64b] S. E. Miller. Directional control in light-wave guidance. The Bell System Technical Journal, 43(4):1727–1739, July 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-4-1727.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-4-1727.pdf.

Mills:1964:CPE

[Mil64c] G. W. Mills. A comparison of permanent electrical connections. The Bell System Technical Journal, 43(3):1067–1102, May 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-3-1067.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-3-1067.pdf.

Miller:1965:LPG

[Mil65] S. E. Miller. Light propagation in generalized lens-like media. The Bell System Technical Journal, 44(9):2017–2064, November 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol44/bstj44-9-2017.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol44-1965/
articles/bstj44-9-2017.pdf.

Miller:1968:STW

[Mil68] S. E. Miller. On solutions for two waves with periodic coupling. The Bell System Technical Journal, 47(8):1801–1822, October 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-8-1801.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-8-1801.pdf.

Miller:1969:IOI

[Mil69a] Stewart E. Miller. Integrated optics: An introduction. The Bell System Technical Journal, 48(7):2059–2069, September 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-7-2059.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-7-2059.pdf.

Miller:1969:STA

[Mil69b] Stewart E. Miller. Some theory and applications of periodically coupled waves. The Bell System Technical Journal, 48(7):2189–2219, September 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-7-2189.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-7-2189.pdf.

Murray-Lasso:1968:UMT

[ML68]  M. A. Murray-Lasso.  Uni-

fied matrix theory of lumped and distributed directional couplers. The Bell System Technical Journal, 47(1):39–71, January 1968. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol47/bstj47-1-
                                        39.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-1-39.pdf.

Miller:1967:BJB

[MLW67] R. C. Miller, E. F. Labuda, and C. E. Webb. B.S.T.J. briefs: Excited level populations in high current density argon discharges. The Bell System Technical Journal, 46(1):281–284, January 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol46/bstj46-1-281.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-1-281.pdf.

Martel:1961:FRD

[MM61] H. C. Martel and M. V. Mathews. Further results on the detectability of known signals in Gaussian noise. The Bell System Technical Journal, 40(2):423–451, March 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-2-423.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-2-423.pdf.

Marcuse:1964:ATG

[MM64] D. Marcuse and S. E. Miller.

Analysis of a tubular gas lens. The Bell System Technical Journal, 43(4):1759–1782, July 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-4-1759.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-4-1759.pdf.

See erratum [Ano64k].

Misawa:1966:BJB

[MM66a] T. Misawa and L. P. Marinaccio. B.S.T.J. briefs: A 1/4-watt Si PvN X-band IMPATT (IM-Pact Avalanche Transit Time) diode. The Bell System Technical Journal, 45(6):989–991, July/ August 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-6-989.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-6-989.pdf.

Morrison:1966:EBH

[MM66b] J. A. Morrison and S. P. Morgan.

Electron beam heating of a thin film on a highly conducting substrate. The Bell System Technical Journal, 45(5):661–687, May/ June 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-5-661.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-5-661.pdf.

McKenna:1968:ESE

[MM68] James McKenna and J. A. Morrison. The effects of strain on electromagnetic modes of anisotropic dielectric waveguides at p–n junctions. The Bell System Technical Journal, 47(9):1933–1956, November 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-9-1933.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-9-1933.pdf.

Marcatili:1969:IRD

[MM69] E. A. J. Marcatili and S. E. Miller. Improved relations describing directional control in electromagnetic wave guidance. The Bell System Technical Journal, 48(7):2161–2188, September 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-7-2161.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-7-2161.pdf.

Mayo:1963:CSM

[MMW^{+}63] J. S. Mayo, H. Mann, F. J. Witt, D. S. Peck, H. K. Gummel, and W. L. Brown. The command system malfunction of the Telstar satellite. The Bell System Technical Journal, 42(4):1631–1657, July 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-4-1631.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-4-1631.pdf.

Martellotto:1964:PIC

[MOP64] N. A. Martellotto, H. Oehring, and M. C. Paull. PRO-CESS III — A compilerassembler for no. 1 ESS. The Bell System Technical Journal, 43(5):2457–2481, September 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-5-2457.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-5-2457.pdf.

Morris:1962:FAE

[Mor62a] Robert Morris. Further analysis of errors reported in ‘capabilities of the telephone network for data transmission’. The Bell System Technical Journal, 41(4):1399–1414, July 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/

                                        images/Vol41/bstj41-4-1399.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol41-1962/
                                        articles/bstj41-4-1399.pdf.

Morrison:1962:MFP

[Mor62b] J. A. Morrison. Maximization of the fundamental power in nonlinear capacitance diodes. The Bell System Technical Journal, 41(2):677–721, March 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol41/bstj41-2-677.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol41-1962/
                                        articles/bstj41-2-677.pdf.

Morgan:1964:MHL

[Mor64a] Samuel P. Morgan. Microwave heating of a luneberg lens. The Bell System Technical Journal, 43(2):659–677, March 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-2-659.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-2-659.pdf.

Morgan:1964:SLL

[Mor64b] Samuel P. Morgan. Scattering losses in a large luneberg lens due to random dielectric inhomogeneities. The Bell System Technical Journal, 43(2):679–696, March 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-2-679.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-2-679.pdf.

Morgan:1965:IAA

[Mor65] Samuel P. Morgan. Interaction of adaptive antenna arrays in an arbitrary environment. The Bell System Technical Journal, 44(1):23–47, January 1965. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol44/bstj44-1-
23.pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-1-23.pdf.

Moschytz:1965:MRF

[Mos65] G. S. Moschytz. Miniaturized RC filters using phase-locked loop. The Bell System Technical Journal, 44(5):823–870, May/ June 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-5-823.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-5-823.pdf.

Mounts:1967:LRT

[Mou67] F. W. Mounts. Low-resolution TV: An experimental digital system for evaluating bandwidth reduction techniques. The Bell System Technical Journal, 46(1):167–198, January 1967. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol46/bstj46-1-167.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-1-167.pdf.

Mounts:1969:VES

[Mou69] F. W. Mounts. A video encoding system with conditional picture-element replenishment. The Bell System Technical Journal, 48(7):2545–2554, September 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-7-2545.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-7-2545.pdf.

Maurer:1967:CMC

[MP67] D. W. Maurer and C. M. Pleass. The CPC: A medium current density, high reliability cathode. The Bell System Technical Journal, 46(10):2375–2404, December 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

                                        bell-labs.com/BSTJ/images/
                                        Vol46/bstj46-10-2375.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol46-1967/articles/
                                        bstj46-10-2375.pdf.

Mounts:1969:AIN

[MP69] F. W. Mounts and D. E. Pearson. Apparent increase in noise level when television pictures are frame-repeated. The Bell System Technical Journal, 48(3):527–539, March 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-3-527.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-3-527.pdf.

Malthaner:1960:CRL

[MR60] W. A. Malthaner and J. P.

Runyon. Controller for a remote line concentrator in a time-separation switching experiment. The Bell System Technical Journal, 39(1):59–86, January 1960. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol39/bstj39-1-
59.pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-1-59.pdf.

Mazo:1969:ROD

[MRS69] J. E. Mazo, Harrison E. Rowe, and J. Salz. Rate optimization for digital frequency modulation. The Bell System Technical Journal, 48(9):3021–3030, November 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-9-3021.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-9-3021.pdf.

Morgan:1960:PDC

[MS60] S. P. Morgan and F. M. Smits. Potential distribution and capacitance of a graded p–n junction. The Bell System Technical Journal, 39(6):1573–1602, November 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol39/bstj39-6-1573.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol39-1960/
                                        articles/bstj39-6-1573.pdf.

Miedema:1962:TQE

[MS62] H. Miedema and M. G. Schachtman. TASI quality — effect of speech detectors and interpolation. The Bell System Technical Journal, 41(4):1455–1473, July 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol41/bstj41-4-1455.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol41-1962/
                                        articles/bstj41-4-1455.pdf.

Marcatili:1964:HMD

[MS64] E. A. J. Marcatili and R. A. Schmeltzer. Hollow metallic and dielectric waveguides for long distance optical transmission and lasers. The Bell System Technical Journal, 43(4):1783–1809, July 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-4-1783.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol43-1964/
articles/bstj43-4-1783.pdf.

Mazo:1965:PEQ

[MS65] J. E. Mazo and J. Salz. Probability of errors for quadratic detectors. The Bell System Technical Journal, 44(9): 2165–2186, November 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-9-2165.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-9-2165.pdf.

Mazo:1966:TER

[MS66] J. E. Mazo and J. Salz. Theory of error rates for digital FM. The Bell System Technical Journal, 45(9):1511–1535, November 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-9-1511.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-9-1511.pdf.

Mann:1962:CCE

[MSV62] H. Mann, H. M. Straube, and C. P. Villars. A companded coder for an experimental PCM terminal. The Bell System Technical Journal, 41(1):173–226, January 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol41/bstj41-1-173.

                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol41-1962/
                                        articles/bstj41-1-173.pdf.

Murphy:1968:DTS

[Mur68] Bernard T. Murphy. Diode and transistor self-analogues for circuit analysis. The Bell System Technical Journal, 47(4):487–502, April 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol47/bstj47-4-487.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-4-487.pdf.

Musa:1963:DSF

[Mus63] J. D. Musa. Discrete smoothing filters for correlated noise. The Bell System Technical Journal, 42(5):2121–2151, September 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-5-2121.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-5-2121.pdf.

Mitchell:1961:MPR

[MV61] Doren Mitchell and K. G. Van Wynen. A 150-mc personal radio signaling system. The Bell System Technical Journal, 40(5):1239–1257, September 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/

images/Vol40/bstj40-5-1239.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-5-1239.pdf.

Mitchell:1967:SER

[MW67] J. P. Mitchell and D. K. Wilson. Surface effects of radiation on semiconductor devices. The Bell System Technical Journal, 46(1):1–80, January 1967. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol46/bstj46-1-
1.pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-1-1.pdf.

Nagelberg:1967:PPC

[Nag67] Elliott R. Nagelberg. Phase progression in conical waveguides. The Bell System Technical Journal, 46(10):2453–2466, December 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

bell-labs.com/BSTJ/images/
Vol46/bstj46-10-2453.pdf;
http://www.alcatel-lucent.
com/bstj/vol46-1967/articles/
bstj46-10-2453.pdf.

Nagelberg:1968:RPD

[Nag68] Elliott R. Nagelberg. Radiating properties of dielectric covered apertures. The Bell System Technical Journal, 47(10): 2169–2181, December 1968. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.

                                        bell-labs.com/BSTJ/images/
                                        Vol47/bstj47-10-2169.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol47-1968/articles/
                                        bstj47-10-2169.pdf.

Narayanan:1967:TDA

[Nar67] S. Narayanan. Transistor distortion analysis using Volterra series representation. The Bell System Technical Journal, 46(5):991–1024, May/June 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol46/bstj46-5-991.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-5-991.pdf.

Nasell:1964:SNL

[Nas64] Ingemar Nasell. The 1962 survey of noise and loss on toll connections. The Bell System Technical Journal, 43(2):697–718, March 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-2-697.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-2-697.pdf.

Nasell:1967:SPP

[Nas67] Ingemar Nasell. Some properties of power sums of truncated normal random variables. The Bell System Technical Journal, 46(9):2091–2110, November 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-9-2091.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-9-2091.pdf.

Nasell:1968:STC

[Nas68]  I. Nasell.  Some transmis-

sion characteristics of Bell System toll connections. The Bell System Technical Journal, 47(6):1001–1018, July/ August 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-6-1001.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-6-1001.pdf.

Nelson:1965:ACS

[NC65] W. L. Nelson and W. J.

Cole. Autotrack control systems for antenna mounts with non-orthogonal axes. The Bell System Technical Journal, 44(7):1367–1403, September 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-7-1367.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-7-1367.pdf.

Nasell:1968:TPB

[NEH68] I. Nasell, C. R. Ellison, Jr., and R. Holmstrom. The transmission performance of Bell System intertoll trunks. The Bell System Technical Journal, 47(8):1561–1613, October 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol47/bstj47-8-1561.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-8-1561.pdf.

Nelson:1963:PLL

[Nel63]  W. L. Nelson.  Phase-lock

loop design for coherent angleerror detection in the Telstar satellite tracking system. The Bell System Technical Journal, 42(5):1941–1975, September 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-5-1941.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-5-1941.pdf.

Nelson:1964:DLD

[Nel64] T. J. Nelson. Digital light deflection. The Bell System Technical Journal, 43(3):821–845, May 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-3-821.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-3-821.pdf.

Nicollian:1967:SSI

[NG67] E. H. Nicollian and A. Goetzberger. The Si–SiO_{2} interface — electrical properties as determined by the metal-insulator-silicon conductance technique. The Bell System Technical Journal, 46(6):1055–1133, July/ August 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-6-1055.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-6-1055.pdf.

Newman:1964:EEC

[NM64] D. J. Newman and S. P.

Morgan. Existence of eigenvalues of a class of integral equations arising in laser theory. The Bell System Technical Journal, 43(1):113–126, January 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-1-113.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-1-113.pdf.

Nicollian:1967:QTT

[NM67] E. H. Nicollian and H. Melchior.

A quantitative theory of 1/f type noise due to interface states in thermally oxidized silicon. The Bell System Technical Journal, 46(9):2019–2033, November 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-9-2019.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-9-2019.pdf.

Nagelberg:1965:MCC

[NS65] E. R. Nagelberg and J. Shefer. Mode conversion in circular waveguides. The Bell System Technical Journal, 44(7):1321–1338, September 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol44/bstj44-7-1321.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol44-1965/
                                        articles/bstj44-7-1321.pdf.

Ohm:1961:PER

[Ohm61] E. A. Ohm. Project Echo: Receiving system. The Bell System Technical Journal, 40(4): 1065–1094, July 1961. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol40/bstj40-4-1065.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol40-1961/articles/
                                        bstj40-4-1065.pdf.

Osborne:1969:LNR

[OKS69] T. L. Osborne, L. U. Kibler, and W. W. Snell. Low noise receiving downconverter. The Bell System Technical Journal, 48(6):1651–1663, July/ August 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-6-1651.

pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-6-1651.pdf.

ONeal:1966:DMQ

[O’N66a] J. B. O’Neal, Jr. Delta modulation quantizing noise analytical and computer simulation results for Gaussian and television input signals. The Bell System Technical Journal, 45(1):117–141, January 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-1-117.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-1-117.pdf.

ONeal:1966:PQS

[O’N66b] J. B. O’Neal, Jr. Predictive quantizing systems (differential pulse code modulation) for the transmission of television signals. The Bell System Technical Journal, 45(5):689–721, May/ June 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-5-689.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-5-689.pdf.

ONeill:1969:MTT

[O’N69] J. F. O’Neill. Multiplex touchtone detection using time speedup. The Bell System Technical Journal, 48(1):249–254, January 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-1-249.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-1-249.pdf.

Ohm:1963:RSC

[OS63] E. A. Ohm and W. W. Snell. A radiometer for a space communications receiver. The Bell System Technical Journal, 42(5):2047–2080, September 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-5-2047.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-5-2047.pdf.

Osborne:1969:DEB

[Osb69] T. L. Osborne. Design of efficient broadband varactor upconverters. The Bell System Technical Journal, 48(6):1623–1649, July/ August 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-6-1623.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-6-1623.pdf.

Ossanna:1964:MMR

[Oss64] Joseph F. Ossanna, Jr. A model for mobile radio fading due to building reflections: Theoretical and experimental fading waveform power spectra. The Bell System Technical Journal, 43(6):2935–2971, November 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-6-2935.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-6-2935.pdf.

Otterman:1962:GSD

[Ott62] Joseph Otterman. Grade of service of direct traffic mixed with store-and-forward traffic. The Bell System Technical Journal, 41(4):1415–1437, July 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol41/bstj41-4-1415.
pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-4-1415.pdf.

Parkin:1968:SVM

[Par68a] R. E. Parkin. A state variable method of circuit analysis based on a nodal approach. The Bell System Technical Journal, 47(9):1957–1970, November 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-9-1957.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-9-1957.pdf.

Parkin:1968:DAD

[Par68b] Robert E. Parkin. The design of active distributed filters. The Bell System Technical Journal, 47(5):605–622, May/ June 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol47/bstj47-5-605.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-5-605.pdf.

Paull:1962:RCN

[Pau62]  M. C. Paull.  Reswitching

of connection networks. The Bell System Technical Journal, 41(3):833–855, May 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol41/bstj41-3-833.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol41-1962/
                                        articles/bstj41-3-833.pdf.

Peck:1963:SER

[PBBS63] D. S. Peck, R. R. Blair, W. L. Brown, and F. M. Smits. Surface effects of radiation on transistors. The Bell System Technical Journal, 42(1):95–129, January 1963. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol42/bstj42-1-
                                        95.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-1-95.pdf.

Pantell:1964:TDT

[PDS64] R. H. Pantell, M. DiDomenico, Jr., and O. Svelto. The theory of direct transitions in semiconductors. The Bell System Technical Journal, 43(2):805–816, March 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-2-805.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-2-805.pdf.

Prabhu:1969:IIC

[PE69] V. K. Prabhu and L. H. Enloe. Interchannel interference considerations in anglemodulated systems. The Bell System Technical Journal, 48(7):2333–2358, September 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-7-2333.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-7-2333.pdf.

Peek:1961:MPC

[Pee61]  R. L. Peek, Jr.  Magne-

tization and pull characteristics of mating magnetic reeds. The Bell System Technical Journal, 40(2):523–546, March 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-2-523.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-2-523.pdf.

Peek:1963:FVP

[Pee63] R. L. Peek, Jr. Flexural vibrations of a propped cantilever. The Bell System Technical Journal, 42(3):609–636, May 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-3-609.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-3-609.pdf.

Pfahnl:1963:PFD

[Pfa63] Arnold Pfahnl. Properties of fastdecay cathode-ray tube phosphors. The Bell System Technical Journal, 42(1):181–201, January 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-1-181.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-1-181.pdf.

Pfahnl:1964:BJB

[Pfa64]  A. Pfahnl.  B.S.T.J. briefs:

Quantum efficiency of the green and red electroluminescence of GaP. The Bell System Technical Journal, 43(1):333–334, January 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-1-333.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-1-333.pdf.

Pierce:1966:IRC

[Pie66] J. R. Pierce. Information rate of a coaxial cable with various modulation systems. The Bell System Technical Journal, 45(8):1197–1207, October 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-8-1197.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-8-1197.pdf.

Pierce:1968:LRC

[Pie68]  William H. Pierce.  Linear-

real codes and coders. The Bell System Technical Journal, 47(6):1065–1097, July/ August 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-6-1065.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-6-1065.pdf.

Pierce:1969:SDN

[Pie69]  J. R. Pierce.  Synchroniz-

ing digital networks. The Bell System Technical Journal, 48(3):615–636, March 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-3-615.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-3-615.pdf.

Pilc:1968:TDS

[Pil68]  R. J. Pilc.  The transmis-

sion distortion of a source as a function of the encoding block length. The Bell System Technical Journal, 47(6):827–885, July/ August 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol47/bstj47-6-827.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-6-827.pdf.

Pilc:1969:OLM

[Pil69] Randolph J. Pilc. The optimum linear modulator for a Gaussian source used with a Gaussian channel. The Bell System Technical Journal, 48(9):3075–3089, November 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-9-3075.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-9-3075.pdf.

Pope:1968:PRG

[Pop68] D. L. Pope. Parametric representation of ground antennas for communication systems studies. The Bell System Technical Journal, 47(10):2145–2168, December 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

                                        bell-labs.com/BSTJ/images/
                                        Vol47/bstj47-10-2145.pdf;

http://www.alcatel-lucent.
com/bstj/vol47-1968/articles/
bstj47-10-2145.pdf.

Prabhu:1969:SDB

[PR69] V. K. Prabhu and H. E. Rowe.

Spectral density bounds of a PM wave. The Bell System Technical Journal, 48(3):769–811, March 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-3-769.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-3-769.pdf.

Prabhu:1967:SCL

[Pra67] V. K. Prabhu. Stability considerations in lossless varactor frequency multipliers. The Bell System Technical Journal, 46(9):2035–2060, November 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-9-2035.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-9-2035.pdf.

Prabhu:1968:SCB

[Pra68] V. K. Prabhu. Some considerations of broadband noise performance of optical heterodyne receivers. The Bell System Technical Journal, 47(3):429–457, March 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/

                                        images/Vol47/bstj47-3-429.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-3-429.pdf.

Prabhu:1969:ERC

[Pra69] V. K. Prabhu. Error rate considerations for coherent phase-shift keyed systems with co-channel interference. The Bell System Technical Journal, 48(3):743–767, March 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-3-743.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-3-743.pdf.

Protonotarios:1967:BJB

[Pro67a] E. N. Protonotarios. B.S.T.J. briefs: Realizability conditions for the impendance function of the lossless tapered transmission line — A critique. The Bell System Technical Journal, 46(5):1047–1054, May/ June 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol46/bstj46-5-1047.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-5-1047.pdf.

Protonotarios:1967:SON

[Pro67b] E. N. Protonotarios. Slope overload noise in differential pulse code modulation systems. The Bell System Technical Journal, 46(9):2119–2161, November 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-9-2119.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-9-2119.pdf.

Priebe:1969:NES

[PS69] H. F. Priebe, Jr. and E. L. Seley. No. 101 ESS 4A switch unit. The Bell System Technical Journal, 48(2):477–495, February 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-2-477.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-2-477.pdf.

Peck:1963:CDC

[PW63] D. S. Peck and M. C. Wooley. Component design, construction and evaluation for satellites. The Bell System Technical Journal, 42(4):1665–1686, July 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-4-1665.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-4-1665.pdf.

Pasternack:1968:ASD

[PW68] G. Pasternack and R. L. Whalin.

Analysis and synthesis of a digital phase-locked loop for FM demodulation. The Bell System Technical Journal, 47(10): 2207–2237, December 1968. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol47/bstj47-10-2207.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol47-1968/articles/
                                        bstj47-10-2207.pdf.

Paul:1963:PGA

[PWY63] B. Paul, J. W. West, and E. Y. Yu. A passive gravitational attitude control system for satellites. The Bell System Technical Journal, 42(5):2195–2238, September 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-5-2195.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-5-2195.pdf.

Rabiner:1968:SSR

[Rab68] Lawrence Rabiner. Speech synthesis by rule: An acoustic domain approach. The Bell System Technical Journal, 47(1):17– 37, January 1968. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol47/bstj47-1-
                                        17.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-1-17.pdf.

Rado:1962:NCF

[Rad62] T. Radó. On non-computable functions. The Bell System Technical Journal, 41(3):877–884, May 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol41/bstj41-3-877.
pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-3-877.pdf.

Rainal:1965:BJB

[Rai65] A. J. Rainal. B.S.T.J. brief: Axiscrossing intervals of Rayleigh processes. The Bell System Technical Journal, 44(6):1219–1224, July/ August 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-6-1219.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-6-1219.pdf.

Rainal:1966:DFA

[Rai66a]  A. J. Rainal.  Duration of

fades associated with radar clutter. The Bell System Technical Journal, 45(8):1285–1298, October 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-8-1285.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-8-1285.pdf.

Rainal:1966:PPD

[Rai66b] A. J. Rainal. Phase principle for detecting narrow band Gaussian signals. The Bell System Technical Journal, 45(1):143–148, January 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol45/bstj45-1-143.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol45-1966/
                                        articles/bstj45-1-143.pdf.

Rainal:1967:ACP

[Rai67a] A. J. Rainal. Axis-crossings of the phase of sine wave plus noise. The Bell System Technical Journal, 46(4):737–754, April 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol46/bstj46-4-737.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-4-737.pdf.

Rainal:1967:BJB

[Rai67b] A. J. Rainal. B.S.T.J. briefs: Axis-crossing intervals of sine wave plus noise. The Bell System Technical Journal, 46 (7):1655–1658, September 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol46/bstj46-7-1655.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-7-1655.pdf.

Rainal:1968:BJB

[Rai68a] A. J. Rainal. B.S.T.J. briefs: Approximate and exact results concerning zeros of Gaussiannoise. The Bell System Technical Journal, 47(3):461–463, March 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-3-461.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-3-461.pdf.

Rainal:1968:FSP

[Rai68b] A. J. Rainal. First and second passage times of sine wave plus noise. The Bell System Technical Journal, 47(10):2239– 2258, December 1968. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol47/bstj47-10-2239.pdf;
http://www.alcatel-lucent.
com/bstj/vol47-1968/articles/
bstj47-10-2239.pdf.

Rainal:1968:PPM

[Rai68c] A. J. Rainal. Phase principle for measuring location or spectral shape of a discrete radio source. The Bell System Technical Journal, 47(3):415–428, March 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-3-415.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-3-415.pdf.

Rao:1969:MST

[Rao69] T. N. Rao. Minimal synthesis of two-variable reactance matrices. The Bell System Technical Journal, 48(1):163–199, January 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-1-163.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-1-163.pdf.

Rappeport:1964:DCS

[Rap64]  M. A. Rappeport.  Digital

computer simulation of a fourphase data transmission system. The Bell System Technical Journal, 43(3):927–964, May 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-3-927.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-3-927.pdf.

Rappeport:1965:FPD

[Rap65]  M. A. Rappeport.  Four-

phase data systems in combined delay distortion, Gaussian noise, and impulse noise. The Bell System Technical Journal, 44(2):253–260, February 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-2-253.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-2-253.pdf.

Rappaport:1967:CRR

[Rap67] Stephen S. Rappaport. Communications and radar receiver gains for minimum average cost of excluding randomly fluctuating signals in random noise. The Bell System Technical Journal, 46(8):1753–1773, October 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-8-1753.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-8-1753.pdf.

Rappaport:1968:GCD

[Rap68] Stephen S. Rappaport. Gain control for diversity receivers. The Bell System Technical Journal, 47(9):2015–2027, November 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-9-2015.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-9-2015.pdf.

Rack:1964:TMS

[RB64] A. J. Rack and M. R. Biazzo.

A technique for measuring small optical loss using an oscillating spherical mirror interferometer. The Bell System Technical Journal, 43(4):1563–1579, July 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-4-1563.

                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-4-1563.pdf.

Ray-Chaudhuri:1961:CMR

[RC61]  D. K. Ray-Chaudhuri.  On

the construction of minimally redundant reliable system designs. The Bell System Technical Journal, 40(2):595–611, March 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol40/bstj40-2-595.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol40-1961/
                                        articles/bstj40-2-595.pdf.

Renau:1965:MEB

[RC65] Jacques Renau and James A. Collinson. Measurements of electromagnetic back-scattering from known, rough surfaces. The Bell System Technical Journal, 44 (10):2203–2226, December 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol44/bstj44-10-2203.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol44-1965/articles/
                                        bstj44-10-2203.pdf.

Riney:1962:TSB

[RE62] T. D. Riney and J. W. Elek. Thermoelastic stresses in balanced and unbalanced seals. The Bell System Technical Journal, 41(5):1519–1536, September 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol41/bstj41-5-1519.
pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-5-1519.pdf.

Rosenzweig:1963:SCD

[RGS63] W. Rosenzweig, H. K. Gummel, and F. M. Smits. Solar cell degradation under 1-Mev electron bombardment. The Bell System Technical Journal, 42(2):399–414, March 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-2-399.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-2-399.pdf.

Ross:1960:ST

[RH60] Ian C. Ross and Frank Harary.

The square of a tree. The Bell System Technical Journal, 39(3):641–647, May 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-3-641.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-3-641.pdf.

Rice:1962:IBP

[Ric62a]  S. O. Rice.  Intervals be-

tween periods of no service in certain satellite communication systems—analogy with a traffic system. The Bell System Technical Journal, 41 (5):1671–1690, September 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol41/bstj41-5-1671.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol41-1962/
                                        articles/bstj41-5-1671.pdf.

Rice:1962:SSSb

[Ric62b] S. O. Rice. Single server systems – II. busy periods. The Bell System Technical Journal, 41(1):279–310, January 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol41/bstj41-1-279.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol41-1962/
                                        articles/bstj41-1-279.pdf.

Rice:1962:SSSa

[Ric62c] S. O. Rice. Single server systems. I. Relations between some averages. The Bell System Technical Journal, 41(1):269–278, January 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol41/bstj41-1-269.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol41-1962/
                                        articles/bstj41-1-269.pdf.

Rice:1963:IOI

[Ric63] S. O. Rice. Innage and outage intervals in transmission systems composed of links. The Bell System Technical Journal, 42(5):2267–2283, September 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-5-2267.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-5-2267.pdf.

Rice:1968:UAE

[Ric68]  Stephen O. Rice.  Uniform

asymptotic expansions for saddle point integrals—application to a probability distribution occurring in noise theory. The Bell System Technical Journal, 47(9):1971–2013, November 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-9-1971.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-9-1971.pdf.

Rice:1969:STO

[Ric69]  S. O. Rice.  Second and

third order modulation terms in the distortion produced when modulated FM waves are filtered. The Bell System Technical Journal, 48(1):87–141, January 1969. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol48/bstj48-1-
87.pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-1-87.pdf.

Rigrod:1965:ORR

[Rig65]  W. W. Rigrod.  The op-

tical ring resonator. The Bell System Technical Journal, 44(5):907–916, May/June 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol44/bstj44-5-907.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol44-1965/
                                        articles/bstj44-5-907.pdf.

Riordan:1961:DLC

[Rio61]  John Riordan.  Delays for

last-come first-served service and the busy period. The Bell System Technical Journal, 40(3):785–793, May 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol40/bstj40-3-785.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol40-1961/
                                        articles/bstj40-3-785.pdf.

Ruthroff:1961:PES

[RJ61] Clyde L. Ruthroff and William C. Jakes, Jr. Project Echo: System calculations. The Bell System Technical Journal, 40(4):1029–1039, July 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol40/bstj40-4-1029.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol40-1961/
                                        articles/bstj40-4-1029.pdf.

Riesz:1963:SED

[RK63] R. R. Riesz and E. T. Klemmer. Subjective evaluation of delay and echo suppressors in telephone communicaions. The Bell System Technical Journal, 42(6):2919–2941, November 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-6-2919.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-6-2919.pdf.

Roberts:1962:SPL

[Rob62]  S. W. Roberts.  Scheduling

of pole line inspections. The Bell System Technical Journal, 41(6):1737–1758, November 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol41/bstj41-6-1737.
pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-6-1737.pdf.

Robertson:1967:OCL

[Rob67a] G. H. Robertson. Operating characteristics for a linear detector of CW signals in narrow-bandGaussian noise. The Bell System Technical Journal, 46(4):755–774, April 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-4-755.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-4-755.pdf.

Robinson:1967:NOC

[Rob67b] F. N. H. Robinson. Nonlinear optical coefficients. The Bell System Technical Journal, 46(5):913–956, May/June 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol46/bstj46-5-913.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-5-913.pdf.

Robertson:1968:PDP

[Rob68]  George H. Robertson.  Per-

formance degradation by postdetector nonlinearities. The Bell System Technical Journal, 47(3):407–414, March 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol47/bstj47-3-407.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-3-407.pdf.

Robertson:1969:CNC

[Rob69a] G. H. Robertson. Computation of the noncentral chi-square distribution. The Bell System Technical Journal, 48(1):201–207, January 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-1-201.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol48-1969/
articles/bstj48-1-201.pdf.

Robertson:1969:CSQ

[Rob69b] G. H. Robertson. Computer study of quantizer output spectra. The Bell System Technical Journal, 48(7): 2391–2403, September 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-7-2391.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-7-2391.pdf.

Ruthroff:1969:SHR

[ROB69c] C. L. Ruthroff, T. L. Osborne, and W. F. Bodtmann. Short hop radio system experiment. The Bell System Technical Journal, 48(6):1577–1604, July/ August 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-6-1577.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-6-1577.pdf.

Rosenzweig:1962:DLM

[Ros62a] W. Rosenzweig. Diffusion length measurement by means of ionizing radiation. The Bell System Technical Journal, 41(5):1573–1588, September 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/

                                        images/Vol41/bstj41-5-1573.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol41-1962/
                                        articles/bstj41-5-1573.pdf.

Ross:1962:RCC

[Ros62b] I. M. Ross. Reliability of components for communication satellites. The Bell System Technical Journal, 41(2):635–662, March 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol41/bstj41-2-635.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol41-1962/
                                        articles/bstj41-2-635.pdf.

Rosenbaum:1969:PEP

[Ros69] Arnold S. Rosenbaum. PSK error performance with Gaussian noise and interference. The Bell System Technical Journal, 48(2):413–442, February 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-2-413.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-2-413.pdf.

Rowe:1962:ASC

[Row62] H. E. Rowe. Approximate solutions for the coupled line equations. The Bell System Technical Journal, 41(3):1011–1029, May 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/

images/Vol41/bstj41-3-1011.
pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-3-1011.pdf.

Rowe:1964:IAT

[Row64a] H. E. Rowe. Imperfections in active transmission lines. The Bell System Technical Journal, 43(1):261–292, January 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-1-261.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-1-261.pdf.

Rowe:1964:SAT

[Row64b] H. E. Rowe. Stability of active transmission lines with arbitrary imperfections. The Bell System Technical Journal, 43(1):293–328, January 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-1-293.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-1-293.pdf.

Rowe:1968:ESS

[RP68] Harrison E. Rowe and Arno A.

Penzias. Efficient spacing of synchronous communication satellites. The Bell System Technical Journal, 47(10):2379– 2433, December 1968. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol47/bstj47-10-2379.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol47-1968/articles/
                                        bstj47-10-2379.pdf.

Rinehart:1962:CSP

[RR62] J. D. Rinehart and M. F. Robbins. Characteristics of the service provided by communications satellites in uncontrolled orbits. The Bell System Technical Journal, 41(5):1621–1670, September 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol41/bstj41-5-1621.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol41-1962/
                                        articles/bstj41-5-1621.pdf.

Rabiner:1969:CTA

[RSR69] Lawrence R. Rabiner, Ronald W. Schafer, and Charles M. Rader. The chirp z-transform algorithm and its application. The Bell System Technical Journal, 48(5):1249–1292, May/ June 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-5-1249.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-5-1249.pdf.

Ruthroff:1969:IDR

[RT69] Clyde L. Ruthroff and LeRoy C. Tillotson. Interference in a dense radio network. The Bell System Technical Journal, 48(6):1727–1743, July/ August 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-6-1727.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-6-1727.pdf.

Rudin:1969:CAE

[Rud69] H. R. Rudin, Jr. A continuously adaptive equalizer for generalpurpose communication channels. The Bell System Technical Journal, 48(6):1865–1884, July/ August 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-6-1865.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-6-1865.pdf.

Ruscio:1966:PAM

[Rus66] Joseph T. Ruscio. Phase and amplitude measurements of coherent optical wavefronts. The Bell System Technical Journal, 45(9):1583–1597, November 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-9-1583.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-9-1583.pdf.

Ruthroff:1961:PEF

[Rut61] Clyde L. Ruthroff. Project Echo:

FM demodulators with negative feedback. The Bell System Technical Journal, 40(4):1149–1156, July 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-4-1149.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-4-1149.pdf.

Ruthroff:1968:ILO

[Rut68a] C. L. Ruthroff. Injection-locked-oscillator FM receiver analysis. The Bell System Technical Journal, 47(8):1653–1661, October 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-8-1653.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-8-1653.pdf.

Ruthroff:1968:CFD

[Rut68b] Clyde L. Ruthroff. Computation of FM distortion in linear networks for bandlimited periodic signals. The Bell System Technical Journal, 47(6):1043–1063, July/ August 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-6-1043.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-6-1043.pdf.

Rowe:1962:TMW

[RW62] H. E. Rowe and W. D. Warters.

Transmission in multimode waveguide with random imperfections. The Bell System Technical Journal, 41(3):1031–1170, May 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol41/bstj41-3-1031.
pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-3-1031.pdf.

Saltzberg:1967:TRS

[Sal67]  Burton R. Saltzberg.  Tim-

ing recovery for synchronous binary data transmission. The Bell System Technical Journal, 46(3):593–622, March 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-3-593.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-3-593.pdf.

Sandberg:1960:SDP

[San60] I. W. Sandberg. Synthesis of driving-point impedances with active RC networks. The Bell System Technical Journal, 39(4):947–962, July 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-4-947.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol39-1960/
                                        articles/bstj39-4-947.pdf.

Sandberg:1961:STA

[San61]  I. W. Sandberg.  Synthesis

of transformerless active N-port networks. The Bell System Technical Journal, 40(3):761–783, May 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol40/bstj40-3-761.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol40-1961/
                                        articles/bstj40-3-761.pdf.

Sandberg:1962:RMS

[San62a] I. W. Sandberg. The realizability of multiport structures obtained by imbedding a tunnel diode in a lossless reciprocal network. The Bell System Technical Journal, 41(3):857–876, May 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol41/bstj41-3-857.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol41-1962/
                                        articles/bstj41-3-857.pdf.

Sandberg:1962:TDL

[San62b] I. W. Sandberg. The tunnel diode as a linear network element. The Bell System Technical Journal, 41(5):1537–1556, September 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol41/bstj41-5-1537.

pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-5-1537.pdf.

Sandberg:1963:PSS

[San63a] I. W. Sandberg. On the properties of some systems that distort signals. I. The Bell System Technical Journal, 42(5):2033–2046, September 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-5-2033.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-5-2033.pdf.

Sandberg:1963:TLM

[San63b] I. W. Sandberg. On the theory of linear multi-loop feedback systems. The Bell System Technical Journal, 42(2):355–382, March 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-2-355.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-2-355.pdf.

See errata [Ano64j].

Sandberg:1963:SDN

[San63c] I. W. Sandberg. Signal distortion in nonlinear feedback systems. The Bell System Technical Journal, 42(6):2533–2550, November 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/

                                        images/Vol42/bstj42-6-2533.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-6-2533.pdf.

Sandberg:1964:BJBa

[San64a] I. W. Sandberg. B.S.T.J. briefs: A condition for the L8-stability of feedback systems containing a single time-varying nonlinear element. The Bell System Technical Journal, 43(4):1815–1817, July 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-4-1815.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-4-1815.pdf.

Sandberg:1964:BJBb

[San64b] I. W. Sandberg. B.S.T.J. briefs: A note on a signal recovery problem. The Bell System Technical Journal, 43(6):3065–3067, November 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-6-3065.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-6-3065.pdf.

Sandberg:1964:FDC

[San64c] I. W. Sandberg. A frequencydomain condition for the stability of feedback systems containing a single time-varying nonlinear element. The Bell System Technical Journal, 43(4):1601–1608, July 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-4-1601.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-4-1601.pdf.

Sandberg:1964:BSN

[San64d] I. W. Sandberg. On the L_{2}-boundedness of solutions of nonlinear functional equations. The Bell System Technical Journal, 43(4):1581–1599, July 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-4-1581.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-4-1581.pdf.

Sandberg:1964:PSS

[San64e] I. W. Sandberg. On the properties of some systems that distort signals — II. The Bell System Technical Journal, 43(1):91–112, January 1964. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol43/bstj43-1-
91.pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-1-91.pdf.

Sandberg:1964:RNC

[San64f] I. W. Sandberg. On the response of nonlinear control systems to periodic input signals. The Bell System Technical Journal, 43(3):911–926, May 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-3-911.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-3-911.pdf.

Sandberg:1965:BJB

[San65a] I. W. Sandberg. B.S.T.J. brief: An observation concerning the application of the contractionmapping fixed-point theorem, and a result concerning the norm-boundedness of solutions of nonlinear functional equations. The Bell System Technical Journal, 44(8):1809–1812, October 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-8-1809.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-8-1809.pdf.

Sandberg:1965:BSN

[San65b] I. W. Sandberg. On the boundedness of solutions of nonlinear integral equations. The Bell System Technical Journal, 44(3):439–453, March 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-3-439.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-3-439.pdf.

Sandberg:1965:SRT

[San65c] I. W. Sandberg. Some results on the theory of physical systems governed by nonlinear functional equations. The Bell System Technical Journal, 44(5):871–898, May/June 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-5-871.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-5-871.pdf.

Sandberg:1965:SSR

[San65d] I. W. Sandberg. Some stability results related to those of V. M. Popov. The Bell System Technical Journal, 44 (9):2133–2148, November 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-9-2133.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-9-2133.pdf.

Sandberg:1966:BJB

[San66a] I. W. Sandberg. B.S.T.J.

briefs: A note on a type of optimization problem that arises in communication theory. The Bell System Technical Journal, 45(5):761–764, May/ June 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-5-761.

                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol45-1966/
                                        articles/bstj45-5-761.pdf.

See erratum [Ano66l].

Sandberg:1966:SCC

[San66b] I. W. Sandberg. On the sensitivity of channel capacity for the Gaussian bandlimited channel. The Bell System Technical Journal, 45(9):1475–1492, November 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol45/bstj45-9-1475.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol45-1966/
                                        articles/bstj45-9-1475.pdf.

Sandberg:1967:FPR

[San67a] I. W. Sandberg. Floatingpoint-roundoff accumulation in digital-filter realizations. The Bell System Technical Journal, 46(8):1775–1791, October 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol46/bstj46-8-1775.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-8-1775.pdf.

Sandberg:1967:SPC

[San67b] I. W. Sandberg. Some properties of a classic numerical integration formula. The Bell System Technical Journal, 46(9):2061–2080, November 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-9-2061.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-9-2061.pdf.

Sandberg:1967:TTA

[San67c] I. W. Sandberg. Two theorems on the accuracy of numerical solutions of systems of ordinary differential equations. The Bell System Technical Journal, 46(6):1243–1266, July/ August 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-6-1243.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-6-1243.pdf.

Sandberg:1969:CUW

[San69a] I. W. Sandberg. On conditions under which it is possible to synchronize digital transmission systems. The Bell System Technical Journal, 48(6):1999–2022, July/ August 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-6-1999.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-6-1999.pdf.

Sandberg:1969:STD

[San69b] I. W. Sandberg. Some theorems on the dynamic response of nonlinear transistor networks. The Bell System Technical Journal, 48(1):35–54, January 1969. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol48/bstj48-1-
                                        35.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-1-35.pdf.

Sandberg:1969:SPN

[San69c] Irwin W. Sandberg. Some properties of a nonlinear model of a system for synchronizing digital transmission networks. The Bell System Technical Journal, 48(9):2975–2997, November 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-9-2975.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-9-2975.pdf.

Savage:1966:SDC

[Sav66a] E. J. Savage. Sequential decoding — the computation problem. The Bell System Technical Journal, 45(1):149–175, January 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol45/bstj45-1-149.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol45-1966/
                                        articles/bstj45-1-149.pdf.

Savage:1966:BRB

[Sav66b] J. E. Savage. A bound on the reliability of block coding with feedback. The Bell System Technical Journal, 45(6):967–977, July/ August 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-6-967.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-6-967.pdf.

Savage:1967:SSS

[Sav67] J. E. Savage. Some simple selfsynchronizing digital data scramblers. The Bell System Technical Journal, 46(2):449–487, February 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-2-449.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-2-449.pdf.

Schafer:1961:PEM

[SB61] J. P. Schafer and R. H.

Brandt. Project Echo: 960mc, 10-kw transmitter. The Bell System Technical Journal, 40(4):1041–1064, July 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-4-1041.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-4-1041.pdf.

Sandberg:1964:PNI

[SB64] I. W. Sandberg and V. E. Beneš.

On the properties of nonlinear integral equations that arise in the theory of dynamical systems. The Bell System Technical Journal, 43(6):2839–2853, November 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-6-2839.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-6-2839.pdf.

Seidel:1968:ECH

[SBF68] H. Seidel, H. R. Beurrier, and A. N. Friedman. Error-controlled high power linear amplifiers at VHF. The Bell System Technical Journal, 47(5):651–722, May/ June 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-5-651.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-5-651.pdf.

Subramanian:1965:BJB

[SC65] M. Subramanian and J. A. Collinson. B.S.T.J. briefs: Modulation of laser beams by atmospheric turbulence. The Bell System Technical Journal, 44(3):543–546, March 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-3-543.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-3-543.pdf.

Subramanian:1967:MLB

[SC67] M. Subramanian and J. A.

Collinson. Modulation of laser beams by atmospheric turbulence — depth of modulation. The Bell System Technical Journal, 46(3):623–648, March 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-3-623.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-3-623.pdf.

Scattaglia:1963:BJB

[Sca63]  J. V. Scattaglia.  B.S.T.J.

briefs: A self-reorganizing synchronization network. The Bell System Technical Journal, 42(5):2493–2496, September 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-5-2493.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-5-2493.pdf.

Schenker:1960:PCT

[Sch60] L. Schenker. Pushbutton calling with a two-group voice-frequency code. The Bell System Technical Journal, 39(1):235–255, January 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-1-235.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol39-1960/
                                        articles/bstj39-1-235.pdf.

Scheinman:1962:MSB

[Sch62a] A. H. Scheinman. A method for simplifying Boolean functions. The Bell System Technical Journal, 41(4):1337–1346, July 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol41/bstj41-4-1337.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol41-1962/
                                        articles/bstj41-4-1337.pdf.

Schepis:1962:TSF

[Sch62b] A. J. Schepis. On the theory of shrink fits with application to waveguide pressure seals. The Bell System Technical Journal, 41(3):885–907, May 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol41/bstj41-3-885.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol41-1962/
                                        articles/bstj41-3-885.pdf.

Schneider:1963:BJB

[Sch63] M. V. Schneider. B.S.T.J. briefs: A 5-gigacycle tunnel diode oscillator with 9-milliwatt output from a single diode. The Bell System Technical Journal, 42(6):2972–2974, November 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/

images/Vol42/bstj42-6-2972.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-6-2972.pdf.

Schwartz:1965:RBS

[Sch65] M. I. Schwartz. On the reception of binary signals in the presence of a small random delay. The Bell System Technical Journal, 44(6):1123–1138, July/ August 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-6-1123.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-6-1123.pdf.

Schneider:1966:SBP

[Sch66]  M. V. Schneider.  Schot-

tky barrier photodiodes with antireflection coating. The Bell System Technical Journal, 45(9):1611–1638, November 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-9-1611.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-9-1611.pdf.

Schneider:1968:CCD

[Sch68a] H. L. Schneider. Click comparison of digital and matched filter receivers. The Bell System Technical Journal, 47(3):301–313, March 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol47/bstj47-3-301.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-3-301.pdf.

Schneider:1968:DTF

[Sch68b] H. L. Schneider. Data transmission with FSK permutation modulation. The Bell System Technical Journal, 47(6): 1131–1138, July/August 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol47/bstj47-6-1131.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-6-1131.pdf.

Schneider:1969:MLM

[Sch69a] M. V. Schneider. Microstrip lines for microwave integrated circuits. The Bell System Technical Journal, 48(5):1421–1444, May/ June 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-5-1421.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-5-1421.pdf.

Schneider:1969:RLI

[Sch69b] M. V. Schneider. Resonances loss in integrated microwave circuits. The Bell System Technical Journal, 48(7):2325–2332, September 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-7-2325.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-7-2325.pdf.

Smith:1963:POE

[SCM^{+}63] D. H. Smith, C. P. Carlson, R. J. McCune, R. E. Elicker, and R. E. Sageman. Planning, operation and external communications of the Andover Earth Station. The Bell System Technical Journal, 42(4):1383–1420, July 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-4-1383.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-4-1383.pdf.

Susen:1961:TRM

[SD61] C. P. Susen and J. J. Degan. The TH radio microwave carrier supply system. The Bell System Technical Journal, 40(6):1569–1586, November 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-6-1569.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-6-1569.pdf.

Schulz-Dubois:1964:PSG

[SD64] E. O. Schulz-Dubois. Pulse sharpening and gain saturation in traveling-wave masers. The Bell System Technical Journal, 43(2):625–658, March 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-2-625.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-2-625.pdf.

Semplak:1965:SMT

[Sem65] R. A. Semplak. 0.63μ scatter measurements from teflon and various metallic surfaces. The Bell System Technical Journal, 44(8):1659–1674, October 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol44/bstj44-8-1659.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol44-1965/
                                        articles/bstj44-8-1659.pdf.

Sproul:1961:TBR

[SG61] P. T. Sproul and H. D. Griffiths. The TH broadband radio transmitter and receiver. The Bell System Technical Journal, 40(6):1521–1568, November 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol40/bstj40-6-1521.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol40-1961/
                                        articles/bstj40-6-1521.pdf.

Shennum:1962:PLP

[SG62] R. H. Shennum and J. R. Gray. Performance limitations of a practical PCM terminal. The Bell System Technical Journal, 41(1):143–171, January 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol41/bstj41-1-143.
pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-1-143.pdf.

Smits:1967:DCS

[SG67] F. M. Smits and L. E. Gallaher. Design considerations for a semipermanent optical memory. The Bell System Technical Journal, 46(6):1267–1278, July/ August 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-6-1267.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-6-1267.pdf.

Smith:1963:SCT

[SGB^{+}63] K. D. Smith, H. K. Gummel, J. D. Bode, D. B. Cuttriss, R. J. Nielsen, and W. Rosenzweig. The solar cells and their mounting. The Bell System Technical Journal, 42(4):1765–1816, July 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-4-1765.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-4-1765.pdf.

Schneider:1969:MMW

[SGB69] M. V. Schneider, Bernard Glance, and W. F. Bodtmann. Microwave and millimeter wave hybrid integrated circuits for radio systems. The Bell System Technical Journal, 48(6):1703–1726, July/ August 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-6-1703.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-6-1703.pdf.

Shennum:1963:GDT

[SH63] R. H. Shennum and P. T. Haury. A general description of the Telstar satellite. The Bell System Technical Journal, 42(4):801–830, July 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-4-801.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-4-801.pdf.

Schroeder:1967:BJB

[SH67] M. R. Schroeder and S. L. Hanauer. B.S.T.J. briefs: Interpolation of data with continuous speech signals. The Bell System Technical Journal, 46(8):1931–1933, October 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/

images/Vol46/bstj46-8-1931.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-8-1931.pdf.

Sharpless:1963:BJB

[Sha63]  W. M. Sharpless.  B.S.T.J.

briefs: Point-contact wafer diodes for use in the 90- to 140kilomegacycle frequency range. The Bell System Technical Journal, 42(5):2496–2499, September 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-5-2496.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-5-2496.pdf.

Shnidman:1967:GNC

[Shn67] D. A. Shnidman. A generalized Nyquist criterion and an optimum linear receiver for a pulse modulation system. The Bell System Technical Journal, 46(9):2163–2177, November 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-9-2163.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-9-2163.pdf.

Simon:1967:EAR

[Sim67] M. K. Simon. Extensions to the analysis of regenerative repeaters with quantized feedback. The Bell System Technical Journal, 46(8):1831–1851, October 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol46/bstj46-8-1831.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-8-1831.pdf.

Sinden:1966:TTF

[Sin66] F. W. Sinden. Topology of thin film RC circuits. The Bell System Technical Journal, 45(9):1639–1662, November 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol45/bstj45-9-1639.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol45-1966/
                                        articles/bstj45-9-1639.pdf.

Salovey:1961:ICCa

[SK61a] R. Salovey and A. Keller. The influence of crystallization conditions on radiation effects in polyethylene — I. crystallization from dilute solution and from the melt. The Bell System Technical Journal, 40(5):1397–1408, September 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol40/bstj40-5-1397.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol40-1961/
                                        articles/bstj40-5-1397.pdf.

Salovey:1961:ICCb

[SK61b] R. Salovey and A. Keller. The influence of crystallization conditions on radiation effects in polyethylene — II. crystallization from concentrated solutions. The Bell System Technical Journal, 40(5):1409–1419, September 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-5-1409.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-5-1409.pdf.

Saltzberg:1965:DBS

[SK65] B. R. Saltzberg and L. Kurz.

Design of bandlimited signals for binary communication using simple correlation detection. The Bell System Technical Journal, 44(2):235–252, February 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-2-235.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-2-235.pdf.

Strakhov:1968:UBZ

[SK68] Nicholas A. Strakhov and Ludwik Kurz. An upper bound on the zero-crossing distribution. The Bell System Technical Journal, 47(4):529–547, April 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-4-529.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-4-529.pdf.

Semplak:1969:DNR

[SK69] R. A. Semplak and H. E. Keller. A dense network for rapid measurement of rainfall rate. The Bell System Technical Journal, 48(6):1745–1756, July/ August 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-6-1745.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-6-1745.pdf.

Skinner:1966:IMC

[Ski66] J. G. Skinner. Increasing the memory capacity of the digital light deflector by ‘color coding’. The Bell System Technical Journal, 45(4):597–608, April 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol45/bstj45-4-597.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol45-1966/
                                        articles/bstj45-4-597.pdf.

Slauson:1964:SLL

[Sla64] William C. Slauson. Substitution of laminated low-carbon steel for silicon steel in the cores for wire spring relays. The Bell System Technical Journal, 43(6):2905–2933, November 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-6-2905.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-6-2905.pdf.

Slepian:1960:SFT

[Sle60] David Slepian. Some further theory of group codes. The Bell System Technical Journal, 39(5):1219–1252, September 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-5-1219.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-5-1219.pdf.

Slepian:1962:OSB

[Sle62] David Slepian. The one-sided barrier problem for Gaussian noise. The Bell System Technical Journal, 41(2):463–501, March 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol41/bstj41-2-463.
pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-2-463.pdf.

Slepian:1963:BC

[Sle63] David Slepian. Bounds of communication. The Bell System Technical Journal, 42(3):681– 707, May 1963. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol42/bstj42-3-

                                        681.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-3-681.pdf.

Slepian:1964:PSW

[Sle64] David Slepian. Prolate spheroidal wave functions, Fourier analysis and uncertainty — IV: Extensions to many dimensions; generalized prolate spheroidal functions. The Bell System Technical Journal, 43(6): 3009–3057, November 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-6-3009.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-6-3009.pdf.

Slepian:1967:RPB

[Sle67]  David Slepian.  Restoration

of photographs blurred by image motion. The Bell System Technical Journal, 46(10):2353– 2362, December 1967. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol46/bstj46-10-2353.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol46-1967/articles/
                                        bstj46-10-2353.pdf.

Slepian:1968:GCG

[Sle68] David Slepian. Group codes for Gaussian channel. The Bell System Technical Journal, 47(4):575–602, April 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-4-575.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-4-575.pdf.

Smith:1965:JOT

[Smi65] J. W. Smith. The joint optimization of transmitted signal and receiving filter for data transmission systems. The Bell System Technical Journal, 44(10): 2363–2392, December 1965. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol44/bstj44-10-2363.pdf;
http://www.alcatel-lucent.
com/bstj/vol44-1965/articles/
bstj44-10-2363.pdf.

Smith:1968:UVS

[Smi68] James W. Smith. A unified view of synchronous data transmission system design. The Bell System Technical Journal, 47(3):273–300, March 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-3-273.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-3-273.pdf.

Sondhi:1967:AEC

[Son67] M. M. Sondhi. An adaptive echo canceller. The Bell System Technical Journal, 46(3):497–511, March 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol46/bstj46-3-497.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-3-497.pdf.

Slepian:1961:PSW

[SP61] D. Slepian and H. O. Pollak. Prolate spheroidal wave functions, Fourier analysis and uncertainty. I. The Bell System Technical Journal, 40(1):43–63, January 1961. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol40/bstj40-1-
                                        43.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol40-1961/
                                        articles/bstj40-1-43.pdf.

Sondhi:1966:BJB

[SP66] M. M. Sondhi and A. J. Presti. B.S.T.J. briefs: A self-adaptive echo canceller. The Bell System Technical Journal, 45(10): 1851–1854, December 1966. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol45/bstj45-10-1851.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol45-1966/articles/
                                        bstj45-10-1851.pdf.

Spaulding:1968:SRT

[Spa68] David A. Spaulding. Synthesis of rational transfer function approximations using a tapped distributed RC line with feedback. The Bell System Technical Journal, 47(9):2051–2070, November 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-9-2051.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-9-2051.pdf.

Spaulding:1969:SPS

[Spa69] David A. Spaulding. Synthesis of pulse-shaping networks in the time domain. The Bell System Technical Journal, 48(7):2425–2444, September 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-7-2425.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-7-2425.pdf.

Shennum:1963:DCE

[SR63] R. H. Shennum and E. J.

Reid. The design and construction of the electronics package. The Bell System Technical Journal, 42(4):1749–1763, July 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-4-1749.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-4-1749.pdf.

Sze:1967:NRC

[SR67] S. M. Sze and R. M. Ryder. The nonlinearity of the reverse current-voltage characteristics of a p–n junction near avalanche breakdown. The Bell System Technical Journal, 46(6):1135–1139, July/ August 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol46/bstj46-6-1135.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-6-1135.pdf.

Sperry:1960:TET

[SS60] R. V. Sperry and D. Surenian. A transversal equalizer for television circuits. The Bell System Technical Journal, 39(2):405–422, March 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol39/bstj39-2-405.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol39-1960/
                                        articles/bstj39-2-405.pdf.

Slepian:1965:EAP

[SS65] David Slepian and Estelle Sonnenblick. Eigenvalues associated with prolate spheroidal wave functions of zero order. The Bell System Technical Journal, 44(8):1745–1759, October 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/

images/Vol44/bstj44-8-1745.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-8-1745.pdf.

Sze:1967:UCE

[SS67] S. M. Sze and W. Shockley. Unit-cube expression for space-charge resistance. The Bell System Technical Journal, 46(5):837–842, May/June 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol46/bstj46-5-837.
pdf;   http://www.alcatel-
lucent.com/bstj/vol46-1967/
articles/bstj46-5-837.pdf.

Sandberg:1968:NIS

[SS68] I. W. Sandberg and H. Shichman.

Numerical integration of systems of stiff nonlinear differential equations. The Bell System Technical Journal, 47(4):511–527, April 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-4-511.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-4-511.pdf.

Saltzberg:1969:DTE

[SS69a] B. R. Saltzberg and M. K. Simon. Data transmission error probabilities in the presence of low-frequency removal and noise. The Bell System Technical Journal, 48(1):255–273, January 1969. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-1-255.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-1-255.pdf.

Sittig:1969:RUS

[SS69b] E. K. Sittig and F. M. Smits. Recirculating ultrasonic stores: An economical approach to sequential storage with bit rates beyond 100 MHz. The Bell System Technical Journal, 48(3):659–674, March 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-3-659.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-3-659.pdf.

Smith:1969:HPS

[SSD69] Peter W. Smith, M. V. Schneider, and Hans G. Danielmeyer. Highpower single-frequency lasers using thin metal film modeselection filters. The Bell System Technical Journal, 48 (5):1405–1419, May/June 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-5-1405.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-5-1405.pdf.

Semplak:1969:SMA

[ST69] R. A. Semplak and R. H. Turrin. Some measurements of attenuation by rainfall at 18.5 GHz. The Bell System Technical Journal, 48(6):1767–1787, July/ August 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-6-1767.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-6-1767.pdf.

Stagg:1965:LC

[Sta65] U. K. Stagg. The 2A line concentrator. The Bell System Technical Journal, 44(8):1535–1566, October 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-8-1535.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-8-1535.pdf.

Standley:1967:MWT

[Sta67] R. D. Standley. A millimeter wave, two-pole, circularelectric mode, channel-dropping filter structure. The Bell System Technical Journal, 46(10): 2261–2276, December 1967. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol46/bstj46-10-2261.pdf;
http://www.alcatel-lucent.
com/bstj/vol46-1967/articles/
bstj46-10-2261.pdf.

Steinberg:1961:ROM

[Ste61]  Priscilla L. Steinberg.  The

resistance of organic materials to attack by marine bacteria at low temperatures. The Bell System Technical Journal, 40(5):1369–1395, September 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol40/bstj40-5-1369.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol40-1961/
                                        articles/bstj40-5-1369.pdf.

Steier:1965:AHH

[Ste65] William H. Steier. The attenuation of the Holmdel helix waveguide in the 100–125-Kmc band. The Bell System Technical Journal, 44(5):899–906, May/ June 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol44/bstj44-5-899.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol44-1965/
                                        articles/bstj44-5-899.pdf.

Steier:1966:SER

[Ste66] William H. Steier. The statistical effects of random variations in the components of a beam waveguide. The Bell System Technical Journal, 45(3):451–471, March 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol45/bstj45-3-451.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol45-1966/
articles/bstj45-3-451.pdf.

Sunde:1961:PTF

[Sun61] E. D. Sunde. Pulse transmission by AM, FM, and PM in the presence of phase distortion. The Bell System Technical Journal, 40(2):353–422, March 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-2-353.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-2-353.pdf.

Sunde:1964:DTT

[Sun64a] E. D. Sunde. Digital troposcatter transmission and modulation theory. The Bell System Technical Journal, 43(1):143–214, January 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-1-143.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-1-143.pdf.

Sunde:1964:IDA

[Sun64b] E. D. Sunde. Intermodulation distortion in analog FM troposcatter systems. The Bell System Technical Journal, 43(1):399–435, January 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-1-399.

                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-1-399.pdf.

Suyderhoud:1965:GSD

[Suy65] H. G. Suyderhoud. A general statistical determination of transmission characteristics applied to L multiplex. The Bell System Technical Journal, 44(6):1031–1059, July/ August 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol44/bstj44-6-1031.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol44-1965/
                                        articles/bstj44-6-1031.pdf.

Spencer:1969:SOO

[SV69] A. E. Spencer and F. S. Vigilante. System organization and objectives. The Bell System Technical Journal, 48(8):2607–2618, October 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-8-2607.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-8-2607.pdf.

Skrabal:1968:MRE

[SW68] R. J. Skrabal and J. A. Word. Microwave radio equipment and building considerations. The Bell System Technical Journal, 47(7):1511–1530, September 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-7-1511.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-7-1511.pdf.

Sandberg:1969:SNT

[SW69a] I. W. Sandberg and A. N.

Willson, Jr. Some networktheoretic properties of nonlinear DC transistor networks. The Bell System Technical Journal, 48(5):1293–1311, May/ June 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-5-1293.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-5-1293.pdf.

Sandberg:1969:STP

[SW69b] I. W. Sandberg and A. N. Willson, Jr. Some theorems on properties of DC equations of nonlinear networks. The Bell System Technical Journal, 48(1):1– 34, January 1969. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol48/bstj48-1-
1.pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-1-1.pdf.

Tabor:1967:HCD

[Tab67] W. J. Tabor. A high-capacity digital light deflector using wollaston prisms. The Bell System Technical Journal, 46(5):957–990, May/ June 1967. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol46/bstj46-5-957.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-5-957.pdf.

Takacs:1961:SPB

[Tak61] Lajos Takács. Stochastic processes with balking in the theory of telephone traffic. The Bell System Technical Journal, 40(3):795–820, May 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol40/bstj40-3-795.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol40-1961/
                                        articles/bstj40-3-795.pdf.

Takacs:1962:DDS

[Tak62] Lajos Takács. Delay distributions for simple trunk groups with recurrent input and exponential service times. The Bell System Technical Journal, 41(1):311–320, January 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol41/bstj41-1-311.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol41-1962/
                                        articles/bstj41-1-311.pdf.

Takacs:1963:DDO

[Tak63a] L. Takács. Delay distributions for one line with Poisson input, general holding times, and various orders of service. The Bell System Technical Journal, 42(2):487–503, March 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-2-487.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-2-487.pdf.

Takacs:1963:SSQ

[Tak63b] L. Takács. A single-server queue with feedback. The Bell System Technical Journal, 42(2):505–519, March 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-2-505.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-2-505.pdf.

Taylor:1968:RCC

[Tay68a] Michael G. Taylor. Reliable computation in computing systems designed from unreliable components. The Bell System Technical Journal, 47(10):2339– 2366, December 1968. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol47/bstj47-10-2339.pdf;
http://www.alcatel-lucent.
com/bstj/vol47-1968/articles/
bstj47-10-2339.pdf.

Taylor:1968:RIS

[Tay68b] Michael G. Taylor. Reliable information storage in memories designed from unreliable components. The Bell System Technical Journal, 47(10):2299– 2337, December 1968. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol47/bstj47-10-2299.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol47-1968/articles/
                                        bstj47-10-2299.pdf.

Thomas:1964:ADP

[TC64] L. C. Thomas and J. O. Cappellari. Attitude determination and prediction of spin-stabilized satellites. The Bell System Technical Journal, 43(4):1657–1726, July 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-4-1657.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-4-1657.pdf.

Thiele:1969:TCM

[Thi69] A. A. Thiele. The theory of cylindrical magnetic domains. The Bell System Technical Journal, 48(10):3287–3335, December 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

                                        bell-labs.com/BSTJ/images/
                                        Vol48/bstj48-10-3287.pdf;
                                        http://www.alcatel-lucent.

com/bstj/vol48-1969/articles/
bstj48-10-3287.pdf.

Thomas:1963:NCP

[Tho63]  D. E. Thomas.  Numerical

computation of phase from amplitude at optical frequencies. The Bell System Technical Journal, 42(3):637–679, May 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-3-637.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-3-637.pdf.

Thomas:1965:GOE

[Tho65] L. C. Thomas. Generation of orbital elements for the TEL-STAR communications satellites. The Bell System Technical Journal, 44(4):603–673, April 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-4-603.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-4-603.pdf.

Thomas:1968:EBD

[Tho68]  Lee C. Thomas.  Eliminat-

ing broadband distortion in transistor amplifiers. The Bell System Technical Journal, 47(3):315–342, March 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/

                                        images/Vol47/bstj47-3-315.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-3-315.pdf.

Tillotson:1968:MDS

[Til68] Leroy C. Tillotson. A model of a domestic satellite communication system. The Bell System Technical Journal, 47(10): 2111–2137, December 1968. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol47/bstj47-10-2111.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol47-1968/articles/
                                        bstj47-10-2111.pdf.

Tillotson:1969:UFA

[Til69]  LeRoy C. Tillotson.  Use

of frequencies above 10 GHz for common carrier applications. The Bell System Technical Journal, 48(6):1563–1576, July/ August 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-6-1563.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-6-1563.pdf.

Thurmond:1960:GSL

[TK60] C. D. Thurmond and M. Kowalchik. Germanium and silicon liquidus curves. The Bell System Technical Journal, 39(1):169–204, January 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-1-169.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-1-169.pdf.

Tong:1966:SRT

[Ton66]  S. Y. Tong.  Synchroniza-

tion recovery techniques for binary cyclic codes. The Bell System Technical Journal, 45(4):561–596, April 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-4-561.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-4-561.pdf.

Trumbore:1960:SSI

[Tru60] F. A. Trumbore. Solid solubilities of impurity elements in germanium and silicon. The Bell System Technical Journal, 39(1):205–233, January 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-1-205.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-1-205.pdf.

Tabor:1963:MTS

[TS63] W. J. Tabor and J. T. Sibilia.

Masers for the Telstar satellite communications experiment. The Bell System Technical Journal, 42(4):1863–1886, July 1963. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-4-1863.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-4-1863.pdf.

Tsiang:1962:ATD

[TU62] S. H. Tsiang and W. Ulrich. Automatic trouble diagnosis of complex logic circuits. The Bell System Technical Journal, 41(4):1177–1200, July 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol41/bstj41-4-1177.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol41-1962/
                                        articles/bstj41-4-1177.pdf.

Townsend:1964:EEC

[TW64] R. L. Townsend and R. N. Watts. Effectiveness of error control in data communication over the switched telephone network. The Bell System Technical Journal, 43(6):2611–2638, November 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-6-2611.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-6-2611.pdf.

Travis:1965:WDT

[TY65] L. F. Travis and R. E. Yaeger. Wideband data on T1 carrier. The Bell System Technical Journal, 44(8):1567–1604, October 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-8-1567.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-8-1567.pdf.

Uenohara:1963:GPA

[UCE^{+}63] M. Uenohara, M. Chruney, K. M.

Eisele, D. C. Hanson, and A. L. Stillwell. 4-gc parametric amplifier for satellite communication ground station receiver. The Bell System Technical Journal, 42(4):1887–1908, July 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol42/bstj42-4-1887.
pdf;   http://www.alcatel-
lucent.com/bstj/vol42-1963/
articles/bstj42-4-1887.pdf.

Upthegrove:1963:LTS

[UDKM63] H. N. Upthegrove, J. B. D’Albora, Jr., A. R. Kolding, and B. A. McLeod. Launching of the Telstar satellite. The Bell System Technical Journal, 42(4): 1449–1473, July 1963. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol42/bstj42-4-1449.pdf;
http://www.alcatel-lucent.
com/bstj/vol42-1963/articles/
bstj42-4-1449.pdf.

Unger:1960:RWD

[Ung60]  Hans-Georg Unger.  Round

waveguide with double lining. The Bell System Technical Journal, 39(1):161–167, January 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol39/bstj39-1-161.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol39-1960/
                                        articles/bstj39-1-161.pdf.

Unger:1961:MCM

[Ung61a] H. G. Unger. Mode conversion in metallic and helix waveguide. The Bell System Technical Journal, 40(2):613–626, March 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol40/bstj40-2-613.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol40-1961/
                                        articles/bstj40-2-613.pdf.

Unger:1961:NHW

[Ung61b] H. G. Unger. Noncylindrical helix waveguide. The Bell System Technical Journal, 40(1):233–254, January 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol40/bstj40-1-233.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol40-1961/
                                        articles/bstj40-1-233.pdf.

Unger:1961:NMM

[Ung61c] H. G. Unger. Normal modes and mode conversion in helix waveguide. The Bell System Technical Journal, 40(1):255–280, January 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-1-255.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-1-255.pdf.

Unger:1961:WTH

[Ung61d] H. G. Unger. Winding tolerances in helix waveguide. The Bell System Technical Journal, 40(2):627–643, March 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-2-627.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-2-627.pdf.

Unger:1962:LW

[Ung62] H. G. Unger. Lined waveguide. The Bell System Technical Journal, 41(2):745–768, March 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol41/bstj41-2-745.
pdf;   http://www.alcatel-
lucent.com/bstj/vol41-1962/
articles/bstj41-2-745.pdf.

Unger:1966:RTA

[Ung66]  J. H. W. Unger.  Ran-

dom tropospheric angle errors in microwave observations of the early bird satellite. The Bell System Technical Journal, 45(9):1439–1474, November 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-9-1439.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-9-1439.pdf.

Uenohara:1961:PEM

[US61] M. Uenohara and H. Seidel. Project Echo: 961-mc lower-sideband up-converter for satellite-tracking radar. The Bell System Technical Journal, 40(4):1183–1205, July 1961. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol40/bstj40-4-1183.
pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-4-1183.pdf.

Urlich:1964:TNE

[UV64] W. Urlich and Mrs. H. M. Vellenzer. Translations in the no. 1 Electronic Switching System. The Bell System Technical Journal, 43(5):2533–2573, September 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-5-2533.

pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-5-2533.pdf.

Ullman:1969:USS

[UW69] J. D. Ullman and Peter Weiner.

Uniform synthesis of sequential circuits. The Bell System Technical Journal, 48(5): 1115–1127, May/June 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-5-1115.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-5-1115.pdf.

VanRoosbroeck:1960:TCC

[Van60]  W. Van Roosbroeck.  The-

ory of current-carrier transport and photoconductivity in semiconductors with trapping. The Bell System Technical Journal, 39(3):515–613, May 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-3-515.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-3-515.pdf.

Vartabedian:1966:ETD

[Var66] A. G. Vartabedian. The effects of transmission delay in four-wire teleconferencing. The Bell System Technical Journal, 45(10): 1673–1688, December 1966. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol45/bstj45-10-1673.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol45-1966/articles/
                                        bstj45-10-1673.pdf.

Varnerin:1963:SFP

[VCDK63] L. J. Varnerin, R. L. Comstock, W. A. Dean, and R. W. Kordos. The satellite ferrimagnetic power limiter. The Bell System Technical Journal, 42(4):1817–1827, July 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-4-1817.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-4-1817.pdf.

Varon:1967:SPL

[VZ67] D. Varon and G. I. Zysman. Some properties and limitations of electronically steerable phased array antennas. The Bell System Technical Journal, 46(7):1561–1586, September 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol46/bstj46-7-1561.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-7-1561.pdf.

Wang:1961:PTP

[Wan61]  Hao Wang.  Proving theo-

rems by pattern recognition — II. The Bell System Technical Journal, 40(1):1–41, January 1961. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol40/bstj40-1-
1.pdf;   http://www.alcatel-
lucent.com/bstj/vol40-1961/
articles/bstj40-1-1.pdf.

Wang:1968:IDW

[Wan68] H. C. Wang. An improved design of waveguide band-rejection filters. The Bell System Technical Journal, 47(1):1–15, January 1968. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol47/bstj47-1-
1.pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-1-1.pdf.

Warner:1960:DPU

[War60] A. W. Warner. Design and performance of ultraprecise 2.5mc quartz crystal units. The Bell System Technical Journal, 39(5):1193–1217, September 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-5-1193.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-5-1193.pdf.

Warthman:1961:PEB

[War61]  K. L. Warthman.  Project

Echo: Boresight cameras. The Bell System Technical Journal, 40(4):1227–1233, July 1961.

CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol40/bstj40-4-1227.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol40-1961/
                                        articles/bstj40-4-1227.pdf.

Williams:1969:ATU

[WC69] W. H. Williams and H. Chen. An analysis of time usage in Bell System business offices. The Bell System Technical Journal, 48(7):2493–2506, September 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-7-2493.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-7-2493.pdf.

Weber:1962:STC

[Web62] J. H. Weber. Some traffic characteristics of communications networks with automatic alternate routing. The Bell System Technical Journal, 41(2):769–796, March 1962. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol41/bstj41-2-769.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol41-1962/
                                        articles/bstj41-2-769.pdf.

Weber:1964:SSR

[Web64] J. H. Weber. A simulation study of routing and control in communications networks. The Bell System Technical Journal, 43(6):2639–2676, November 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-6-2639.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-6-2639.pdf.

Weger:1960:PEP

[Weg60] M. Weger. Passage effects in paramagnetic resonance experiments. The Bell System Technical Journal, 39(4):1013–1112, July 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol39/bstj39-4-1013.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-4-1013.pdf.

Weiner:1965:TMC

[Wei65] R. S. Weiner. Transient motion of circular elastic plates subjected to impulsive and moving loads. The Bell System Technical Journal, 44(10):2409– 2431, December 1965. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol44/bstj44-10-2409.pdf;
http://www.alcatel-lucent.
com/bstj/vol44-1965/articles/
bstj44-10-2409.pdf.

Weissmann:1966:MMV

[Wei66] G. F. Weissmann. A mathematical model of a vibrating soil-foundation system. The Bell System Technical Journal, 45(1):177–228, January 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-1-177.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-1-177.pdf.

Weldon:1966:BJB

[Wel66a] E. J. Weldon, Jr. B.S.T.J. briefs: Performance of a forwardacting error-control system on the switched telephone network. The Bell System Technical Journal, 45(5):759–761, May/ June 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-5-759.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-5-759.pdf.

Weldon:1966:DSC

[Wel66b] E. J. Weldon, Jr. Differenceset cyclic codes. The Bell System Technical Journal, 45 (7):1045–1055, September 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-7-1045.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-7-1045.pdf.

Wu:1968:SWE

[WG68] C. P. Wu and V. Galindo.

Surface-wave effects on dielectric sheathed phased arrays of rectangular waveguides. The Bell System Technical Journal, 47(1):117–142, January 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-1-117.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-1-117.pdf.

White:1966:BJB

[Whi66] A. D. White. B.S.T.J. briefs:

Laser cavities with increased axial mode separation. The Bell System Technical Journal, 45(2):339–343, February 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-2-339.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-2-339.pdf.

Willson:1968:SEN

[Wil68] A. N. Willson, Jr. On the solutions of equations for nonlinear resistive networks. The Bell System Technical Journal, 47(8):1755–1773, October 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-8-1755.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-8-1755.pdf.

Wilson:1969:STM

[Wil69] Robert W. Wilson. Sun tracker measurements of attenuation by rain at 16 and 30 GHz. The Bell System Technical Journal, 48(5):1383–1404, May/ June 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-5-1383.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-5-1383.pdf.

Witt:1965:EMD

[Wit65] F. J. Witt. An experimental 224 Mb/s digital multiplexerdemultiplexer using pulse stuffing synchronization. The Bell System Technical Journal, 44(9):1843–1885, November 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol44/bstj44-9-1843.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol44-1965/
                                        articles/bstj44-9-1843.pdf.

Wojciechowski:1960:TFS

[Woj60] B. M. Wojciechowski. Theory of a frequency-synthesizing network. The Bell System Technical Journal, 39(3):649–673, May 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/

images/Vol39/bstj39-3-649.
pdf;   http://www.alcatel-
lucent.com/bstj/vol39-1960/
articles/bstj39-3-649.pdf.

Wolman:1965:DCC

[Wol65] Eric Wolman. On definitions of congestion in communication networks. The Bell System Technical Journal, 44(10): 2271–2294, December 1965. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol44/bstj44-10-2271.pdf;
http://www.alcatel-lucent.
com/bstj/vol44-1965/articles/
bstj44-10-2271.pdf.

Wolf:1969:ATI

[Wol69] Jack Keil Wolf. Adding two information symbols to certain nonbinary BCH codes and some applications. The Bell System Technical Journal, 48(7):2405–2424, September 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-7-2405.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-7-2405.pdf.

Wood:1964:DGT

[Woo64] Elizabeth A. Wood. The 80 diperiodic groups in three dimensions. The Bell System Technical Journal, 43(1):541–559, January 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol43/bstj43-1-541.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol43-1964/
                                        articles/bstj43-1-541.pdf.

Wu:1969:RWA

[Wu69]  C. P. Wu.  Resonances in

waveguide antennas with dielectric plugs. The Bell System Technical Journal, 48(7): 2305–2324, September 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol48/bstj48-7-2305.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol48-1969/
                                        articles/bstj48-7-2305.pdf.

Warters:1968:PGM

[WY68] W. D. Warters and D. T. Young. Precise 50 to 60 GHz measurements on a two-mile loop of helix waveguide. The Bell System Technical Journal, 47(6):933–955, July/August 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol47/bstj47-6-933.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-6-933.pdf.

Wyner:1964:BJB

[Wyn64] A. D. Wyner. B.S.T.J. briefs: An improved error bound for Gaussian channels. The Bell System Technical Journal, 43(6):3070–3075, November 1964. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol43/bstj43-6-3070.
pdf;   http://www.alcatel-
lucent.com/bstj/vol43-1964/
articles/bstj43-6-3070.pdf.

Wyner:1965:CBD

[Wyn65] A. D. Wyner. Capabilities of bounded discrepancy decoding. The Bell System Technical Journal, 44(6):1061–1122, July/ August 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-6-1061.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-6-1061.pdf.

Wyner:1966:BCP

[Wyn66a] A. D. Wyner. Bounds on communication with polyphase coding. The Bell System Technical Journal, 45(4):523–559, April 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-4-523.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-4-523.pdf.

Wyner:1966:CBL

[Wyn66b] A. D. Wyner. The capacity of the band-limited Gaussian channel. The Bell System Technical Journal, 45(3):359–395, March 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol45/bstj45-3-359.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol45-1966/
                                        articles/bstj45-3-359.pdf.

Wyner:1967:RPC

[Wyn67] A. D. Wyner. Random packings and coverings of the unit n-sphere. The Bell System Technical Journal, 46(9): 2111–2118, November 1967. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol46/bstj46-9-2111.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol46-1967/
                                        articles/bstj46-9-2111.pdf.

Wyner:1968:CAD

[Wyn68] A. D. Wyner. Communication of analog data from a Gaussian source over a noisy channel. The Bell System Technical Journal, 47(5):801–812, May/ June 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol47/bstj47-5-801.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-5-801.pdf.

Wyner:1969:PEU

[Wyn69a] A. D. Wyner. On the probability of error using a repeatrequest strategy on the additive white Gaussian noise channel. The Bell System Technical Journal, 48(1):71–86, January 1969. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol48/bstj48-1-
71.pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-1-71.pdf.

Wyner:1969:DCD

[Wyn69b] Aaron D. Wyner. On digital communication over a discretetime Gaussian channel with noisy feedback. The Bell System Technical Journal, 48(10):3173– 3186, December 1969. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol48/bstj48-10-3173.pdf;
http://www.alcatel-lucent.
com/bstj/vol48-1969/articles/
bstj48-10-3173.pdf.

Wynn:1969:SOS

[Wyn69c] W. D. Wynn. A second order statistical analysis of the operation of a limiter-phase detector-filter cascade. The Bell System Technical Journal, 48(1):233–248, January 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-1-233.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-1-233.pdf.

Wyner:1969:CAD

[WZ69] Aaron D. Wyner and Jacob Ziv. On communication of analog data from a bounded source space. The Bell System Technical Journal, 48(10): 3139–3172, December 1969. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol48/bstj48-10-3139.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol48-1969/articles/
                                        bstj48-10-3139.pdf.

Young:1963:EDL

[You63a] D. T. Young. Effect of differential loss on approximate solutions to the coupled line equations. The Bell System Technical Journal, 42(6):2787–2793, November 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-6-2787.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-6-2787.pdf.

Young:1963:MRC

[You63b] D. T. Young. Model for relating coupled power equations to coupled amplitude equations. The Bell System Technical Journal, 42(6):2761–2764, November 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol42/bstj42-6-2761.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol42-1963/
articles/bstj42-6-2761.pdf.

Young:1965:MTA

[You65a] D. T. Young. Measured TE01 attenuation in helix waveguide with controlled straightness deviations. The Bell System Technical Journal, 44(2):273–282, February 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-2-273.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-2-273.pdf.

Young:1965:EDM

[You65b] Tzay Y. Young. Epoch detection. A method for resolving overlapping signals. The Bell System Technical Journal, 44(3):401–426, March 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-3-401.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-3-401.pdf.

Yu:1963:SDS

[Yu63] E. Y. Yu. Spin decay, spinprecession damping, and spinaxis drift of the Telstar satellite. The Bell System Technical Journal, 42(5):2169–2193, September 1963. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/

                                        images/Vol42/bstj42-5-2169.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol42-1963/
                                        articles/bstj42-5-2169.pdf.

Yu:1965:ODG

[Yu65] E. Y. Yu. Optimum design of a gravitationally oriented twobody satellite. The Bell System Technical Journal, 44(1):49–76, January 1965. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol44/bstj44-1-
                                        49.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol44-1965/
                                        articles/bstj44-1-49.pdf.

Zador:1966:BJB

[Zad66a] P. L. Zador. B.S.T.J. briefs: Realizability conditions for the impedance function of the lossless tapered transmission line. The Bell System Technical Journal, 45(9):1667–1669, November 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol45/bstj45-9-1667.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol45-1966/
                                        articles/bstj45-9-1667.pdf.

Zador:1966:EPD

[Zad66b] P. L. Zador. Error probabilities in data system pulse regenerator with DC restoration. The Bell System Technical Journal, 45(6):979–984, July/ August 1966. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol45/bstj45-6-979.
pdf;   http://www.alcatel-
lucent.com/bstj/vol45-1966/
articles/bstj45-6-979.pdf.

Zucker:1968:CAA

[ZI68] H. Zucker and W. H. Ierley. Computer-aided analysis of Cassegrain antennas. The Bell System Technical Journal, 47(6):897–932, July/August 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol47/bstj47-6-897.
pdf;   http://www.alcatel-
lucent.com/bstj/vol47-1968/
articles/bstj47-6-897.pdf.

Ziv:1969:BBS

[Ziv69] Jacob Ziv. Bounds on the bias of signal parameter estimators. The Bell System Technical Journal, 48(6):2023–2030, July/ August 1969. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol48/bstj48-6-2023.
pdf;   http://www.alcatel-
lucent.com/bstj/vol48-1969/
articles/bstj48-6-2023.pdf.

Zucker:1968:GAR

[Zuc68] H. Zucker. Gain of antennas with random surface deviations. The Bell System Technical Journal, 47(8):1637–1651, October 1968. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol47/bstj47-8-1637.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol47-1968/
                                        articles/bstj47-8-1637.pdf.

Zupa:1960:MLC

[Zup60] F. A. Zupa. Magnetic latching crossbar switches: A new development in magnetic properties of tool steel. The Bell System Technical Journal, 39(5):1351–1374, September 1960. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol39/bstj39-5-1351.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol39-1960/
                                        articles/bstj39-5-1351.pdf.